### v1

# Undergraduate Course Allocation

Abstract
Prevailing methods of course allocation at undergraduate institutions involve reserving seats to give priority to designated groups of students. We introduce a competitive equilibrium-based mechanism that assigns course seats using student preferences and course priorities. This mechanism satisfies approximate notions of stability, efficiency, envy-freeness, and strategy-proofness. We evaluate its performance relative to a mechanism widely used in practice using preferences estimated from university data. Our empirical findings demonstrate an improvement in student satisfaction and allocation fairness. The number of students who envy another student of weakly lower priority declines by 8 percent, or roughly 500 students.
JEL classification: D47, D63, D82, C63, C78, I21
Keywords: market design, many-to-many matching, course allocation, approximate competitive equilibrium, random serial dictatorship, reserved seats, priorities
Every academic term, over 6,500 undergraduate institutions across the United States assign course schedules to a total of nearly twenty million students. [^1] Based on National Center for Education Statistics, 2017-2018. Based on course timing and prerequisites, the set of schedules a student can be assigned in a given term is limited. Moreover, student preferences over possible schedules are heterogeneous. In the face of room size and teaching constraints, universities use factors such as seniority and student department to complete the challenging task of allocating seats in overdemanded courses.
Course allocation is a well-studied problem in the context of graduate business schools. Many institutions employ course bidding systems that require strategic decision-making from students, often leading to inefficient allocations (Sönmez and Ünver, 2010). Recently, several business schools replaced course bidding systems with the approximate competitive equilibrium from equal incomes mechanism (Budish, 2011; Budish et al., 2023), which seeks to alleviate these issues. This mechanism elicits student preferences over course schedules, assigns students almost-equal budgets of fake money, and internally determines an approximate competitive equilibrium, in which each student receives her most-preferred affordable course schedule subject to course prices. The competitive equilibrium is approximate due to a small market-clearing error, measured by the assignment of a course’s seats over and under its capacity. This mechanism treats all business school students equally, as students are randomly assigned budgets and face the same course prices.
At undergraduate institutions, students are not treated equally, making course allocation a two-sided matching problem. Each course divides students into priority levels based on factors such as seniority and student department. In commonly-used course allocation mechanisms, students select courses in order of seniority (with ties broken randomly), and priorities based on student department are enforced by reserving seats in each course. Top US universities that closely follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. [^2] In general, there is some heterogeneity in how universities elect to assign seats to students. Some universities use priorities (e.g., Dartmouth) or reserves (e.g., Vanderbilt) to assign seats, whereas others delegate the decision to departments (e.g., Cornell). Some use two phases (e.g., UCLA), only allowing students to enroll in a limited number of courses in the first phase. Seniority can be based on academic year (e.g., Carnegie Mellon), number of earned credits (e.g., University of North Carolina), or time to graduation (e.g., Washington University in St. Louis). At some universities, few classes are over-enrolled (e.g., Caltech), whereas at others, the process is rather congested (e.g., UC Berkeley). In most cases, there are a small number of priority levels, rather than one (Budish, 2011) or complete strict priorities Roth (1984).
In this paper, we approach the task of allocating course schedules to students in the context of a many-to-many two-sided matching problem with heterogeneous student preferences and weak course priorities. We propose a novel deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that uses fake money and competitive equilibrium to allocate course seats to students without transfers. This mechanism elicits student preferences for course schedules, assigns almost equal budgets to students, and finds an approximate competitive equilibrium allocation under priority-specific prices. The equilibrium prices respect course priorities by setting a cutoff priority level for each course, where students at the cutoff priority level pay a non-negative price, students above the cutoff pay nothing, and students below the cutoff cannot afford the course.
The approximate competitive equilibrium from equal incomes mechanism is the special case of our mechanism in which no course prioritizes any student over any others. Our first result shows that, with arbitrary course priorities, the PMP mechanism maintains the same, small worst-case bound on the market-clearing error as Budish (2011). This worst-case bound is increasing in the dimension of the space of prices, which equals the number of courses in settings without priorities. Our result holds despite the expansion from one price per course to priority-specific prices in each course. The key insight of the proof is the introduction of a parameterization within the priority-specific price space. The parameterization simplifies the space of prices to a single value per course, enabling a simultaneous search for both the cutoff priority level and the corresponding price. This strategy simplifies the complexity of the problem and allows for the application of some previously established results.
The PMP mechanism has several advantageous theoretical properties. Adjusting course capacities to accommodate over-enrolled courses, the PMP outcome ensures that no student can be assigned a preferred schedule without violating course priorities or capacity constraints (approximate stability). It is also not possible to reassign course seats among a group of students, benefitting some members and hurting none, while ensuring that the outcome respects course priorities to the same degree (approximate priority-constrained efficiency). Among similarly prioritized students, the PMP outcome limits envy; if a student prefers the schedule assigned to a student with the same or a lower priority level in each course, by removing one course from the preferred schedule, the student now prefers her own (schedule envy bounded by a single course). Finally, in large populations, the PMP mechanism is resistant to strategic manipulations by students (strategy-proof in the large).
To better guide practitioners in the course allocation process, we compare the performance of the PMP mechanism with several alternative mechanisms. As a benchmark, we consider a mechanism, the Random Serial Dictatorship with course reserves (RSD) mechanism, which closely resembles what is used in practice. Students select courses in order of seniority (with ties broken randomly) and priorities based on student department and year of study are enforced by reserving seats in each course. [^3] In practice, courses set a large number of reserved seats at the beginning of the registration process. After students initially register, university departments relax the course reserve constraints to ensure full enrollment. We do not model this period or the process of relaxing reserved seats. Instead, we calculate the optimal course reserves that a university registrar should set at the beginning of the process if the number of reserved seats cannot be changed. We also consider the Deferred Acceptance mechanisms with single tie-breaking and multiple tie-breakings (denoted as DA-STB and DA-MTB, respectively). These mechanisms allocate courses to students by extending the Gale-Shapley algorithm (Gale and Shapley, 1962) to many-to-many matching, where priority ties are resolved by either a single tie-breaking rule or course-specific tie-breakings. The Deferred Acceptance mechanism has been successful in many practical applications, including matching applicants to residency programs, students to schools, and workers to job positions (see Roth, 2018). We analyze six additional mechanisms in Appendix D.
We use university data on the final allocation of 6 02360236\,0236 023 students to $756$ courses from the Spring 2018 term at a private institution in the mid-Atlantic region. In total, 23 3692336923\,36923 369 course seats are occupied, with $11.2\%$ of courses enrolled at or above their reported capacities. We use student and course characteristics to estimate a model of student utilities and evaluate the performance of the PMP, RSD, DA-STB, and DA-MTB mechanisms in terms of each student’s preferred mechanism, the standard deviation of student utilities, and the percentage of students who experience schedule envy. The first metric compares each student’s satisfaction from her course schedule across the four mechanisms. The two other metrics compare the fairness of each mechanism’s outcome across students.
Our results demonstrate that the PMP mechanism delivers a substantially fairer outcome than the RSD benchmark. Considerably fewer students - $8\%$ of the total population, or almost $500$ individuals - have envy toward a student of weakly lower priority in each course. No students have envy greater than a single course, as opposed to $3\%$ in the benchmark mechanism. Furthermore, the standard deviation of student utilities declines between $1.1\%$ and $5.6\%$ across years of study. Each of these results are weaker for the DA-STB and DA-MTB mechanisms, with the DA-MTB mechanism delivering a slightly better performance.
These fairness improvements come without sacrificing student satisfaction. For second- through fourth-year students, a larger percentage of students prefer their assigned schedule in the PMP mechanism than students who prefer the RSD benchmark. The difference between these groups ranges from $5.0\%$ and $8.9\%$ across years of study, with $59.2\%-79.1\%$ of students being indifferent. Among students receiving different schedules, the average increase in utility ranges from $5.1\%$ and $9.4\%$. For first-year students, $0.9\%$ more prefer the benchmark. This result is attributed to the gains for older students, who are assigned an average of $135$ more seats. Still, the PMP mechanism provides an $0.8\%$ average increase in their utility.
Literature review. This paper contributes to the literature analyzing the allocation of objects to agents through competitive equilibria. As competitive equilibrium from equal incomes are not guaranteed to exist, many papers focus on random allocation mechanisms. Hylland and Zeckhauser (1979) first proposed the use of fake money and competitive equilibrium to randomly allocate objects to agents through what are referred to as pseudo-market mechanisms. He et al. (2018) incorporate a priority structure into pseudo-markets and analyze random mechanisms with a focus on unit-demand settings such as school choice. [^4] He et al. (2018) explain how their results extend in many-to-many settings with additive utility. Miralles and Pycia (2021), Echenique, Miralles and Zhang (2021) and Nguyen, Nguyen and Teytelboym (2021) consider pseudo-market solutions to random allocation problems with complex constraints. Pycia (2021) provides an excellent survey of the use of pseudo-markets for random allocations in environments without transfers.
In contrast to these papers, we only consider deterministic mechanisms and study a many-to-many matching problem. Budish (2011) introduces the idea of competitive markets that may only approximately clear. He proposes a deterministic mechanism, the approximate competitive equilibrium from equal incomes mechanism, which is approximately Pareto-efficient, strategy-proof in the large, and bounds student envy by a single course. This mechanism was successfully implemented at Wharton Business School and Columbia Business School (see Budish et al., 2017; Budish and Kessler, 2022). [^5] There are few other studies of real-world mechanisms. Budish and Cantillon (2012) draw on theory and field data to study a non-strategy proof course allocation mechanism at Harvard Business School. Bichler and Merting (2021) compare the performance of two random mechanisms at the Technical University of Munich. Rusznák, Biró and Fleiner (2021) analyze course allocation at a large university in Hungary. A key contribution of the present study is extending Budish’s approach to two-sided matching settings with weak course priorities, a central feature of the undergraduate course allocation problem.
Two recent papers also analyze deterministic assignments in many-to-many matching settings. Nguyen and Vohra (2022) establish the existence of a competitive equilibrium when all agent preferences satisfy a novel property of geometric substitutes, which is a strict generalization of the gross-substitutes property (see Kelso and Crawford, 1982). When there is an upper limit on the number of goods $k$ an agent can acquire, they prove that there always exists an approximate competitive equilibrium with no good over-enrolled by more than $k-1$ units. This “good-by-good” bound dominates the aggregate market-clearing bound of Budish (2011) when preferences are close substitutes, whereas the aggregate market-clearing bound dominates when student preferences are close complements. Though our main theoretical results follow the aggregate market-clearing bound approach, we ensure in our simulations that each approximate competitive equilibrium allocation found has no courses that are over-enrolled by more than $k-1$ seats.
Similarly, Lin et al. (2022) analyze deterministic allocation mechanisms through approximate competitive equilibria with a good-by-good bound. They consider complex feasibility constraints and prove the existence of an approximate competitive equilibrium that satisfies stability and fairness properties similar to the ones analyzed in this paper. We establish some additional properties of our mechanism, proving that it is strategy-proof in the large and results in an outcome that is approximately priority-constrained efficient. With simulations on randomly generated data, they compare their approach to draft mechanisms in the context of reassigning sporting event season tickets to families. We address the problem of allocating courses to undergraduate students by estimating preferences from real-world university data and comparing our proposed mechanism to a version of the mechanism used at many top universities in the United States. In Appendix D, we also provide a comparison of our mechanism with the deferred-acceptance mechanism with single and multiple tie-breakings Gale and Shapley (1962), the deferred-acceptance mechanism with minority reserves Hafalir, Yenmez and Yildirim (2013), some alternative pseudo-market mechanisms Budish (2011), and a variant of the probabilistic serial mechanism Bogomolnaia and Moulin (2001). Many of these mechanisms have been studied in many-to-one matching settings, but their properties in many-to-many matching with indifferences are not well-understood.
The remainder of the paper is organized as follows. We model undergraduate course allocation in Section 1 and investigate the properties of the PMP mechanism in Section 2. Our simulation results are presented in Section 3. We offer concluding remarks in Section 4. Appendices A and B present proofs, Appendix C details the student utility model calibration, and Appendices D and E provide additional simulation results and materials.
1 Environment
Course allocation is a many-to-many matching problem described by the tuple $({\cal S},{\cal C},Q,V,\mathcal{R})$.
  • ${\cal S}=\left\{1,...,N\right\}$ is a set of students; in reference to students, we use she/her pronouns.
  • ${\cal C}=\left\{1,...,M\right\}$ is a set of courses.
  • $Q=(q_{1},...,q_{M})$ is a vector of course capacities; each course $c\in{\cal C}$ has $q_{c}$ students.
  • $V=(\succsim_{1},....,\succsim_{N})$ is a vector of student preferences over course schedules: subsets of ${\cal C}$. Students typically consider only a subset of all possible course schedules due to factors such as course meeting times and prerequisites. Additionally, each student can take at most $k$ total course seats. We assume that these restrictions are incorporated into student preferences and that student preferences are strict. We also assume, for simplicity of exposition, that $21\leq k\leq M/$ 2. These restrictions still permit students to view courses as substitutes or complements.
  • =$\mathcal{R}=\{r_{s,c}\}_{s\in{\cal S},c\in{\cal C}}$ is a course priority structure. $r_{s,c}\in\{1,...,R\}$ specifies the level of priority of student $s$ in course $c$, with $R\geq 1$ and a larger number meaning a higher level of priority. Multiple students can share the same level of priority in a course.
Our setting connects two well-studied problems in market design in a natural way. If $r_{s,c}=1$ for all students $s\in{\cal S}$ and courses $c\in{\cal C}$, this is the combinatorial assignment problem of Budish (2011). If $\{r_{s,c}\}_{s\in{\cal S}}$ forms a strict ordering over all students for each course $c$, this is a two-sided matching problem in which students have strict preferences over subsets of courses and courses strictly rank individual students.
Two examples of course priority structures are year-specific and department-specific priorities. Universities often categorize students into four years of study based on semesters or credits completed. In addition, universities place students into departments, each of which offers a set of courses. In a year-specific priority structure, $R=4$, with $r_{s,c}$ equal to the year of study of student $s$ in every course $c$. In a department-specific priority structure, $R=2$, with $r_{s,c}=2$ if student $s$ and course $c$ are in the same department and $r_{s,c}=1$ otherwise. Year-specific priorities give an advantage to students closer to graduation in every course, whereas department-specific priorities give an advantage to students that require a given course to graduate over those that do not. We discuss hybrid priority structures that combine year-specific priorities and department-specific priorities in Section 3.
We consider deterministic allocations of courses to students. An allocation $x=(x_{s})_{s\in{\cal S}}$ assigns the course schedule $x_{s}$ to each student $s\in{\cal S}$. For ease of notation, we view $x_{s}$ as both a subset of courses in ${\cal C}$ and a vector from the set $\{0,1\}^{M}$. Allocation $x$ is feasible if $\sum_{s\in{\cal S}}x_{s,c}\leq q_{c}$ for each $c\in{\cal C}$, assigning no course a greater number of students than its capacity. Allocation $x$ is individually rational if, for every $s\in{\cal S}$, $x_{s}=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\}$, ensuring that $s$ does not want to drop any of her assigned courses. A pair $(s,C)$ of student $s\in{\cal S}$ and course schedule $C\subset{\cal C}$ is a block of $x$ if a student prefers $C$ to her current course schedule $x_{s}$ and each course in $C$ not in $x_{s}$ has an available seat or a lower priority student assigned to a seat. Formally, $(s,C)$ is a block if $C=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\cup C\}$, $C\neq x_{s}$, and for each $c\in C\setminus x_{s}$, $\sum_{s\in{\cal S}}x_{s,c}<q_{c}$ or there is a student $s^{\prime}\in{\cal S}$ with $c\in x_{s^{\prime}}$ and $r_{s,c}>r_{s^{\prime},c}$.
We evaluate allocations based on stability, efficiency, and fairness. An allocation $x$ is stable if it is feasible, individually rational, and admits no blocks (see Roth and Sotomayor, 1990; Echenique and Oviedo, 2006). [^6] A related concept is justified course envy. Allocation $x$ prevents justified course envy if there are no students $s,s^{\prime}\in{\cal S}$ and course $c\in{\cal C}$ such that $r_{s,c}>r_{s^{\prime},c},\,c\notin x_{s},\,c\in x_{s^{\prime}}$ and $c\in\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\cup c\}$. The absence of justified course envy prevents envy towards students of a lower priority, but does not account for individual rationality and the possibility of envy towards several students assigned to several courses.
In the next section, we analyze approximate market equilibria, in which some courses might have seats assigned slightly above capacity or empty demanded seats unfilled. We call such allocations having a market-clearing error. To account for the market-clearing error, we use an approximate version of stability.
Definition 1.
An allocation $x$ is approximately stable if it is stable in the environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$.
The relationship between stability and approximate stability is similar to the one between the notions of Pareto efficiency and approximate Pareto efficiency in Budish (2011). An allocation $y$ Pareto dominates an allocation $x$ if there is at least one student who strictly prefers her course schedule in $y$ and all other students weakly prefer their course schedules in $y$. Allocation $x$ is Pareto efficient if no allocation $y$ Pareto dominates $x$, and approximately Pareto efficient if no $y$ that Pareto dominates $x$ assigns weakly fewer seats in each course. Approximate Pareto efficiency is a meaningful notion of efficiency in environments where a market-clearing error is present.
In our setting, a meaningful notion of efficiency should also account for course priorities. Following Schlegel and Mamageishvili (2020), we extend course priorities over individual students to priorities over subsets of students using first-order stochastic dominance. We say that allocation $y$ dominates allocation $x$ for course $c$, and write $y_{c}\succeq_{c}x_{c}$, if for all levels of priority $r\in{\{1,...,R\}}$, $\sum_{s\in{\cal S}:r_{s,c}\geq r}y_{s,c}\geq\sum_{s\in{\cal S}:r_{s,c}\geq r}x_{s,c}$. Using this definition, we consider the following notion of priority-constrained efficiency.
Definition 2.
An allocation $x$ is approximately priority-constrained efficient if, in the environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$, for each allocation $y$ that Pareto dominates $x$, there is a course $c^{\prime}\in{\cal C}$ for which $y_{c^{\prime}}\nsucceq_{c^{\prime}}x_{c^{\prime}}$.
As with approximate stability, we account for a possible market-clearing error in allocation $x$ by adjusting the course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$. Except for this adjustment, our definition coincides with the priority-constrained efficiency introduced by Schlegel and Mamageishvili (2020).
The seminal measure of fairness is envy-freeness, as introduced by Foley (1967): an allocation $x$ prevents schedule envy if there are no students $s,s^{\prime}\in{\cal S}$ such that $x_{s^{\prime}}\;\succ_{s}\;x_{s}$. However, without using lotteries, we cannot hope to allocate courses in a completely envy-free way among students (e.g., if two students with the same priorities desire a course with a single seat). Extending the notion originally proposed by Budish (2011), we consider the more permissive concept of schedule envy bounded by a single course toward students of weakly lower priority in every course (see also Lin et al., 2022).
Definition 3.
An allocation $x$ has schedule envy bounded by a single course toward students of weakly lower priority if, for any $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$, either $x_{s}\;\succsim_{s}\;x_{s^{\prime}}$ or there exists some course $c^{*}$ such that $x_{s}\;\succsim_{s}\;(x_{s^{\prime}}\setminus\{c^{*}\})$.
This definition provides a criterion of fairness for students whose priorities are weakly ordered in the same way across all courses and minimizes envy between such students.
Allocations are found through mechanisms, which systematically elicit student preferences over course schedules. A mechanism is strategy-proof if there is no student, who, by reporting manipulated preferences, receives an allocation she strictly prefers to the course schedule she would get from reporting her true preferences. Evidence from business schools demonstrates that requiring strategic play on behalf of students can complicate efficiency (see Budish and Cantillon, 2012; Budish and Kessler, 2022; Sönmez and Ünver, 2010). We assume that strategic play is only a concern for students, as course priorities are set based on commonly observable factors such as student year of study or department.
Mechanisms that utilize market equilibria fail to be strategy-proof when students can influence course prices. A mechanism is strategy-proof in the large if truthful reporting is approximately optimal in an environment with a large number of students. For market equilibria, this amounts to being strategy-proof in a limit market in which each student regards the market prices as exogenous to her reported preferences (see Azevedo and Budish, 2019; Budish, 2011). [^7] He et al. (2018) considers the related notion of asymptotic incentive compatibility for random allocations. To avoid unnecessary heavy notation early in the paper, we introduce this concept formally in the proof of Theorem 5.
2 Pseudo-Markets with Priorities
In this section, we present our novel mechanism, which allocates courses to students by extending the concept of approximate competitive equilibrium from equal incomes from Budish (2011) to settings with course priorities. For this purpose, we assign each student $s\in{\cal S}$ a budget of fake money $b^{*}_{s}$, with $1\leq\min_{s}b^{*}_{s}\leq\max_{s}b_{s}^{*}\leq 1+\beta$ for some small $\beta>0$. The parameter $\beta$ is the maximum allowable budget inequality across students. We also allow for slack in the market-clearing condition, which is bounded by $\alpha\geq 0.$
Definition 4.
The allocation $x^{*},$ priority-specific prices $p^{*},$ and budgets $b^{*}$ constitute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities if:
  1. 1.
    Each student $s\in{\cal S}$ is assigned her most-preferred affordable course schedule:
    $x_{s}^{*}=\max_{\left(\succsim_{s}\right)}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}p_{c,r_{s,c}}^{*}x_{c}^{\prime}\leq b_{s}^{*}\right\}$
  2. 2.
    Each course $c\in{\cal C}$ has a cutoff priority level $r^{*}_{c}$ such that:
    $p^{*}_{c,r}\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}\\ {}[0,\overline{b})&r=r^{*}_{c}\\ [\overline{b},+\infty)&r<r^{*}_{c}\end{cases},$ ($\ast$)
    where $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ if $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ and $\overline{b}=1+\beta+\epsilon$ for some small $\epsilon>0$.
  3. 3.
    The market-clearing error is less than $\alpha$: $||z^{*}||_{2}\leq\alpha,$ where $z^{*}=(z_{1}^{*},....,z_{M}^{*})$ and:
    1. (a)
      $z_{c}^{*}=\sum_{s\in{\cal S}}x_{s,c}^{*}-q_{c}$ if $p_{c,1}^{*}>0,$
    2. (b)
      $z_{c}^{*}=\max(\sum_{s\in{\cal S}}x_{s,c}^{*}-q_{c},0)$ if $p_{c,1}^{*}=0$.
  4. 4.
    Student budgets are almost equal: $1\leq\min_{s\in{\cal S}}b_{s}^{*}\leq\max_{s\in{\cal S}}b_{s}^{*}\leq 1+\beta$.
Unlike Budish (2011), the above definition of an approximate competitive equilibrium allows for course prices to depend on priority levels, making the vector of prices $p^{*}=\left\{p^{*}_{c,r}\right\}_{c\in{\cal C},r\in\{1,...,R\}}\in\mathbb{R}^{M\!R}$. Condition ($\ast$2) ensures that each course has a cutoff level of priority, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face non-negative prices, and lower priority students cannot afford a seat. In doing so, condition ($\ast$2) guarantees no students are allocated a seat in a course when a higher priority student who would like a seat does not receive one. A similar condition appeared in He et al. (2018) in the context of random allocation with priorities.
Requiring that $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ if $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ ensures course $c$ has some students at the cutoff level demand seats, if there is any demand. If we allow $r^{*}_{c}$ to satisfy $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}=0$, raising the cutoff level of priority to $r^{*}_{c}+1$ does not change any student’s demand for the course. In turn, choosing $r^{*}_{c}$ so that $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}\}}x^{*}_{s,c}>0$ guarantees that the cutoff level of priority cannot be raised without removing a student from the course.
The market-clearing error for a course depends only on its price at $r=1$, the lowest priority level. Under-enrollment is only counted as an error if the price at level $r=1$ is positive. This requirement is a non-trivial extension of Budish (2011)’s definition to settings with course priorities. Alternatively, if a course’s market-clearing error depends on the price at the cutoff priority level, one can change the cutoff level and artificially lower the error. For example, consider an under-enrolled course with a price of zero at and above the cutoff priority level and a price of $\overline{b}$ below the cutoff. By this definition, the market-clearing error is zero. However, the unfilled course seats should be counted towards the market-clearing error, as they may be eliminated if the price below the cutoff level is decreased.
Budish (2011) establishes the existence of an approximate competitive equilibrium from equal incomes for any budget inequality $\beta>0$ with an upper bound on the market-clearing error of $\alpha=\sqrt{kM/2}$. [^8] Budish (2011) shows the existence of an approximate equilibrium from equal incomes with a market-clearing error at most $\sqrt{\min\{2k,M\}M}/2$. As we assume $k\leq M/2$, the upper bound reduces to $\sqrt{kM/2}$. This bound is proportional to the square root of the dimension of the price space. In the presence of $R$ priority levels for each of the $M$ courses, the dimension of the price space is $M\!R$. This could potentially require a large market-clearing error for a Pseudo-Market Equilibrium with Priorities to exist. The main idea of the proof of Theorem 1 is to reduce the effective dimension of the price space from $M\!R$ to $M$, where we can establish the existence of a Pseudo-Market Equilibrium with Priorities. In particular, for any vector $t\in\mathbb{R}^{M}$, we can define priority-specific prices $p\in\mathbb{R}^{M\!R}$ by setting:
$p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$
for each course $c\in{\cal C}$ and priority level $r\in\{1,...,R\}$. This parameterization guarantees that, for any $t\in[0,R\overline{b})$, the corresponding price vector $p$ satisfies condition ($\ast$2). We then prove the existence of an approximate equilibrium in the $M$-dimensional space.
Theorem 1 (Existence).
For any $\beta>0$, there exists a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities.
Theorem 1 establishes the existence of a Pseudo-Market Equilibrium with Priorities, extending the main result of Budish (2011) to environments with course priorities. This equilibrium has the same upper bound on the market-clearing error as the approximate competitive equilibrium from equal incomes in the same environment without priorities. The proof of Theorem 1 is presented in Appendix A.
2.1 The Pseudo-Market with Priorities Mechanism
Next, using the Pseudo-Market Equilibrium with Priorities, we introduce the Pseudo-Market with Priorities (PMP) mechanism and investigate approximate notions of stability, efficiency, fairness, and strategy-proofness it satisfies.
Definition 5.
The Pseudo-Market with Priorities mechanism with market-clearing error $\alpha$ and budget inequality $\beta$ is defined through the following steps:
  1. 1.
    Each student $s$ reports preferences $\succsim^{\prime}_{s}$ over course schedules.
  2. 2.
    Each student $s$ is assigned a random budget $b^{*}_{s}$ drawn uniformly from $[1,1+\beta]$.
  3. 3.
    Compute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities $(x^{*},p^{*},b^{*})$. Allocate courses to students according to $x^{*}$.
In general, the PMP mechanism depends on the level of allowable market-clearing error $\alpha$ and the level of budget inequality $\beta$. For results that hold for all non-negative $\alpha\geq 0$ and $\beta\geq 0$, we avoid this dependence in our exposition.
We first establish that the PMP mechanism results in an approximately stable allocation. Definitions 4 and 5 ensure that the allocation is individually rational. With courses regarded as being at their full capacity, condition ($\ast$2) on priority-specific prices prevents blocks. [^9] An alternative statement holds. Let $(x^{*},p^{*},b^{*})$ be an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities, and, for each $c\in{\cal C}$, let $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{sc}$ if $p^{*}_{c,1}>0$ and $q^{\prime}_{c}=\max(\sum_{s\in{\cal S}}x^{*}_{sc},q_{c})$ if $p^{*}_{c,1}=0$. Then, allocation $x^{*}$ is stable in the environment with adjusted capacities $q^{\prime}_{c},$ $c\in{\cal C}$.
Theorem 2 (Stability).
Every Pseudo-Market with Priorities mechanism results in an allocation that is approximately stable.
The proof is postponed to Appendix A. The main reason that the outcome of the PMP mechanism is not stable is the possibility of over- and under-enrolled courses. Stability requires a feasible allocation and only allows a course to have unfilled seats if no student can benefit from taking one. The PMP mechanism can violate these conditions when the market-clearing error is nonzero. However, as Theorem 1 shows, such instances are rare.
The presence of course priorities prevents the PMP mechanism from being Pareto efficient. This is similar to how stable allocations in matching markets can fail to be Pareto efficient (see Roth and Sotomayor, 1990). We illustrate this in the following example.
Example 1.
Consider a setting with two students ${\cal S}=\{1,2\}$ and two courses ${\cal C}=\{A,B\}$, with $q_{A}=q_{B}=k=1$. Student preferences are $\{A\}\succ_{1}\{B\}\succ_{1}\varnothing$ and $\{B\}\succ_{2}\{A\}\succ_{2}\varnothing$. Student budgets are $b^{*}_{1}=1$ and $b^{*}_{2}=1+\beta$ for some $10<\beta<$ 1. Course priorities are the opposite of student preferences, with $r_{1,B}=r_{2,A}=2$ and $r_{1,A}=r_{2,B}=1$. Budgets $b^{*}$, cutoff priority levels $r^{*}_{A}=r^{*}_{B}=2$, priority-specific prices $p^{*}_{A}=p^{*}_{B}=(2,1)$, and allocation:
Student $x^{*}_{s,A}$ $x^{*}_{s,B}$
1 0 1
2 1 0
constitute a $(0,\beta)$-Pseudo-Market Equilibrium with Priorities. The allocation is not Pareto efficient, as both students would be made better off by exchanging their assigned seats.
The next result shows that the outcome of the PMP mechanism satisfies a notion of constrained efficiency. Adjusting course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$, no allocation that Pareto dominates the outcome of the PMP mechanism respects priorities to the same degree. For each course, the distribution of assigned student priority levels cannot first-order stochastically dominate the corresponding distribution for the PMP outcome (see Definition 2). [^10] Similarly to Theorem 2, an alternative statement holds. Let $(x^{*},p^{*},b^{*})$ be an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities, and, for each $c\in{\cal C}$, let $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{sc}$ if $p^{*}_{c,1}>0$ and $q^{\prime}_{c}=\max(\sum_{s\in{\cal S}}x^{*}_{sc},q_{c})$ if $p^{*}_{c,1}=0$. Then, allocation $x^{*}$ is priority-constrained efficient in the environment with adjusted capacities $q^{\prime}_{c},$ $c\in{\cal C}$. In environments without course priorities, this result reduces to Proposition 2 in Budish (2011).
Theorem 3 (Efficiency).
Every Pseudo-Market with Priorities mechanism results in an allocation that is approximately priority-constrained efficient.
The proof of the above result resembles the proof of Schlegel and Mamageishvili (2020) for single-unit settings with random allocations and is postponed to Appendix A. [^11] We assume each agent has strict preferences over course schedules. So, unlike Schlegel and Mamageishvili (2020) and Miralles and Pycia (2021), we do not require that a student chooses the cheapest course schedule when multiple course schedules are optimal.
Though the concepts of approximate stability and priority-constrained efficiency are related, neither implies the other. If we assume in Example 1 that both students have the same priority in each course, then there are two (approximately) stable allocations: $x_{1}=\{A\}$, $x_{2}=\{B\}$ and $x_{1}=\{B\}$, $x_{2}=\{A\}$. However, only the first allocation is approximately priority-constrained efficient. Here, approximate priority-constrained efficiency selects among approximately stable allocations, favoring more Pareto optimal outcomes. In turn, suppose that student $1$ is the only student in the environment and that $x_{1}=\{A,B\}$. This allocation is approximately priority-constrained efficient, but not approximately stable, as it is not individually rational.
Theorem 2 prevents the possibility that a student can benefit from taking a seat in a course that a student with lower priority obtains. However, priorities in undergraduate course allocation are weak, with hundreds of students often at the same level of priority across many courses. The next result addresses the question of fairness among students whose priorities are weakly ordered in terms of their priorities across all courses.
Theorem 4 (Fairness).
If $\beta\leq\frac{1}{k-1}$, the Pseudo-Market with Priorities mechanism results in an allocation that has schedule envy bounded by a single course toward students of weakly lower levels of priority.
The result of Theorem 4 extends the one established in Budish (2011) for settings where all students are on the same priority level. With priorities based on seniority and student department, the above result implies that the PMP mechanism with a small budget inequality produces an allocation in which any two students from the same department and the same year of study have schedule envy bounded by a single course toward one another. [^12] This result can be strengthened. The PMP mechanism results in an allocation $x^{*}$ such that for any two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for each $c\in x^{*}_{s^{\prime}}$, envy is bounded by a single course. In particular, rather than all courses, $s$ needs only to be at a weakly higher priority level in the courses being taken by $s^{\prime}$.
Our final theorem establishes that students have almost no incentive to manipulate the PMP mechanism in large populations. This is commonly the case at undergraduate institutions, where student bodies often consist of thousands of students.
Theorem 5 (Strategy-Proofness).
The Pseudo-Market with Priorities mechanism is strategy-proof in the large.
We define the concept of strategy-proofness in the large in the proof of Theorem 5 in Appendix A. To prove this theorem, we leverage on a result from Azevedo and Budish (2019), which establishes a connection between strategy-proofness in the large and envy-freeness. Students can be partitioned into groups based on their course priorities, where each student in a group is at the same level of priority in every course. The PMP mechanism is a semi-anonymous mechanism, as, within groups, each student faces the same course prices and the same random lottery over budgets. As a student cannot prefer the allocation of any lower-budget student in the same group, the PMP mechanism is envy-free but for tie-breaking. Our result then follows from Azevedo and Budish (2019), which shows that any semi-anonymous mechanism that is envy-free but for tie-breaking is strategy-proof in the large. [^13] One can also consider a continuum replication of a course allocation problem with course priorities. By leveraging on the price structure as in Theorem 1, the steps of Theorem 4 of Budish (2011) can be adapted to obtain the result. The alternative proof is available upon request.
Last, we relate the PMP mechanism to some special cases discussed in previous literature. In the case of no priorities, our mechanism reduces to the approximate competitive equilibrium from equal incomes mechanism of Budish (2011). On the other hand, courses may have strict priorities over students. In this case, we have a two-sided matching problem. As students can view sets of courses as complements, no stable matching is guaranteed to exist. Still, Theorem 1 confirms a Pseudo-Market Equilibrium with Priorities exists with a small error. So, our mechanism does not necessarily coincide with any stable matching mechanism.
If students have strict preferences over individual courses as well, the student-proposing deferred acceptance mechanism results in the student-optimal stable matching (see Roth, 1984), which can be supported with a Pseudo-Market Equilibrium with Priorities where the market clears exactly, the cutoff level of priority is at the lowest-priority student to receive a seat, and the price of each course is zero at the cutoff. The set of equilibria with zero market-clearing error exactly coincides with the set of stable matchings. [^14] See Miralles and Pycia (2021) for a related result showing that every efficient assignment can be decentralized through prices in random allocation settings. In large markets, we expect the set of stable matchings to be small (see Kojima and Pathak, 2009) and, hence, PMP outcomes should not differ much from each other.
When each student can only enroll in one course, our setting reduces to the unit demand variation of the problem, known in the literature as school choice. The standard mechanisms in school choice are ordinal (Abdulkadiroğlu and Sönmez, 2003) or elicit only restricted information about cardinal preferences from agents (Abdulkadiroğlu, Che and Yasuda, 2015). With weak course priorities, the deferred acceptance algorithm may not return a student-optimal stable matching (Erdil and Ergin, 2008). At the same time, any stable matching that is priority-constrained efficient, student-optimal or not, can be supported by a Pseudo-Market Equilibrium with Priorities with zero market-clearing error.
Finally, He et al. (2018) show how to extend the pseudo-market approach of Hylland and Zeckhauser (1979) to school choice settings with random allocations. Their pricing structure ensures that their pseudo-market equilibrium allocation is fair. We adopt a similar pricing structure but analyze deterministic allocations rather than random allocations. While deterministic allocations are preferable from a practical market design point of view, they are associated with additional complications in the existence of an equilibrium.
3 Simulations
Next, we analyze the performance of the Pseudo-Market with Priorities mechanism with student preferences estimated using course allocation data from a private institution. We compare the Pseudo-Market with Priorities mechanism with a version of the Random Serial Dictatorship with course reserves, the mechanism used for course allocation at many U.S. universities, and Deferred Acceptance with single and multiple tie-breakings, mechanisms actively used in the allocation of students to public schools.
3.1 Student and Course Data
Table 1: The number of students and courses in each college.
College A B C D E F G
# of students 853 1642 259 1274 745 741 509
# of courses 180 84 12 269 88 84 39
Notes: We use capital letters to denote college names.
We utilize university data containing the final allocation of courses to students from the Spring 2018 semester at a private institution in the mid-Atlantic region. There are 6 02360236\,0236 023 students divided into four years of study and $756$ courses. The university is made up of seven colleges. Colleges are the largest constituent units of the university and typically contain several departments. For example, the Department of Economics and the Department of English are housed in the College of Liberal Arts at many U.S. universities. [^15] We only consider undergraduate students. We merge fifth-year and fourth-year students, as there are few fifth-year students in the data, and both groups are treated equally in the course registration process. Some graduate and exchange students take undergraduate courses, but we exclude these students. Colleges differ in the size of their student population and the number and type of courses they offer, as shown in Table 1. Some colleges serve the whole university, offering general education courses in math and science, whereas others mainly offer courses only for their students.
Course enrollments. Table 2 shows that $11.2\%$ of all courses are at or above their listed capacity. This percentage serves as an important indicator of overlap in student interest and the need for a proper mechanism to allocate seats in over-demanded courses. Moreover, the table indicates that the university allows a significant number of courses to be over-enrolled (around $7.3\%$). The listed maximum capacity for such courses is intended to ensure proper class dynamics and balanced course sections. For example, teaching business communication or drama courses of a large size is difficult, but these courses can often accommodate a few students above their listed capacity. This fits our treatment of capacity constraints in the Pseudo-Market with Priorities mechanism, which allows enrollment above a course’s capacity. However, in some courses, the maximum capacity is strictly binding. These situations are typically associated with the physical limitations of the classrooms. We set each course’s capacity equal to the maximum of the capacity observed in the data and the true enrollment.
Course seat reservations. The data records course seat reservations, each specifying a course and a number of seats reserved for students in certain departments and years of study. The concept of course seat reservations is somewhat different from the concept of a priority structure used in the theoretical part of the paper. Rather than saving a fixed number of seats for specific groups of students, a priority structure ensures that certain student groups have a claim on course seats over other student groups. Table 3 presents quantiles of the distributions of course capacities, enrollment, and total number of reserved seats from the university data. In total, 33 4553345533\,45533 455 course seats were available across courses, 23 3692336923\,36923 369 seats were occupied by students, and 13 9221392213\,92213 922 were reserved at the start of the course registration process.
Table 2: The percentage of courses at or above maximum capacity.
Seats taken above capacity $\geq 0$ $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses 11.2 7.3 4.1 3.3 2.5 1.9
Notes: The table reports information about over-enrollment of courses in the university data. Maximum capacity is a soft constraint on enrollment for many courses.
Table 3: Course capacities, enrollment, and reservations.
Quantile 0.10 0.25 0.50 0.75 0.90
Maximum Capacity 8 15 25 50 98
Actual Enrollment 3 7 15 35 72
# of Reserved Seats 0 0 3 20 53
Notes: This table presents some quantiles of the distributions of course capacities, enrollment, and number of reserved seats from the university data. Courses with zero enrollment are dropped from the data.
Table 4 provides some examples of course seat reservations. For instance, Microeconomics reserves $25$ seats for Department $1$ and Department $2$ students in Year $1$. In the data, it is possible that these reservations are done independently by the two departments (e.g., among Year $1$ students, $15$ seats are reserved for Department $1$ students and $10$ seats reserved for Department $2$). Whenever this is the case, we combine these reservations into one by uniting the department names and summing the number of reserved seats. After this simplification, each course has at most one reservation for each year of study and one for “all years of study”.
Table 4: Examples of course seat reservations.
Course Departments Year of Study Total Reserved Seats
Microeconomics Dept 1, Dept 2 Year 1 25
Real Analysis Dept 2 Year 3, Year 4 50
Machine Learning All Year 4 10
Communication Dept 3 All 5
Notes: Each row of this table corresponds to a seat reservation for a different course. “All” denotes that any student, regardless of department/year of study, is eligible for a reserved seat.
Course priorities. Using the data on course seat reservations, we create a course priority structure that reflects the university administration’s intentions. [^16] There is a tradeoff between using a course priority structure and using course seat reservations. Reservations allow the university administration to specify distributional concerns by reserving a certain amount of a course’s seats to multiple, different groups of students without prioritizing one of the other. However, they require specifying an exact number of reserved seats, which is inflexible in environments with uncertain course demand. With priorities, if there are few high priority students, the seats are free to be allocated to students of lower priority. However, priorities do not address distributional concerns. See Celebi and Flynn (2024) for an innovative analysis merging priorities and reserves into the concept of adaptive priorities. In practice, the course registration process closely follows a Random Serial Dictatorship with course reserves, where students select courses in order of seniority (with ties broken randomly) and courses make seat reservations based on department and year of study. Hence, the process favors students in later years of study by giving them earlier time slots to choose courses, resembling year-specific priorities (see Section 1). In addition, courses reserve seats for students in specific departments and years of study, resembling department-specific priorities. [^17] Universities that follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. However, each university operates slightly differently. Within a year of study, many universities (e.g., Duke) randomly stagger students across time slots rather than having all register at once. In addition, some universities (e.g., Vanderbilt) reserve seats for students in specific majors or degree programs rather than departments or years of study. We combine these notions by using a hybrid priority structure in our simulations, with year-specific priorities taking precedence over department-specific priorities. We analyze the case in which department-specific priorities take precedence over year-specific priorities in Appendix D. [^18] There are more than two ways how the year-specific and department-specific priorities can be combined.
Example 2 provides an illustration for Real Analysis in Table 4. We drop the exact number of reserved seats and assign the highest priority to 4th-year students in Department 2, followed by 4th-year students in other departments, 3rd-year students in Department 2, and 3rd-year students in other departments. No seats are reserved for 1st- and 2nd-year students, making all 2nd-year and all 1st-year students the bottom two levels.
Example 2.
The course seat reservations for Real Analysis in Table 4 results in the following hybrid priority structure for the course:
  • $r_{s,c}=8$: 4th-year students in Dept 2
  • $r_{s,c}=7$: 4th-year students in all departments except Dept 2
  • $r_{s,c}=6$: 3rd-year students in Dept 2
  • $r_{s,c}=5$: 3rd-year students in all departments except Dept 2
  • $r_{s,c}=4$: $\varnothing$
  • $r_{s,c}=3$: 2nd-year students in all departments
  • $r_{s,c}=2$: $\varnothing$
  • $r_{s,c}=1$: 1st-year students in all departments
Student utilities. The university data does not contain information on students’ preferences over course schedules. To recover this information, we use a parametric form for student utilities and the method of simulated moments to calibrate its parameters (see, e.g., Agarwal, 2015). We assume that each student can take up to $k=5$ courses and that students’ preferences are additive over courses. [^19] Some students are enrolled in six or even seven courses in the data. These students account for less than $3\%$ of the total student population. Student $s$’s utility from taking course $c$ is denoted by $u_{sc}$ and has the following form:
$u_{sc}={\color[rgb]{0,0,0}Y^{\prime}_{s}\theta}+Y^{\prime}_{s}\Gamma H_{c}+{\color[rgb]{0,0,0}z_{c}}+\varepsilon_{sc},$ (1)
where $Y_{s}$ and $H_{c}$ are binary vectors of observed student and course characteristics, respectively, $\theta$ is a vector of unobserved coefficients, $\Gamma$ is a matrix of unobserved coefficients conformable with vectors $Y_{s}$ and $H_{c}$, $z_{c}$ is an unobserved course-specific component, and $\varepsilon_{sc}$ is random utility component that is assumed to have a normal distribution $\varepsilon_{sc}\sim N(0,\sigma^{2})$ with unobserved parameter $\sigma$. As normalizations, the value of the outside option from not taking a course is $u_{s0}=0$ and the standard deviation of the noise parameter is $\sigma=1$.
The parametric form in equation (1) contains three non-random components. The first, $Y^{\prime}_{s}\theta$, is a horizontal component determined by the student’s characteristics. The sole nonzero element in $Y_{s}$ indicates student $s$’s college $a\in{\cal A}=\{A,...,G\}$ and student $s$’s year of study $y\in{\cal Y}=\{1,...,4\}$. That is, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ and $Y_{s}^{\prime}\theta=\theta_{ay}$, where vector $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $|{\cal A}||{\cal Y}|=28$ elements to calibrate. The larger this horizontal component is, the more courses student $s$ wants to enroll in. This accounts for variability in the number of courses taken by students from different colleges and years of study (see Table C.1 in Appendix C).
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. This accounts for large heterogeneity in patterns of course enrollment across colleges (see Table C.2 in Appendix C). The third component, ${\color[rgb]{0,0,0}z_{c}}$, is a vertical component that measures the popularity or quality of course $c$. The vector $z$ has length $756$, with one entry for each course.
We cannot jointly identify the parameters $\theta$, $\Gamma$, and $z$, as they enter in equation (1) in an additive way. We normalize the average horizontal component across students in the same college to $\overline{\theta}_{a}=0$ and set the interaction terms $\gamma_{aa}$ to 00 for each $a\in{\cal A}$. These normalizations allow for an interpretation of $z_{c}$ as the average student utility from taking course $c$ among students in the same college as course $c$. The value of the parameter $\gamma_{aa^{\prime}},a\neq a^{\prime}$ is the average increase in student utility from taking course $c^{\prime}$ in college $a^{\prime}$ over taking course $c$ in student’s own college $a$ when both courses have the same values of vertical components (i.e., $z_{c}=z_{c^{\prime}}$). We provide more details on the parameter normalization on p. C.1 in the Appendix C. Overall, we have $819$ independent parameters $\theta$, $\gamma$, and $z$ to calibrate.
Last, we assume that each student’s choice set is limited to $80$ courses. This assumption is motivated by the observation that any given student typically only considers taking a subset of all offered courses in a given term and is made for computational purposes (see also Budish and Cantillon, 2012; Diebold and Bichler, 2017). We randomly draw these $80$ courses among those taken by at least one student from the same college and year. The probability that a student in college $a$ and year $y$ draws course $c$ is proportional to the share of seats taken in course $c$ by students in college $a$ and year $y$, among all seats taken by such students. [^20] We expand some students’ choice sets to make sure that courses have enough reserve-eligible students with the course in their choice sets. This ensures that our calibration process runs smoothly. We elaborate on this on p. C.1 in Appendix C. Not all $80$ utility entries are necessarily positive.
We use the method of simulated moments to identify the model’s parameters. A description of the calibration procedure, interpretations of the calibrated parameter values, and model fit results are left to Appendix C. Here, we discuss two major complications that arose during the calibration process. The first obstacle in recovering student utilities was that, in practice, course reservations do not remain constant during the registration process. After the first week of course registration, student demand for courses becomes more or less clear. To ensure complete course enrollment, departments start relaxing course reservations by admitting students from waiting lists. While there are best practices for handling waiting lists, this phase of enrollment is done solely at each department’s discretion. [^21] The best practices include prioritizing more senior students and students in the course’s department. As a result, a course’s final enrollment can violate its seat reservations set at the beginning of the process. We use Hall’s Marriage Theorem to identify course reserve violations and minimally adjust the number of reserved seats. The seat reservations are lowered in 170 of the 756 courses to ensure the number of reserved seats is consistent with the allocation observed in the data.
The second obstacle concerned calibrating the vertical components $z_{c}$ for courses at maximum capacity. For such courses, no value of $z_{c}$ can lead to a higher enrollment than we observe in the data. While we could have used modern econometrics techniques to partially identify course fixed effects and provide inference on parameters using moment inequalities (see Chernozhukov, Chetverikov and Kato, 2019), it is not clear which parameter values from the partially identified set to choose for the student utility model. Instead, we identified the vertical components for courses at maximum capacity using data on how quickly the seats in those courses were occupied during the course enrollment process. Appendices C.2-C.3 provide a detailed discussion of both obstacles.
3.2 Course Allocation Mechanisms
Using the student utility model, we simulate the outcomes of four mechanisms across 100 random draws of students’ utilities. We assume that students report their utilities truthfully, as all considered mechanisms are either strategy-proof or strategy-proof in the large.
The Pseudo-Market with Priorities (PMP) mechanism. Students are randomly assigned evenly spaced budgets between $1$ and $1+\beta$, with $\beta=\frac{1}{k-1}=0.25$. Students are placed at one of $R=8$ levels of priority in each course. Odd levels 1, 3, 5, and 7 correspond to first, second, third, and fourth-year students ineligible for a course’s reserves, respectively. Even levels correspond to students eligible for reserved seats (see Example 2). Some courses have no students at some levels of priority. For example, if a course has no seat reservations, there are effectively only four levels, one for each year of study.
Our algorithm that searches for a Pseudo-Market Equilibrium with Priorities departs from the one used in Budish (2011) to accommodate priority-specific prices and obtain a tighter bound on the number of over-enrolled courses. As in the proof of Theorem 1, we parameterize prices with $t\in[0,R\overline{b}]^{M}$, where $\overline{b}=1.251>1+\beta$. For each $t\in[0,R\overline{b}]^{M}$, course $c\in{\cal C}$, and priority $r\in{\{1,...,R\}}$, we define priority-specific prices as $p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$. Then, we look for an equilibrium in this lower-dimensional space.
The algorithm has two phases. Phase I starts with an educated guess for the equilibrium prices based on students’ utility-maximizing course schedules. Then, the algorithm finds each student’s utility-maximizing schedule subject to course prices and adjusts prices proportional to the number of over and under-enrolled seats. The algorithm searches in the lower-dimensional space of $t$-parameters for student demand with a market-clearing error smaller than the theoretical bound $\alpha=\sqrt{kM/2}\approx 43.5$. If the algorithm fails to improve in a given iteration, the current price vector is retained and the size of the price adjustment is reduced for the next iteration. We move to Phase II if the market-clearing error fails to improve within a certain number of iterations, or if the improvement is smaller than $1\%$ when the theoretical bound has been reached. The main innovation of the search process is that course demand is only updated for the subset of students whose demand could be potentially affected by price change on each step. This innovation has proven particularly effective for large student populations, significantly reducing the time required for the search process. [^22] Starting with an educated guess also improves the speed of convergence. In addition, our price adjustment process is a version of the Walras’ tâtonnement procedure, a simpler algorithm than the one used in Othman, Sandholm and Budish (2010). At the same time, we do not consider course complementarities, which is an important part of their paper.
Phase II takes the course prices obtained in Phase I and works to reduce the excess course demand. The phase ensures that no courses are over-enrolled by more than $k-1=4$ seats and that the distribution of over-enrolled courses is smaller than the one observed in the data (see Table 2). [^23] Doing so guarantees that the allocation both has an aggregate market-clearing error below the bound in Theorem 1 and meets the good-by-good bound of Nguyen and Vohra (2022). As a result of Phase II, an allocation with a market-clearing error greater than the theoretical bound $\alpha$ might be obtained. In this case, the algorithm repeats Phases I and II. If there are more than $6$ iterations of Phases I and II, the algorithm restarts with a new set of initial prices. Appendix E.2 provides more details about Phases I and II.
The Deferred Acceptance mechanism with single and multiple tie-breakings. These two mechanisms use the student-proposing deferred acceptance algorithm to allocate course seats to students. The only difference between the two variants is how they break ties between students at the same level of priority in a course. The first variant uses a single tie-breaking order across all courses. To break ties, we use the budgets assigned in the PMP mechanism, placing students with the largest budgets first. We refer to this version of the mechanism as DA-STB. The second variant uses multiple, course-specific tie-breakings. We refer to this version of the mechanism as DA-MTB. Both versions have received a thorough analysis in the many-to-one matching literature (see, e.g., Erdil and Ergin, 2008, 2017). [^24] As students can take more than one course, neither DA-STB nor DA-MTB is strategy-proof (see Roth, 1982). However, Azevedo and Budish (2019) establish that both mechanisms are strategy-proof in the large.
The Random Serial Dictatorship (RSD) with optimal course reserves. Students are ordered by seniority. To break ties, we use the same order as in PMP and DA-STB. Using this order, the mechanism assigns each student the $k$ courses available that offer her the highest utility (or less than $k$, if there are not $k$ such courses). Course seat reservations are treated as described in Section 3.1: for each course and each year of study (including the “All” option), there is a set of departments and a number of reserved seats, which could be zero. If a student is eligible for one of a course’s seat reservations, she can be enrolled in a reserved or a regular seat. Otherwise, the student can only be enrolled in a regular seat. All reserved seats must be occupied before a reserve-eligible student occupies a regular seat.
Using the number of reserved seats observed in the data leads to significant under-performance of RSD with course reserves (see Appendix D.2). The RSD mechanism treats this number of reserves as fixed, resulting in numerous inefficiencies. In practice, the process is more flexible. University departments relax course reservation constraints during the course allocation process. Doing so releases pressure on the department heads to set course reservations precisely at the beginning of the process. In turn, the PMP, DA-STB, and DA-MTB mechanisms compute an effective number of reserved seats in the course allocation process. Hence, comparing the performance of the other three mechanisms to the performance of RSD with a fixed number of reserved seats would not be a fair exercise.
Instead, we estimate the optimal reserved seats for each course in the RSD mechanism. For this purpose, we generate $100$ environments with random student utilities according to the student utility model and a random tie-breaking and run the DA-STB mechanism. To determine the optimal reserved seats for a given course and set of departments, we take the average across all environments (rounded to the nearest integer) of the number of reserve-eligible students assigned to the course. We call the result of the RSD mechanism operated with these calculated reserved seats the RSD with optimal course reserves.
3.3 Simulation results
Table 5: Each student’s preferred mechanism by year of study.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
[…]
Year 1 12.9 (1.9) 13.7 (1.7) 3.8 (1.0) 6.7 (1.2) 16.1 (1.2) 18.8 (1.4)
Year 2 24.1 (2.1) 16.7 (1.4) 8.8 (1.2) 7.5 (1.0) 26.4 (1.2) 25.5 (1.1)
Year 3 21.3 (1.8) 12.4 (1.2) 8.5 (1.0) 3.8 (0.7) 22.5 (1.1) 19.9 (1.2)
Year 4 12.9 (1.2) 7.9 (0.9) 3.9 (0.5) 1.6 (0.5) 16.4 (0.9) 14.8 (0.9)
Notes: This table reports the percentage of students who strictly prefer a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who strictly prefer the benchmark. Results are averages across $100$ runs with different random component draws.
We evaluate the performance of the Pseudo-Market with Priorities mechanism, Deferred Acceptance with single and multiple tie-breakings mechanisms, and Random Serial Dictatorship mechanism with optimal course reserves across 100 simulation runs. [^25] The program was written in Mathematica. The majority of the simulations were run on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GBs of RAM, 70 GBs of SSD hard drive space, and Windows Server 2019 Standard operating system. On average, it takes about forty-five minutes to calculate the allocation outcomes of all four mechanisms at one core. Each run corresponds to a random draw of students’ utilities and a random tie-breaking order, which is kept the same across PMP, RSD, and DA-STB. DA-MTB has multiple course-specific tie-breakings.
Student satisfaction. We consider the Random Serial Dictatorship mechanism with optimal course reserves as a benchmark. Table 5 reports the percentage of students who strictly prefer a given mechanism to the benchmark and the percentage of students who strictly prefer the benchmark. The remaining share of students is attributed to indifferences, which account for $48.1\%-94.5\%$ of students across mechanisms and years of study. The large share of indifferences can be explained by students taking under-enrolled courses, as only $11.2\%$ of courses are at or above full capacity in the data.
For second-, third-, and fourth-year students, the share of students who strictly prefer the PMP mechanism to the RSD benchmark is larger than the share of students with the opposite preference. Among students with changing schedules, the mean improvement in utility is $5.1\%$, $8.0\%$, and $9.4\%$, respectively. While slightly more first-year students prefer the benchmark, the PMP mechanism provides an $0.8\%$ mean improvement in utility. This weaker result can be attributed to the gains for more senior students. On average, the PMP mechanism assigns $135$ more course seats to second- through fourth-year students than the RSD benchmark. The DA-STB and DA-MTB mechanisms also perform well relative to the benchmark for second- through fourth-year students. For each DA mechanism, the share of students who strictly prefer it to the RSD benchmark is larger than the share of students with the opposite preference. The difference in these shares is smaller than that for the PMP mechanism for students in each year of study. In general, the efficiency gains for the PMP mechanism are likely understated, as we do not consider any complementarities or more complex student preferences, which the PMP mechanism can easily accommodate.
Table 6: Over-enrolled courses in the Pseudo-Market with Priorities mechanism.
# seats above max capacity $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses $2.5$ $0.4$ $0.1$ $0.02$ $0.0$
Notes: This table provides information about the distribution of over-enrolled courses in the Pseudo-Market with Priorities mechanism. Results are averages across 100 runs with different random component draws.
In principle, the PMP mechanism may be delivering better outcomes to students due to the market-clearing error, which allows some courses to be over-enrolled. [^26] Alternatively, we could have compared the performance of the four mechanisms with quotas equal to the number of seats assigned in the PMP mechanism for each run. We thank Olivier Tercieux for this suggestion. Table 6 reports the distribution of over-enrollment in courses. In line with the good-by-good clearing bound of Nguyen and Vohra (2022), no courses are over-enrolled by more than $k-1=4$ seats. In addition, the percent of over-enrolled courses is smaller than that observed in the university’s final allocation. Less than $2.5\%$ of courses are over-enrolled by more than one seat, as opposed to $7.3\%$ in the data. The average market-clearing error across outcomes is about $21.4$ seats, less than half of the worst-case bound $\alpha\approx 43.5$.
Table 7: The change in the standard deviation of students’ utilities by year of study.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
Year 1 -3.3% -0.6% -2.4%
Year 2 -5.6% -1.5% -4.0%
Year 3 -3.4% -0.2% -1.9%
Year 4 -1.1% 0.2% -0.7%
Notes: This table reports the percent change in the standard deviation of students’ utilities compared to the benchmark of to the Random Serial Dictatorship with optimal course reserves for three mechanisms. Results are averages across $100$ runs with different random component draws.
Allocation fairness. We compare allocation fairness across mechanisms using two metrics. Table 7 shows that, within each year of study, the PMP and DA-MTB mechanisms decrease the standard deviation of students’ utilities relative to the RSD benchmark. The same result holds for first- through third-year student in the DA-STB mechanism. For each year of study, this finding is the strongest for the PMP mechanism. Between the two DA mechanisms, the DA-MTB mechanism performs slightly better. For each mechanism, the largest reduction occurs for second-year students. This is in line with Table 5, where Year 2 had the fewest percent of students who are indifferent between each mechanism and the benchmark.
Table 8 reports results on students who experience schedule envy towards students of weakly lower priority in each course. The first column shows the percentage of students with no such envy. The values in each subsequent column are obtained by removing one course from the envied schedules and checking whether the student still experiences envy. In line with the result of Theorem 4, in the PMP mechanism, no students experience schedule envy greater than a single course toward students of weakly lower priority in each course. None of the other three mechanisms satisfy this property. About $9.1\%$ of students experience envy in the PMP mechanism, substantially less than in the RSD, DA-STB, and DA-MTB mechanisms ($17.2\%$, $14.1\%$, and $13.3\%$, respectively). [^27] Ours is the first paper to compare DA-STB and DA-MTB in terms of the envy students experience to others of a weakly lower priority. See Abdulkadiroğlu, Pathak and Roth (2009) and Ashlagi, Nikzad and Romm (2019) for a comparison of these mechanisms using other measures of fairness and efficiency. Relative to the RSD benchmark, the PMP mechanism reduces the percentage of students who experience any envy toward students of weakly lower priority by 8 points, or almost $500$ students.
Table 8: The percentage of students who experience schedule envy.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
Pseudo-Market with Priorities 90.9 (0.6) 9.1 (0.6) 0 (0.0) 0 (0.0) 0 (0.0) 0 (0.0)
Random Serial Dictatorship with optimal course reserves 82.8 (1.2) 13.9 (0.9) 2.6 (0.3) 0.6 (0.2) 0.06 (0.05) 0.002 (0.0)
Deferred Acceptance with single tie-breaking 85.9 (0.6) 11.6 (0.5) 2.1 (0.2) 0.4 (0.2) 0.03 (0.03) 0.001 (0.0)
Deferred Acceptance with multiple tie-breakings 86.7 (0.6) 12.6 (0.6) 0.7 (0.2) 0.02 (0.02) 0 (0.0) 0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the four mechanisms. Results are averages across 100 runs with different random component draws. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by $1,...,5$ courses.
We can also study allocation fairness between priority levels. In the PMP, DA-STB, and DA-MTB mechanisms, no student can ever improve her course schedule by taking a seat that is occupied by a student of strictly lower priority. However, in the RSD mechanism, $16.2\%$ of students can. These instances occur when a course reserves too few of its seats for a designated group of students. Then, a reserve-eligible student late in the tie-breaking order may miss out on a seat when a reserve-ineligible student obtains one.
Limitations. Our student utility model has limitations that a richer data set and a more detailed empirical analysis could address. Most notably, the additive utility specification may be too restrictive, as it fails to consider some complementarities and substitutabilities across courses. For example, courses covering similar content may be substitutes, while a course requiring another as a corequisite (i.e., to be taken at the same time) may be complements. Still, the additive structure of student preferences allows us to compare the performance of our new mechanism with the DA-STB and DA-MTB mechanisms. Unlike these mechanisms, the PMP mechanism can easily be extended to more general preference specifications.
A first step in expanding this analysis beyond additive utilities is to consider preferences over course sections rather than courses. Universities often offer the same course in different time slots or with different instructors. A section-level analysis would relax the assumption of additive utility, as each student cannot be enrolled in two sections of the same course. In the data, many courses have several sections, with some being much more popular than others. Since we aggregate enrollment and capacities across sections, if one of a course’s sections is not at maximum capacity, the course cannot be either. Overall, $30\%$ of course sections are at maximum capacity, in comparison to only $11.2\%$ of courses. Consequently, we expect the gains from the PMP mechanism presented earlier in this section to be understated.
Another approach would be to maintain the additive utility specification, but consider more factors that influence student preferences. Our student utility model assumes that a student’s preferences depend only on her college and year of study. Allowing utilities to depend on a student’s department would capture additional heterogeneity in preferences. Additionally, information about past enrollment would help to disentangle preferences over courses from constraints based on the course’s prerequisites. This information would also allow us to pin down students’ choice sets. However, past enrollment, along with other characteristics (e.g., demographics) are not available in the data.
We conclude this section by summarizing the main insights from our empirical analysis. The Pseudo-Market with Priorities mechanism delivers a substantially fairer outcome than the Random Serial Dictatorship with optimal course reserves, both in terms of the standard deviation of student utilities and the percentage of students with envy toward other students of weakly lower priority. Furthermore, student satisfaction improves for second- through fourth-year students. The two variants of the DA mechanism share some benefits of the PMP mechanism, but have weaker fairness results and cannot be extended to accommodate more general student preferences. Appendix D presents additional simulation results, detailing the equilibrium prices in the PMP mechanism (Section D.1), the performance of six alternative mechanisms (Section D.2), results under an alternative priority structure (Section D.3), and robustness checks with respect to the size of student choice sets and the standard deviation of the noise parameter (Section D.4).
4 Discussion
This paper explores the many-to-many matching problem that arises in undergraduate course allocation. We allow courses to prioritize students based on factors such as year of study and department and design a deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that respects course priorities. The PMP mechanism extends the approximate competitive equilibrium from equal incomes mechanism of Budish (2011) to settings with course priorities and has the same, small worst-case bound on the market-clearing error. It also satisfies desirable approximate notions of stability, efficiency, fairness, and strategy-proofness. Our empirical findings demonstrate that the PMP mechanism outperforms the Random Serial Dictatorship with optimal course reserves, a mechanism resembling the one widely used in practice, in terms of student satisfaction and allocation fairness.
The benefits of introducing the PMP mechanism depend on university-specific factors. Universities with only a few popular courses will likely see small gains from implementing any fairer mechanism. In contrast, when many courses are overdemanded, the mechanisms used in practice often lead to significant envy among students at the same level of priority. The PMP mechanism will be most effective in these environments, as it guarantees fairness without compromising priority constraints or student satisfaction.
A university’s course priority structure also plays a considerable role in its potential gains from implementing the PMP mechanism. With priorities based on seniority and student department, the random serial dictatorship and deferred acceptance mechanisms heavily rely on tie-breaking to allocate courses, resulting in a large percentage of students with envy toward others in the same year of study and department. Our Theorem 4 guarantees that, with a small enough budget inequality, the PMP mechanism bounds envy among these students by a single course. In contrast, if schools employ finer priority structures, using factors such as units completed or grade point average, the PMP and RSD mechanisms are likely to result in similar course allocation outcomes.
We conclude by discussing several aspects pertinent to the practical implementation of the PMP mechanism, drawing on previous experience with implementing the A-CEEI mechanism at Wharton and other business schools. The model in Section 1 assumes that students can report their preferences over all possible course schedules. In practice, with thousands of available courses, reporting preferences requires a language simple enough for students to use and expressive enough to elicit students’ actual preferences. The practical implementation of the A-CEEI mechanism in Budish et al. (2017) allows students to report cardinal values for individual courses and adjustments for pairs of courses. However, Budish and Kessler (2022) find evidence of preference reporting errors that negatively impacted student welfare at Wharton. Moreover, some adjustments to the preference reporting language of Budish et al. (2017) may be necessary to make this framework accessible to undergraduate students. For example, as credit limits play a large role in determining course schedules, students need to be able to indicate a maximum number of courses or credits to enroll in while reporting their preferences. Importantly, the deployment of recent machine learning-based techniques should tremendously improve the precision of elicited preferences (see Soumalias et al., 2024).
After taking in students’ reported preferences, the algorithm used to find an equilibrium needs to be able to solve maximization problems that find each student’s most-preferred course schedule subject to her budget and priority-specific prices. With thousands of courses, overlapping lecture times and prerequisite requirements create considerable scheduling constraints that students should not be expected to handle through reported preferences alone. To deal with these constraints, the maximization problem should prevent students from obtaining seats in courses that meet at the same time. Relatedly, it should accommodate course prerequisite constraints, blocking students from enrolling in courses they cannot obtain. As many degree programs require students to take certain courses in a given term, making these assignments outside the mechanism would reduce preference reporting mistakes. These modifications would make our mechanism simpler to implement in practice.
Finding a Pseudo-Market Equilibrium with Priorities with students’ reported preferences is a computationally intense problem. Budish et al. (2017) estimate that 20 times the computational time used for business school students at Wharton would be necessary to implement the A-CEEI mechanism for the full student body at Ohio State University. However, equilibrium computation for the PMP mechanism may be less challenging than this estimate suggests. Students in unrelated programs often have limited overlap in course interests. The structure of priority-specific prices simplifies the problem of finding a student’s demand, as the price is zero or unaffordable at all but one of a course’s priority levels. In our simulations, the average timing of the PMP mechanism improves on that of the A-CEEI mechanism by a factor of two (see Appendix D.2). Still, incorporating scheduling constraints and preference complementarities may be challenging and should be carefully addressed in practice.
Appendix A Appendix: Proofs
Proof of Theorem 1. Consider a course allocation problem $({\cal S},{\cal C},Q,{\cal V},\mathcal{R})$. We show that, for any $\beta>0$, there exists a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities.
Let $b=(b_{1},...,b_{N})$ be a budget vector that satisfies $1\leq\min_{s}(b_{s})\leq\max_{s}(b_{s})\leq 1+\beta$ for some $\beta>0$, and let $\overline{b}=1+\beta+\epsilon$ for a fixed $\epsilon>0$. Consider the $M$-dimensional set ${\cal T}=[0,R\overline{b}]^{M}$, which conveniently parameterizes priority-specific prices and allows us to look for a competitive equilibrium in a lower dimensional space. In particular, for each parameter $t\in{\cal T}$, course $c\in{\cal C}$, and level of priority $r\in{\{1,...,R\}}$, we define priority-specific prices as
$p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0).$ (A1)
With this definition, for each $t\in{\cal T}$ and $c\in{\cal C}$, there is a unique cutoff level of priority $r^{*}_{c}(t)\in{\{1,...,R+1\}}$ [^28] Note that $r^{*}_{c}(t)=R+1$ is only achieved if $t=R\overline{b}$. The cutoff level $r^{*}_{c}$ can always be reduced to $R$. such that for any $r\in\{1,...,R\}$, $p_{c,r}(t)$ satisfies
$p_{c,r}(t)\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}(t)\\ {}[0,\overline{b})&r=r^{*}_{c}(t)\\ [\overline{b},+\infty)&r<r^{*}_{c}(t)\end{cases}.$ (A2)
We will consider an auxiliary enlargement of this set, $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ and define $p_{c,r}(\tilde{t})$ using equation (A1) for $\tilde{t}\in\tilde{\cal T}$. For each $s\in{\cal S}$, define the demand function $d_{s}:\tilde{{\cal T}}\rightarrow 2^{\cal C}$ by
$d_{s}(\tilde{t})=\max_{\succsim_{s}}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}x^{\prime}_{s,c}\max(\tilde{t}_{c}-(r_{s,c}-1)\overline{b},0)\leq b_{s}+\tau_{s,x^{\prime}_{s}}\right\},$
where the $\tau_{s,x_{s}}$ are student- and schedule-specific taxes chosen to ensure that the demand is single-valued. For each course $c\in{\cal C}$, the excess demand function $z_{c}:\tilde{{\cal T}}\rightarrow\mathbb{Z}$ is defined by
$z_{c}(\tilde{t})=\sum_{s\in{\cal S}}x^{*}_{s,c}-q_{c},$
where $x^{*}_{s}=d_{s}(t^{*})$ for all $s\in{\cal S}$. Excess demand is bounded, as $-N\leq z_{c}\leq N-1$ for all $c\in{\cal C}$. We also define a budget surface for each student $s\in{\cal S}$ and schedule $x_{s}\subseteq{\cal C}$ as
$H(s,x_{s})=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s,c}\max(\tilde{t}_{c}-(r_{s,c}-1)\overline{b},0)=b_{s}+\tau_{s,x_{s}}\right\}.$
Unlike the case without priorities, the budget surface $H(s,x_{s})$ may not be a hyperplane (see Budish, 2011). Lemma B1 in Appendix B shows that it is still possible to choose $b_{s}$ and $\tau_{s,x_{s}}$ such that at most $M$ budget constraints intersect for any $\tilde{t}\in\tilde{{\cal T}}$.
Next, we define a truncation function trunc$:\tilde{{\cal T}}\rightarrow{\cal T}$, where for each $c\in{\cal C}$,
$\text{trunc}(\tilde{t})_{c}=\min\{R\overline{b},\max\{0,\tilde{t}_{c}\}\}.$
Also, for $\gamma\in(0,1/N)$, we define a tâttonnement price adjustment function $f:\tilde{{\cal T}}\rightarrow\tilde{{\cal T}}$ by
$f(\tilde{t})=\text{trunc}(\tilde{t})+\gamma z(\text{trunc}(\tilde{t})).$
Suppose that $f$ has a fixed point $\tilde{t}^{*}=f(\tilde{t}^{*}),$ and denote its truncation by $t^{*}=\text{trunc}(\tilde{t}^{*})$. We show that the prices $\{p_{c,r}(t^{*})\}_{c\in{\cal C},{r\in\{1,...,R\}}}$ defined by equation (A1), allocation $x^{*}_{s}=d_{s}(t^{*})$, and budgets $b^{*}_{s}=b_{s}+\tau_{s,x^{*}_{s}}$ for all $s\in{\cal S}$ constitute an exact competitive equilibrium (or ($0,\beta$)-Pseudo-Market Equilibrium with Priorities as in Definition 4). We might only need to slightly adjust $t^{*}$ to find another fixed point that satisfies condition $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}>0$.
  • The definition of the demand function $d_{s}(t^{*})$ implies that any course schedule that student $s$ prefers to $x^{*}_{s}=d_{s}(t^{*})$ must cost strictly more than $b^{*}_{s}=b_{s}+\tau_{s,x^{*}_{s}}$.
  • Prices $\{p_{c,r}(t^{*})\}_{c\in{\cal C},r\in{\{1,...,R\}}}$ and cutoff levels of priority defined by (A1) and (A2) ensure that condition ($\ast$2) of Definition 4 is satisfied.
[…]
  • $p_{c,1}(t^{*})=0$ implies $z_{c}(t^{*})\leq 0$. To see this, consider two cases. If $\tilde{t}_{c}^{*}\in(0,R\overline{b})$, the fixed point equation implies $z_{c}(t^{*})=0$. If $\tilde{t}_{c}^{*}\in[-1,0]$, we have $t_{c}^{*}\equiv\text{trunc}(\tilde{t}^{*}_{c})=0$ and $p_{c,1}(t^{*})=0$, so the fixed point equation ensures that $z_{c}(t^{*})\leq 0$.
  • To make sure the condition $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}>0$ is satisfied for any course $c$ with $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ (see condition ($\ast$2)), assume from the contrary that course $c$ has demand $\sum_{s\in{\cal S}}x^{*}_{s,c}>0$ but demand from students at the cutoff level of priority is 0: $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(t^{*})\}}x^{*}_{s,c}=0$. Then, we can consider $\hat{t}_{c}^{*}=t^{*}_{c}+\overline{b}-p_{c,r_{c}^{*}(t^{*})}$, where the students at the cutoff level of priority a face of price $\overline{b}$ and the prices at lower priority levels do not change. With $\hat{t}^{*}=(\hat{t}^{*}_{c},t^{*}_{-c})$, the cutoff increases to level $r^{*}_{c}(\hat{t}^{*})=r^{*}_{c}(t^{*})+1$, the prices become $p_{c,r^{*}_{c}(\hat{t}^{*})}=0$ and ${p}_{c,r^{*}_{c}(t^{*})}=\overline{b}$, and the adjustment does not change the demand for any student or excess demand for any course. If we still have $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(\hat{t}^{*})\}}x^{*}_{s,c}=0$, we repeat the price adjustment until the condition is satisfied.
    Finally, $z_{c}(\hat{t}^{*})=z_{c}(t^{*})=0$ implies $\hat{t}^{*}_{c}<R\overline{b}$ and $\hat{t}^{*}_{c}>t^{*}_{c}\geq 0$. Therefore, $\hat{t}^{*}_{c}=\text{trunc}(\hat{t}^{*}_{c})+z_{c}(\hat{t}^{*})$, meaning $\hat{t}^{*}$ satisfies both the fixed point equation and $\sum_{\{s\in{\cal S}:r_{s,c}=r^{*}_{c}(\hat{t}^{*})\}}x^{*}_{s,c}>0$.
Overall, if $f$ has a fixed point $\tilde{t}^{*}=f(\tilde{t}^{*})$, its truncation $t^{*}=\text{trunc}(\tilde{t}^{*})$ or adjustment $\hat{t}^{*}$ is a competitive equilibrium price vector for allocation $x^{*}_{s}=d_{s}(t^{*})$ and budgets $b_{s}^{*}$ for all $s\in{\cal S}$. Though, the fixed point of operator $f$ might fail to exist. Following Budish (2011), we define a “convexification” of $f$, $F:\tilde{\cal T}\rightarrow\tilde{\cal T}$, by
$F(\tilde{t})=co\{y:\exists\text{ a sequence }\tilde{t}^{w}\rightarrow\tilde{t},\tilde{t}\neq\tilde{t}^{w}\in\tilde{\cal T}\text{ such that }f(\tilde{t}^{w})\rightarrow y\},$
where $co$ denotes the convex hull of the set. $F$ is nonempty and convex and $\tilde{\cal T}$ is compact and convex. From Lemma 2.4 of Cromme and Diener (1991), $F$ is an upper hemicontinuous correspondence and hence has a fixed point by Kakutani’s fixed point theorem. We denote a fixed point by $\tilde{t}^{*}\in F(\tilde{t}^{*})$, and let again $t^{*}=\text{trunc}(\tilde{t}^{*})$ be its truncation.
The remainder of the proof uses $F$ to show that there exists a vector $t^{*}$ with excess demand at most a distance of $\sqrt{kM/2}$ from a perfect market-clearing excess demand vector. The reduction of the $M\!R$-dimensional space of prices $\{p_{c,r}\}_{c\in{\cal C},{r\in{\{1,...,R\}}}}$ to $M$-dimensional space $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ allows us to use the steps of Budish (2011) to establish the existence of a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities. For completeness, we provide adapted steps in Appendix B.∎
Proof of Theorem 2. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism and suppose each course $c\in{\cal C}$ instead has capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$. We show that the allocation $x^{*}$ is stable. Since each course is at full capacity, $x^{*}$ must be feasible. Definitions 4 and 5 ensure that $x^{*}_{s}$ is individually rational as well. Last, suppose that $(s,C)$ is a block. Then, $C=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x^{*}_{s}\cup C\}$, $C\neq x^{*}_{s}$, and for each $c\in{\cal C}$ with $c\notin x^{*}_{s}$, there is a student $s^{\prime}\in{\cal S}$ with $c\in x^{*}_{s^{\prime}}$ and $r_{s^{\prime},c}<r_{s,c}$. There must exist at least one course $c\in{\cal C}$ with $c\notin x^{*}_{s}$ such that $p^{*}_{c,r_{s,c}}>0$; otherwise, student $s$ should have been assigned $C$ instead of $x^{*}_{s}$. The structure of equilibrium prices implies that for any student $s^{\prime}$ with a lower priority than $s$ for course $c$ we must have $p^{*}_{c,r_{s^{\prime},c}}\geq\overline{b}$. Therefore, we obtain a contradiction: $x^{*}$ cannot possibly assign a student $s^{\prime}$ a seat in course $c$. Hence, there are no blocks, and the outcome of every Pseudo-Market with Priorities mechanism is approximately stable. ∎
Proof of Theorem 3. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism. We show that the allocation $x^{*}$ is approximately priority-constrained efficient. Suppose each course $c\in{\cal C}$ has capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$ and that $y$ is an allocation that Pareto dominates $x^{*}$. Allocation $y$ cannot assign more students to some course $c$ than $x^{*}$ does as it would imply that $\sum_{s\in{\cal S}}y^{*}_{s,c}>q^{\prime}_{c}$. If $y$ assigns fewer students to some course $c$ than $x^{*}$ does, then $\sum_{s\in{\cal S}:r_{s,c}\geq 1}x^{*}_{s,c}>\sum_{s\in{\cal S}:r_{s,c}\geq 1}y_{s,c}$, confirming that $y_{c}\nsucceq_{c}x_{c}^{*}$.
Now, suppose that allocation $y$ has the same number of seats assigned in each course. Then, there is a student $s^{\prime}$ who strictly prefers their assigned course schedule in $y$ to their schedule in $x^{*}$. For all students $s\in{\cal S}$, we must have that
$\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}y_{s,c}\geq b^{*}_{s}\geq\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}x^{*}_{s,c},$
with the first inequality holding strictly for student $s^{\prime}$. Across all students, we have that
$\sum_{s\in{\cal S}}\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}y_{s,c}>\sum_{s\in{\cal S}}\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}x^{*}_{s,c}.$
Rearranging,
$\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}=r}y_{s,c}>\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}=r}x^{*}_{s,c}.$
Hence, we obtain
$0<\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})=\sum_{c\in{\cal C}}\sum_{r=2}^{R}(p^{*}_{c,r}-p^{*}_{c,r-1})\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})+p^{*}_{c,1}\sum_{s:\,r_{s,c}\geq 1}(y_{s,c}-x^{*}_{s,c}).$
The final term must equal zero, as allocations $y$ and $x^{*}$ assign the same number of seats in each course, Also, $p^{*}_{c,r}-p^{*}_{c,r-1}\leq 0$ for all levels $r$, and we must have $\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})<0$ for some course $c$ and level $r$. This implies $y_{c}\nsucceq_{c}x^{*}_{c}$ and confirms that allocation $x^{*}$ is approximately priority-constrained efficient.∎
Proof of Theorem 4. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism and suppose that $\beta\leq\frac{1}{k-1}$. We show that the allocation $x^{*}$ has schedule envy bounded by a single course toward students of weakly lower levels of priority. Consider two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$. Denote the prices faced by $s$ and $s^{\prime}$ as $p^{*}_{s}=\{p^{*}_{c,r_{s,c}}\}_{c\in{\cal C}}$ and $p^{*}_{s^{\prime}}=\{p^{*}_{c,r_{s^{\prime},c}}\}_{c\in{\cal C}}$ and the course schedule assigned to student $s^{\prime}$ as $x^{*}_{s^{\prime}}=\{c_{j_{1}},...,c_{j_{k^{\prime}}}\}$ for $c_{j_{1}},...,c_{j_{k^{\prime}}}\in\{1,...,M\}$ and $k^{\prime}\leq k$. Suppose that $s$ envies $s^{\prime}$ and $s$ cannot afford the course schedule of $s^{\prime}$ even if one course from $x^{*}_{s^{\prime}}$ is dropped. That is,
$p^{*}_{s}\cdot x^{*}_{s^{\prime}}\backslash\{c_{j_{\ell}}\}>b^{*}_{s}$
for $\ell=1,...,k^{\prime}$. Since $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$, $p^{*}_{s^{\prime}}\geq p^{*}_{s}$, guaranteeing that
$p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}\backslash\{c_{j_{\ell}}\}>b^{*}_{s}$
for $\ell=1,...,k^{\prime}$. As $b^{*}_{s^{\prime}}\geq p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}$, summing these inequalities over $\ell$, we obtain
$(k^{\prime}-1)b^{*}_{s^{\prime}}\geq(k^{\prime}-1)p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}>k^{\prime}b^{*}_{s}.$
This implies $\frac{b^{*}_{s^{\prime}}}{b^{*}_{s}}>\frac{k^{\prime}}{k^{\prime}-1}\geq\frac{k}{k-1}\geq 1+\beta$, which contradicts how budgets are allocated.∎
Proof of Theorem 5. We first formally define a direct mechanism, a semi-anonymous direct mechanism, and the property of being strategy-proof in the large. Next, we show that the PMP mechanism is a semi-anonymous direct mechanism that is envy-free but for tie-breaking. The result then follows from Appendix C in Azevedo and Budish (2019).
Consider a sequence of environments labeled by the number of students $|{\cal S}^{N}|\equiv N$. The set of courses ${\cal C}$ is fixed, with the number of courses equal to $M$, but the capacity of each course $q^{N}_{c}$ can vary. Each student $s\in{\cal S}^{N}$ has a type $v_{s}=(\succeq_{s},r_{s})$ that describes her preferences $\succeq_{s}$ over course schedules and her priorities $\{r_{s,c}\}_{c\in\cal C}$. The set of all possible types is denoted as ${\cal V}$. An allocation $x\in X^{N}_{0}\equiv\{0,1\}^{MN}$ specifies a schedule $x_{s}$ for each student $s$. Letting $X=\Delta X_{0}$ denote the set of random allocations, we assume that each student has a von Neumann-Morgenstern utility function $u_{v}:X\rightarrow[0,1]$ consistent with her type $v\in{\cal V}$.
We consider a sequence of direct mechanisms $\{\Phi^{N}\}_{N\in\mathbb{N}}$ that, for each report of student preferences, assigns a distribution over course allocations; that is, $\Phi^{N}:{\cal V}^{N}\rightarrow\Delta(X_{0}^{N})$. The PMP mechanism is an example of a semi-anonymous direct mechanism, where agents are divided into groups, agents within each group are treated the same way, and agents across groups can be treated differently. Formally, we partition ${\cal S}^{N}$ into groups according to their course priorities $\left\{S^{N}_{g}\right\}_{g\in\{1,...,R\}^{M}}$. This partition places students who have the same level of priority in every course in one group. All students in the same group face the same prices in the PMP mechanism and the same lottery over budgets. However, these students might have different preferences $\succeq_{s}$ and, hence, types $v_{s}$. We assume that each student can misrepresent her preferences, but not her priorities.
To define the property of a direct mechanism being strategy-proof in the large, we consider function $\phi^{N}:{\cal V}\times\Delta{\cal V}\rightarrow X$ for each $N$ according to
$\phi^{N}(v_{s},m)=\sum_{v_{-s}\in{\cal V}^{N-1}}\Phi_{s}^{N}(v_{s},v_{-s})Pr(v_{-s}|v_{-s}\sim iid(m)),$
where $\Phi_{s}^{N}(v_{s},v_{-s})$ denotes the course schedule obtained by student $s$ when she reports type $v_{s}$ and all other students report $v_{-s}$. $Pr(v_{-s}|v_{-s}\sim iid(m))$ denotes the probability that profile $v_{-s}$ is realized when other students’ types $v_{-s}$ are independent and identically distributed according to $m\in\Delta{\cal V}$. In other words, $\phi^{N}(v_{s},m)$ describes the random outcome that student $s$ expects to receive when she reports $v_{s}$ and the other students’ types are distributed independently and identically according to $m$—and they report their types truthfully. Using this notation, we have the following definition.
Definition A1.
A semi-anonymous direct mechanism $\{\Phi^{N}\}_{N\in\mathbb{N}}$ is strategy-proof in the large if, for any random distribution of reports by other students $m\in\Delta{\cal V}$ with full support and $\varepsilon>0$, there exists $N_{0}$ such that, for all $N\geq N_{0}$ and all $v^{\prime}_{s},v_{s}\in{\cal V}$ where students can misrepresent only their preferences, we have $u_{v_{s}}[\phi^{N}(v_{s},m)]\geq u_{v_{s}}[\phi^{N}(v^{\prime}_{s},m)]-\varepsilon.$
In other words, in a large enough market, reporting preferences truthfully is approximately optimal, for any independent and identical distribution of the other students’ types that has full support. Students can only misrepresent their preferences, not their priorities.
Appendix C in Azevedo and Budish (2019) provides an argument explaining why any semi-anonymous mechanism that is envy-free but for tie-breaking is strategy-proof in the large. We define envy-freeness but for tie-breaking below.
Definition A2.
[…]
Appendix B Proof of Theorem 1 (Omitted Details)
Appendix A shows how Theorem 1 of Budish (2011) can be extended to accommodate priority-specific course prices. Below, we confirm it is possible to choose budgets $b_{s}$ and $\tau_{s,x_{s}}$ such that at most $M$ budget constraints intersect for any $\tilde{t}\in\tilde{{\cal T}}$. This fact is used on p. A.
Lemma B1.
One can choose taxes $\left\{\tau_{s,x_{s}}\right\}_{s\in{\cal S},x_{s}\subseteq{{\cal C}}}$ that satisfy the following conditions:
  • (i)
    Taxes are small: $-\varepsilon<\tau_{s,x_{s}}<\varepsilon$.
  • (ii)
    Taxes favor more preferred bundles: $\tau_{s,x_{s}}>\tau_{s,x^{\prime}_{s}}$ for $x^{\prime}_{s}\succ_{s}x_{s}$.
  • (iii)
    The inequality bounds are preserved:
    $1\leq\min_{s,x_{s}}(b_{s}+\tau_{s,x_{s}})\leq\max_{s,x_{s}}(b_{s}+\tau_{s,x_{s}})\leq 1+\beta.$
  • (iv)
    No perturbed budgets are equal: $b_{s}+\tau_{s,x_{s}}\neq b_{s^{\prime}}+\tau_{s^{\prime},x_{s^{\prime}}}$.
  • (v)
    There is no auxiliary price vector $\tilde{t}\in\tilde{\cal T}$ at which more than $M$ budget constraints $H(s,x_{s})$ intersect.
Proof.
Fix an allocation $x=(x_{s})_{s\in{\cal S}}$. Budish (2011) shows that it is possible to choose $\left\{\tau_{s,x_{s}}\right\}_{s\in{\cal S},x_{s}\subseteq{\cal C}}$ that satisfy the first four conditions. We now establish that it is always possible to slightly change taxes such that condition (v) is also satisfied. Let us assume that more than $M$ budget constraints $H(s,x_{s})$ intersect and denote
=${\cal I}=\left\{s\in{\cal S}:\cap_{s}H(s,x_{s})\neq\varnothing\right\},$
with $|{\cal I}|>M.$ For each $\tilde{t}\in\tilde{\cal T}$, consider prices $\left\{p_{c,r}(\tilde{t})\right\}_{c\in{\cal C},r\in{\{1,...,R\}}}$ defined by equation (A1) and the cutoffs defined by (A2). The definition of cutoffs $r^{*}_{c}(\tilde{t})$ implies that
[…]
In addition, the definition of cutoffs implies $p_{c,r_{s,c}}\geq\overline{b}$ for $1\leq r_{s,c}<r^{*}_{c}(\tilde{t})$. Therefore, a seat in course $c$ is not allocated to agent $s$ for $r_{s,c}<r^{*}_{c}(\tilde{t})$; that is, $x_{s,c}=0.$ Hence, we also have
[…]
Therefore, we obtain that $x_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})$ might be non-zero only if $r_{s,c}=r^{*}_{c}(\tilde{t})$. That is, entries in agent $s$’s budget constraint, $s\in{\cal I}$, are non-zero only if $r_{s,c}=r^{*}_{c}(\tilde{t})$. Denote $p_{c}(\tilde{t})\equiv p_{r^{*}_{c}}(\tilde{t})$. Hence, we obtain that the set of equations $\left\{s\in{\cal I}:H(s,x_{s})=0\right\}$ is the set of linear equations with coefficients $x_{s,c}\in\{0,1\}$:
$\left\{\begin{array}[]{cc}x_{s,1}\cdot p_{1}(\tilde{t})+x_{s,2}\cdot p_{2}(\tilde{t})+...+x_{s,M}\cdot p_{M}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ x_{s^{\prime},1}\cdot p_{1}(\tilde{t})+x_{s^{\prime},2}\cdot p_{2}(\tilde{t})+...+x_{s^{\prime},M}\cdot p_{M}(\tilde{t})&=b_{s^{\prime}}+\tau_{s^{\prime}x}\end{array}\right.$
for $s,s^{\prime}\in{\cal I}$. Since $|{\cal I}|>M$ for any $\tilde{t}\in{\cal T}$, there are at most $M$ independent linear equations of prices. So, the Rouché–Capelli theorem implies that we can choose $\tau_{s,x_{s}}$ such that only at most $M$ equations are satisfied for any $\tilde{t}$. [^29] Note that when we change $\tau_{s,x_{s}}$, some other budget constraints might start intersecting. Since the set of possible intersecting budget constraints is finite and $\tau_{s,x_{s}}$ varies continuously, we can always choose $\tau_{s,x_{s}}$ without influencing the intersection property of the other budget constraints.
Our proof of Theorem 1 in Appendix A extends steps 1-3 of the proof of Theorem 1 from Budish (2011) to accommodate priority-specific course prices. The rest of this section adapts the remaining steps from this proof to our setting.
Step 4. Suppose that $t^{*}$ is not on any budget constraint. We show that $t^{*}$ is an exact competitive equilibrium price vector. There is a neighborhood around $t^{*}$ where each agent’s demand is unchanging in price. At price $t^{*}$, $f$ is continuous, and as a result, $F(t^{*})=f(t^{*})$.
  • If $t^{*}=\tilde{t}^{*}$, then $F(\tilde{t}^{*})=F(t^{*})=f(t^{*})$ and, thus, $t^{*}=\tilde{t}^{*}\in F(\tilde{t}^{*})=f(\tilde{t}^{*})$. Therefore, $t^{*}$ is a fixed point. Hence, as we established earlier, it is an exact competitive equilibrium price vector.
  • If $t^{*}\neq\tilde{t}^{*}$, we establish the following lemma.
    Lemma B2.
    For any $\tilde{t}\in\tilde{\cal T}\setminus{\cal T}$, (i) $f(\tilde{t})=f($$(\tilde{t}))$ and (ii) $F(\tilde{t})\subseteq F($$(\tilde{t}))$.
    Proof.
[…]
    As a result, since $F(t^{*})=f(t^{*})$ and $\tilde{t}^{*}\in F(\tilde{t}^{*})$, $\tilde{t}^{*}\in F(t^{*})=f(t^{*})=f(\tilde{t}^{*})$ and, thus, $\tilde{t}^{*}=f(\tilde{t}^{*})$. Therefore, $t^{*}$ is an exact competitive equilibrium price vector.
Step 5. Next, suppose that $t^{*}$ is on $1\leq L\leq M$ budget constraints. We denote $\Phi=\{0,1\}^{L}$ and construct a set of $2^{L}$ price vectors $\{t^{\phi}\}_{\phi\in\Phi}$ that satisfy the following conditions:
  1. 1.
    Each $t^{\phi}$ is close enough to $t^{*}$ such that there is a path from $t^{\phi}$ to $t^{*}$ that does not cross any budget constraint.
  2. 2.
    Each $t^{\phi}$ is on the “affordable” side of the $\ell$th budget constraint if $\phi_{\ell}=0$ and is on the “unaffordable” side if $\phi_{\ell}=1$.
To construct vectors $\{t^{\phi}\}_{\phi\in\Phi}$, note that each of the $L$ intersecting budget constraints defines two sets:
$H^{0}_{\ell}$ =$=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)\leq b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
$H^{1}_{\ell}$ =$=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)>b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
The first set delineates the set of prices for which agent $s_{\ell}$ can afford schedule $x_{s_{\ell}}$, whereas the second set delineates the set of prices for which agent $s_{\ell}$ can’t afford $x_{s_{\ell}}$. Let $\phi=(\phi_{1},...,\phi_{L})\in\Phi$ be an $L$-dimensional vector of zeros and ones, and the polytope $\pi(\phi):=\cap_{\ell=1}^{L}H^{\phi_{\ell}}_{\ell}$ be the set of points in ${\cal T}$ that belongs to the intersection of sets indexed by $\phi$. Let $H=\{H(s,x_{s})\}_{s\in{\cal I},x_{s}\subseteq{\cal C}}$ be the finite set of all budget constraints formed by any student-schedule pair $(s,x_{s})$. We then define the distance
$\delta<\inf_{\tilde{t}^{\prime\prime}\in{\cal T},H\in{\cal H}}\{||(t^{*}-\tilde{t}^{\prime\prime})||_{2}:\tilde{t}^{\prime\prime}\in H,t^{*}\notin H\},$
which is such that any budget constraint that $t^{*}$ does not belong to is further than $\delta$ away from $t^{*}$. Now, for each $\phi\in\Phi$, we define $t^{\phi}$ by an arbitrary element of $\pi(\phi)\cap B_{\delta}(t^{*})$, where $B_{\delta}(t^{*})$ is a $\delta$-ball of $t^{*}$. Then, the set of price vectors $\{t^{\phi}\}_{\phi\in\Phi}$ satisfies the two requirements above.
Step 6. We now show that a perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$. For this purpose, we first show that for any $y\in F(t^{*})$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ Take some $y\in F(t^{*})$. Consider a sequence $t^{w}\rightarrow t^{*},t^{*}\neq t^{w}\in\tilde{\cal T}\text{ such that }f(t^{w})\rightarrow y^{\prime}$. Note that the sequence $t^{w}$ consists of a finite number of subsequences $t^{w,\phi}\in\pi(\phi)\cap B_{\delta}(t^{*})$ for some $\phi\in\Phi$ [^30] Some subsequences can have only a finite number of elements. Since all elements of the subsequence $t^{w,\phi}$ are on the same side of set $H^{\phi}_{\ell}$, every agent has the same choice at every point. Hence, if the subsequence has an infinite number of elements, we have $t^{w,\phi}\rightarrow t^{*}$ and
$f(t^{w,\phi})\rightarrow t^{*}+\gamma z(t^{\phi}).$
Therefore, the limit of $f(t^{w})$ for the original sequence $t^{w}$ must be also $t^{*}+\gamma z(t^{\phi^{\prime}})$ for some $\phi^{\prime}\in\Phi$. So, $y^{\prime}=t^{*}+\gamma z(t^{\phi^{\prime}})$. Since, by definition, $y$ is a convex combination of such $y^{\prime}$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$
Lemma B2 implies that $\tilde{t}^{*}\in F(\tilde{t}^{*})\subseteq F(t^{*})$. Hence, we must have
$\tilde{t}^{*}=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi}).$
for some $\{\lambda_{\phi}\}_{\phi\in\Phi}$ with $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ We also denote
$\zeta=\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})=\frac{\tilde{t}^{*}-t^{*}}{\gamma}.$
We note that $\zeta\leq 0$ and $\zeta_{c}<0$ imply $t_{c}^{*}=0$. Hence, vector $\zeta$ is a perfect market-clearing excess demand vector, and it lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$.
Steps 7-9. The structure of excess demand has the same geometric structure as in Budish (2011). In particular, let $L^{\prime}$ be the number of agents whose budget constraints intersect at price $t^{*}$, and renumber agents so that $s=1,...,L^{\prime}$. We denote the number of budgets of student $s$ that intersect at $t^{*}$ as $w_{s}$. Since at most $M$ budgets constraints can intersect, we must have $L\equiv\sum_{s=1}^{L^{\prime}}w_{s}\leq M$. We also denote the bundles pertaining to $s$’s budget constraints as $x_{s}^{1}\succ...\succ x_{s}^{w_{s}}$.
Similarly to Budish (2011), we consider bundles that $s$ demands at prices near $t^{*}$. In the set $H^{0}(s,x^{1}_{s})$, agent $s$ can purchase her favorite bundle $x^{1}_{s}$. Hence, one does not need to know whether prices belong to sets $H^{0}(s,x_{s}^{2}),...,H^{0}(s,x_{s}^{w_{s}})$. Let us denote the demand for prices at $H^{0}(s,x^{1}_{s})\cap B_{\delta}(t^{*})\cap\tilde{\cal T}$ as $d_{s}^{0}$. Similarly, we consider prices in $H^{1}(s,x^{m}_{s})\cap H^{0}(s,x^{m+1}_{s})$ and denote the corresponding demands as $d_{s}^{m}$ for $m=1,...,w_{s}$. Overall, agent $s=1,...,L^{\prime}$ purchases $w_{s}+1$ distinct bundles at prices near to $t^{*}$.
Let us denote the excess demand of the remaining agents as
$z_{S\backslash\{1,...,L^{\prime}\}}(t^{*})=\sum_{s=L^{\prime}+1}^{S}d_{s}(t^{*})-q.$
A perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$ with the elements
$z_{S\backslash\{1,...,L^{\prime}\}}(t^{*})+\sum_{s=1}^{L^{\prime}}\sum_{f=1}^{w_{s}}a_{s}^{f}d^{f}_{s}$
where $0\leq a_{s}^{f}\leq 1,s=1,...,L^{\prime},f=1,...,w_{s}$ and $\sum_{f=1}^{w_{s}}a_{s}^{f}=1,$ $s=1,...,L^{\prime}$. Budish (2011) shows in Step 8 of Theorem 1 that there exists an element of the above geometric structure, $z(t^{\phi^{\prime}})$, within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector. Step 9 explains how to adjust agent budgets to find an approximate competitive equilibrium at $t^{*}$ with excess demands equal to $z(t^{\phi^{\prime}})$. These purely mathematical arguments remain unchanged from the original paper.∎
Appendix C The Student Utility Model Calibration
This appendix provides details on the method of simulated moments approach used to calibrate the student utility model. Section C.1 outlines the student utility model. Section C.2 discusses the necessary adjustments to the university’s course reserves. Section C.3 presents the empirical moments we use to calibrate the model. Section C.4 outlines the iterated process we use to calibrate the model and presents the calibration results. Section C.5 provides details on model fit.
C.1 Model
We consider a model with student horizontal preferences, course vertical components, and an interaction term. In particular, $u_{sc}$, student $s$’s utility from taking course $c$, takes the following parametric form:
$u_{sc}=Y^{\prime}_{s}\theta+Y^{\prime}_{s}\Gamma H_{c}+z_{c}+\varepsilon_{sc},$ (C1)
where:
  • $Y_{s}$ and $H_{c}$ are vectors of student $s$’s observed characteristics and course $c$’s observed characteristics, respectively. Both vectors are binary vectors having all but one element as zero. Denote the set of colleges by ${\cal A}=\{A,...,G\}$ and years of study by ${\cal Y}=\{1,...,4\}$. Then, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ determines the college and year of study of student $s$ and $H_{c}\in\{0,1\}^{|{\cal A}|}$ determines the college of course $c$.
  • $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ and $z=\{z_{c}\}_{c\in{\cal C}}$ are two vectors of unobserved coefficients; $\Gamma$ is a matrix of unobserved coefficients conformable with vectors $Y_{s}$ and $H_{c}$.
  • $\{\varepsilon_{sc}\}_{s\in{\cal S},c\in{\cal C}}$ are random utility components that are assumed to be identically and independently distributed according to the standard normal distribution $\varepsilon_{sc}\sim N(0,1)$.
We assume that the utility from taking multiple courses equals the sum of utilities of taking individual courses. The utility of taking no course is normalized to $u_{s0}=0$. Each student can be enrolled in up to a maximum of five courses.
The parametric form for student utilities in equation (C1) has three non-random components. The first, $Y^{\prime}_{s}\theta$ is a horizontal component that is determined solely by the student’s characteristics. The vector of coefficients $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ that has $28$ elements in total to calibrate. The horizontal component determines the importance of taking courses for student $s$, with a larger value meaning student $s$ wants to enroll in more courses. This accounts for variability in the data in the number of courses taken by students from different colleges and years of study, as shown in Table C.1.
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. The matrix $\Gamma$ accounts for heterogeneity in patterns of course enrollment across colleges observed in the data demonstrated in Table C.2. For example, students from college $A$ take $71\%$ of courses from their home college, while more than $95\%$ of courses taken by students from college $C$ are courses outside their home college.
The third component $z_{c}$ is a vertical component that determines the attractiveness of course $c$. This component is common among students and determines the quality of the course. The number of vertical components coincides with the number of courses, $756$.
Table C.1: The average number of courses taken by students.
Year 1 Year 2 Year 3 Year 4
A 3.6 3.4 3.1 2.9
B 4.0 4.3 3.9 3.0
C 4.7 4.3 4.2 3.5
D 4.4 4.3 4.1 3.4
E 4.3 4.3 3.8 3.3
F 4.2 4.0 3.8 2.9
G 4.6 4.7 4.7 4.1
Notes: Each row of this table presents the average number of courses taken by students from the same college and different years of study.
Table C.2: The percentages of courses taken by students across colleges
% A B C D E F G
A 71 0.47 1.9 17 1.7 5 2.7
B 1.9 44 0.61 18 19 12 4.1
C 10 1.3 4.5 37 29 15 2.4
D 2.5 0.98 0.83 64 12 13 6.8
E 1.2 2.6 0.14 24 54 16 2.1
F 1.5 1.8 0.84 23 18 52 2.3
G 1.3 0.39 0.26 31 7.7 6.1 53
Notes: The table the percentage of courses taken by students across colleges. The rows correspond to students’ colleges, and the columns correspond to courses’ colleges. For example, the cell in row $2$ and column $1$ shows that $1.9\%$ of courses taken by students from college $B$ are courses from college $A$.
Parameter normalization. To correctly interpret the interaction term $\Gamma$ and the vertical component $z$, we need to further normalize the model’s parameters. As described above, for student $s$ from college $a$ and course $c$ from college $a^{\prime}$, we have $Y^{\prime}_{s}\Gamma H_{c}=\gamma_{aa^{\prime}}$. Let $S_{a}$ and $S_{ay}$ be the set of students from college $a$ and the set of students from college $a$ and year $y$, respectively. The average utility of students from college $a$ from taking course $c$ from college $a^{\prime}$ equals
$\overline{u}_{ac}=\frac{1}{|S_{a}|}\sum_{s\in{\cal S}_{a}}u_{sc}=\overline{\theta}_{a}+\gamma_{aa^{\prime}}+z_{c},$
where $\overline{\theta}_{a}=\frac{1}{|S_{a}|}\sum_{y}|S_{ay}|\theta_{ay}$ is the average student-specific component among students in college $a$. We have three additive parameters that lead to the same value of the utility components $\overline{u}_{ac}$ and need a normalization.
One possible normalization is $\overline{\theta}_{a}=0$ and $\gamma_{aa}=0$ for $a=1,...,7$. This normalization leads to the interpretation of $z_{c}$ as the average popularity of the course among students from the same college. Then, $\gamma_{aa^{\prime}}$ is an increase in student’s utility from taking a course in college $a^{\prime}$ compared to a course in student’s college $a$ with the same vertical component.
Students’ choice sets. The student choice sets are unobserved in the data. We do not observe students’ past enrollments, and, as a result, we do not know each student’s set of courses for which she satisfies course prerequisites. In addition, students’ choice sets are influenced by the courses chosen by students with earlier time-slots in the course allocation process, and, hence, students’ choice sets are determined endogenously. There are further complications associated with the changing of course reserves during the course allocation process, which are described in Section C.2.
These complications require us to take a stance on how student’s choice sets are formed. We assume that each student’s choice set is limited to $80$ courses. The courses in the student’s choice set are drawn randomly such that the probability that a course from college $a^{\prime}$ is drawn equals the share of students from the same college-year who are enrolled in courses in college $a^{\prime}$ in the actual data. Additionally, we make adjustments for courses with course reservations. For each course reservation with $r$ reserved seats, we ensure that the course is included in the choice sets of at least $2r$ students who qualify for the reservation. This adjustment prevents course reservations from artificially limiting a course?s capacity due to an insufficient number of qualified students. [^31] If there are fewer than $2r$ qualified students, the course is included in the choice set of all such students.
Random choice sets are a limitation of our model. Any changes in the way we draw the course lists could potentially change the calibrated parameter values. Crawford, Griffith and Iaria (2021) provide an insightful discussion of preference estimation techniques with unobserved choice sets. Still, the size of the student choice sets we consider is reasonable, as previous studies report that students typically consider only a limited subset of courses in a given semester (see Budish and Cantillon, 2012; Diebold and Bichler, 2017). We explore how the performance of the Pseudo-Market with Priorities mechanism changes for various sizes of choice sets (and the size of the noise parameter) using simulations in Appendix D.4.
C.2 Course Reserve Adjustment
We use the method of simulated moments to calibrate the student utility model parameters. We need, however, to perform some adjustments to course reserves. Course reserves are typically set large at the beginning of the course allocation process and relaxed at later stages. Hence, the final course allocation can violate the initial course reserves. In fact, $170$ out of $756$ courses violate initial course reservations in the data.
We next describe the necessary and sufficient conditions for the existence of a matching that assigns a given population of students to the course and satisfies the course reservation criteria. Afterward, we describe the process of adjusting course reserves to ensure that the number of reserved seats is consistent with the course seat assignment observed in the data.
Two course reserve criteria. Consider a course with a capacity of $q$ seats. Denote the number of course reserves for students who satisfy criterion $A_{i}$ (student’s department and year of study) as $\mathfrak{r}_{A_{i}}$, $i=1,2$. We assume that $\mathfrak{r}_{A_{1}}\geq 0$ or $\mathfrak{r}_{A_{2}}\geq 0$ and $\mathfrak{r}_{A_{1}}+\mathfrak{r}_{A_{2}}\leq q$. One should interpret the reservation such that if a student satisfies both $A_{1}$ and $A_{2}$, she is eligible to take a seat satisfying either criterion, but if she takes $A_{1}$-type seats it does not decrease the number of $A_{2}$-type seats. This interpretation is consistent with the actual course enrollment process in the data. The number of students who satisfy criterion $A_{i}$ and who are matched to the course in the final allocation is denoted by $x_{A_{i}}$, $i=1,2$. Students who do not satisfy any course reserve criteria are called 00-students, and we denote the number of such students assigned to the course as $x_{0}$. Then, the course assignment satisfies course reserves if and only if
$\left\{\begin{array}[]{l}x_{0}+x_{A_{1}\cup A_{2}}\leq q\\[-2.84526pt] x_{0}\leq q-\mathfrak{r}_{A_{1}}-\mathfrak{r}_{A_{2}}\\[-2.84526pt] x_{0}+x_{A_{1}\backslash A_{2}}\leq q-\mathfrak{r}_{A_{2}}\\[-2.84526pt] x_{0}+x_{A_{2}\backslash A_{1}}\leq q-\mathfrak{r}_{A_{1}}\end{array}\right..$ (C2)
To see that the above conditions are necessary, we consider the criteria one by one. The first inequality corresponds to the criterion that puts no restrictions on students, the second inequality corresponds to students who do not satisfy $A_{1}$ and $A_{2}$, the third corresponds to students who do not satisfy $A_{2}$, and the fourth corresponds to students who do not satisfy $A_{1}$. To show sufficiency, we first enroll students $x_{0}$, $x_{A_{1}\backslash A_{2}}$, and $x_{A_{2}\backslash A_{1}}$. It is clearly possible to do so satisfying all criteria. The remaining students $x_{A_{1}\cap A_{2}}$ are constrained by the total course capacity. The first inequality guarantees that they could be enrolled as well.
The general case. The number of constraints increases exponentially with the number of separate seat reservations a course has. However, there are 4 or fewer seat reservations per course in the university data, which makes the analysis computationally manageable. Below, we present the necessary and sufficient conditions that guarantee the existence of a matching that satisfies multiple course reserve criteria that are based on the seminal Hall’s Marriage Theorem (see Hall, 1935).
To state the theorem, let us consider a set of students $S$ and course reserves criteria $A_{1},...,A_{I}$ for some $I\geq 1$. The number of reserved seats corresponding to criterion $A_{i}$ equals $\mathfrak{r}_{A_{i}}\geq 0$. We denote the set of available course seats as $Y$ with $|Y|=q$ and assume $\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}\leq q$. Each student could potentially satisfy multiple criteria. We consider a bipartite graph $(S,Y,E)$ where $E$ is the set of edges connecting student and seats. A student $s\in{\cal S}$ is connected to seat $y\in Y$ with an edge if student $s$ is eligible to take seat $y$. For example, if there are no course reserves, each student is connected with an edge to each seat ($q$ edges). If there is one course reservation criterion $A$, each 00-student has $q-\mathfrak{r}_{A}$ edges, and each student who satisfies criterion $A$ has $q$ edges.
For a subset of students $S^{\prime}\subseteq S$, let $N(S^{\prime})$ denote the set of seats in $Y$ that are connected to at least one student in $S$. We call these seats the neighbors of $S^{\prime}$. A perfect matching is a matching between students and seats in which each student is assigned exactly one seat. We are now ready to state the theorem.
Theorem (Hall’s Marriage Theorem).
The necessary and sufficient condition for the existence of a perfect matching is that every subset $S^{\prime}\subseteq S$ is connected to as many neighbors as the number of its elements
$|S^{\prime}|\leq|N(S^{\prime})|.$ (C3)
It is not hard to see that the conditions of Hall’s Marriage Theorem transform to the system of inequalities (C2) when there are only two criteria. To obtain the necessary and sufficient constraints in the general case, we need to consider only those inequalities where for a fixed set of seats $N(S^{\prime})$ there is the maximum possible $|S^{\prime}|$.
  • The number of students who satisfy no criteria, that is, 00-students
    $x_{0}\leq q-\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}.$
  • The number of students who satisfy only criterion $A_{i}$ or no criteria (adding 00-students only increases the number on the left-hand side of inequality (C3) and does not change the number on the right-hand side of the inequality)
    $x_{A_{i}\backslash\cup_{j\neq i}A_{j}}+x_{0}\leq q-\sum_{j\neq i}\mathfrak{r}_{A_{j}}$
  • Let us now denote =${\cal I}=\{1,...,I\}$ and take some subset of available course reserve criteria ${\cal I^{\prime}}\subseteq{\cal I}$. We consider the number of students who satisfy at least one criterion in the set $\{A_{i}:i\in{\cal I^{\prime}}\}$ or no criteria and who do not satisfy the criteria in the complementary set $\superscript\{A_{i}:i\in{\cal I}\backslash{\cal I^{\prime}}\}$. Hall’s Marriage Theorem implies that the number of these students must satisfy
    $x_{\cup_{i\in{\cal I^{\prime}}}A_{i}\backslash\cup_{j\notin{\cal I^{\prime}}}A_{j}}+x_{0}\leq q-\sum_{j\notin{\cal I^{\prime}}}\mathfrak{r}_{A_{j}}.$ (C4)
Overall, Hall’s Marriage Theorem implies that the set of inequalities (C4) for all possible subsets of criteria ${\cal I^{\prime}}\subseteq{\cal I}$ provides the necessary and sufficient conditions for the existence of a matching of students to course seats that satisfies the course reserve criteria.
Course reserve adjustment. Using the above conditions, we have identified 170 out of 756 courses that violate the initial course reserve criteria in the university data. For these courses, we order course seat reservations by year of study in decreasing order. We then take the first course reserve and decrease the number of its reserves by one. If the Hall’s inequalities are still violated, we move to the next one and also decrease the number of its reserves by one. If decreasing all course reserves by one does not make matching possible, we repeat this procedure. Clearly, the procedure terminates in a finite number of steps, as a matching with zero course reserves always possible. [^32] For over-enrolled courses, we first set $q$ equal the maximum between course capacity and the number of students assigned in the university data.
C.3 Moment Conditions
For each set of parameter values and the realization of random components of the student utility model described in Section C.1, we use the Random Serial Dictatorship with adjusted course reserves to obtain the allocations of courses to students. Then, we use these allocations to calculate the simulated moments specified below.
For each horizontal component of students’ preferences $\theta_{ay},a\in{\cal A},y\in{\cal Y}$, we consider the number of courses taken by students with the same characteristics as a moment, i.e.,
$\breve{m}_{ay}=\sum_{s\in{\cal S}_{ay}}\sum_{c\in{\cal C}}x_{s,c},$
where $x_{s,c}\in\{0,1\}$ indicates student $s$’s enrollment into course $c$ and $S_{ay}$ is the set of students from college $a$ and year $y$.
Similarly, the number of students from college $a$ who enrolled in courses in college $a^{\prime}$ is a natural moment for the interaction term $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$; that is,
$\hat{m}_{aa^{\prime}}=\sum_{s\in{\cal S}_{a}}\sum_{c\in{\cal C}_{a^{\prime}}}x_{s,c},$
where $C_{a^{\prime}}$ is the set of courses in college $a^{\prime}$ and $S_{a}$ is the set of students from college $a$.
For the vertical component of course-specific qualities $z_{c},c\in{\cal C}$, it is natural to consider the course enrollment as a moment, i.e.,
$\tilde{m}_{c}=\sum_{s\in{\cal S}}x_{s,c}.$
Courses at maximum capacity. A critical aspect of the data is that $85$ out of $756$ courses are at maximum capacity in the data, which creates a problem with identifying course-specific parameters $z_{c}$ for such courses. To address this issue, we replace moments $\tilde{m}_{c}$ for such courses with other moments based on the timing of course enrollment. We discuss these moments in more detail below.
To address the problem of identifying the vertical components $z_{c}$ for courses $c\in{\cal C}$ at maximum capacity, we use a crucial aspect of the data—the timing of student course enrollment. To track student course enrollment, we first define student ranking.
Definition C1.
The rank of student $s$, $\text{rank}_{s}$, is her position in the order of students to enroll in the Random Serial Dictatorship with course reserves after breaking ties.
Hence, more popular courses are filled up earlier with students of a smaller rank. Less popular courses are filled up later with students of a larger rank. [^33] More senior students have earlier time slots in the registration process. Hence, the measuring popularity by how fast a course fills up is biased towards the popularity of courses among more senior students.
To define student ranking in the university data, we considered the earliest timestamp of each student in our dataset; that is, the time students registered for their first course. Using these timestamps, we rank students in the university data. [^34] An alternative measure of how fast courses fill up is course registration times. Both student ranking and course registration time are closely related in our simulations, as each student registers for all desired courses simultaneously. This connection is less evident in the data. Many students do not register for courses at the same time. This choice pattern is probably connected with students changing their minds or new possibilities arising during the actual course registration process. We found that student ranking better connects our simulations to the actual university data than course registration time. To measure course popularity, we use the mean student rank enrolled in the course:
$Mean(\{\text{rank}_{s}:x_{s,c}=1,\,s\in{\cal S}\}).$
For courses at maximum capacity, we replace moments $\tilde{m}_{c}$ with the mean rank across the registered students. [^35] We also tried to use the largest student rank and some quantiles of students registered for the course to measure course popularity. These measures did not work very well. Last-minute course registrations are a noisy measure that do not reflect course popularity well. Measures based on quantiles proved to be discontinuous, with a small change in utility parameters often resulting in jumps in the student ranks at a specific quantile and a bad model fit. We keep using the symbol $\tilde{m}_{c}$ for all moments corresponding to vertical components $z$ to avoid clutter in notation.
Overall, there are $756+28+49=833$ moments $m=(\tilde{m},\breve{m},\hat{m})$ and $819$ independent parameters $\pi=(\theta,\Gamma,z)$ to calibrate. Hence, the model is over-identified, allowing us to test for model fit in Section C.5.
C.4 Calibration
We use the Random Serial Dictatorship with adjusted course reserves to calculate simulated moments and find $\pi^{*}$ that minimize a criterion function:
$||m-m^{s}(\pi)||_{W}^{2}=(m-m^{s}(\pi))^{\prime}W(m-m^{s}(\pi)),$ (C5)
where $m$ is a set of moments constructed from the empirical data, $m^{s}(\pi)$ is the average of moments constructed from $100$ simulation rounds for a given vector of parameters $\pi$, and $W$ is a weight matrix. We draw random components of student utility and student choice sets for each simulation round. We use the following procedure in our simulations (see Jalali, Rahmandad and Ghoddusi, 2015):
  • First Stage. We calibrate the optimal parameters $\pi$ using an initial weight matrix $W_{0}$. $W_{0}$ is a diagonal matrix with diagonal element $d$ equal to $1/m^{2}_{d}$, where $m_{d}$ is the value of the moment constructed from the empirical data. Hence, the first stage minimizes the sum of percentage errors for each moment.
    Using estimated $\pi$, we calculate the variance-covariance matrix of simulated moments
    $\hat{S}=\frac{1}{L_{1}}\sum^{L_{1}}_{\ell_{1}=1}\left[m^{\ell_{1}}(\pi,\varepsilon^{\ell_{1}})-\frac{1}{L_{2}}\sum^{L_{2}}_{\ell_{2}=1}m^{\ell_{2}}(\pi,\varepsilon^{\ell_{2}})\right]\left[m^{\ell_{1}}(\pi,\varepsilon^{\ell_{1}})-\frac{1}{L_{2}}\sum^{L_{2}}_{\ell_{2}=1}m^{\ell_{2}}(\pi,\varepsilon^{\ell_{2}})\right],$ (C6)
    where $\varepsilon^{\ell_{1}}$ and $\varepsilon^{\ell_{1}}$ are draws of random component for $\ell_{1}=1,...,L_{1}$ and $\ell_{2}=1,...,L_{2}$. Following best practices, we take $L_{1}=L_{2}=1000$ (see Adda and Cooper, 2003). The optimal weight matrix is the inverse of the variance-covariance matrix $W^{*}=\hat{S}^{-1}.$
  • Second Stage. The second stage of the calibration process is iterative. We start with the optimal weight matrix $W^{*}_{\ell}=W^{*}$ for $\ell=1$. We then use the optimal weight matrix to obtain more precise estimates for parameters $\pi^{*}_{\ell}$. Using the more precise estimate $\pi^{*}_{\ell}$, we compute the optimal weight matrix $W^{*}_{\ell+1}$ and repeat the process $\ell=2,3,...$ until the maximum change of calibrated parameters $|\pi^{*}_{l+1}-\pi^{*}_{l}|$ becomes smaller than $0.01$ to ensure convergence.
Calibration results. Table C.3 provides the calibrated parameters of the student utility model. The top portion of the table displays the horizontal components $\theta$, with rows corresponding to colleges and columns corresponding to year of study. The coefficient $\theta_{ay}$ represents the overall importance of taking courses for students in college $a$ and year $y$. Our normalization ensures that the sum of the coefficients in each row, weighted by the share of the student population across years, equals zero.
We observe that the horizontal components are positive for students in their first and second years and mostly negative in their third and fourth years. This observation suggests that students prefer enrolling in more courses during their initial years of study. The tendency to enroll in fewer than five courses in the later years may be due to several factors. Such students might choose not full-term courses that are not in our data or have already accumulated enough credits for their degree, allowing them to ease their course load in the final years.
The middle portion of the table presents interaction terms of matrix $\Gamma$. Each non-zero interaction coefficient $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$ corresponds to the increase in student’s utility from taking a course in college $a^{\prime}$ (column) compared to a course in student’s college $a$ (row) with the same vertical component. Note that we normalize $\gamma_{aa}=0,a\in{\cal A}$. The general trend of negative utility changes emphasizes a preference for internal college courses, but specific positive changes indicate potential areas for beneficial inter-college course offerings. For example, students from colleges $C$ and $D$ value highly courses in college $A$.
The bottom portion of the table presents quantiles of $756$ coefficients of vertical component $z$, with each coefficient $z_{c}$ being responsible for the average popularity of the course among students from the same college. The values range from a minimum of $-2.32$ to a maximum of $2.28$. This wide range indicates significant variation in the popularity of courses within colleges. Most of the quantile values are negative, up to the $90th$ percentile, indicating that a majority of the courses have negative values. This result is most likely manifested by the student utility model that relies on college-year parameters to identify student preferences. Students randomly draw $80$ courses in their choice set but no more than $5$ courses. Our model, which relies on college-level and year-level parameters, matches this empirical pattern by putting the values of the vertical components of most courses in the negative region.
C.5 Model Fit
We test for model fit using the calibrated parameters. Our student utility model is over-identified, as we have $833$ moments and $819$ independent parameters. The $J$-test statistic is given by
$J=\frac{K}{K+1}(m-m^{s}(\pi^{*}))^{\prime}W^{*}(m-m^{s}(\pi^{*}))\sim\chi^{2}_{14}.$
where $W^{*}$ is the variance-covariance matrix of moments estimated at the terminal step of stage $2$ in the calibration process. This statistic has chi-square distribution with $14$ degrees of freedom under the null hypothesis that the model fits the data. The $J$-test statistic is $J=2.42$, which is less than the critical value of $29.14$ at $99\%$ cut-off level. Hence, the $J$-test does not reject the null hypothesis that the model fits the data at $99\%$ confidence level.
Table C.3: The calibrated parameters of the student utility model.
Coefficients $\theta$
Year 1 Year 2 Year 3 Year 4
A 0.12 0.20 -0.04 -0.27
B 0.13 0.19 0.01 -0.31
C 0.28 0.21 0.01 -0.44
D 0.09 0.19 -0.03 -0.32
E 0.20 0.15 -0.13 -0.29
F 0.17 0.08 -0.09 -0.28
G 0.19 0.11 0.01 -0.37
Coefficients $\gamma$
A B C D E G F
A 0.00 -0.65 -0.58 -0.28 -0.55 -0.70 -0.52
B -0.11 0.00 -0.54 -0.24 -0.09 -0.46 -0.48
C 0.37 -0.22 0.00 -0.01 0.01 -0.28 -0.26
D 0.14 -0.12 -0.39 0.00 -0.16 -0.32 -0.27
E 0.02 -0.55 -0.34 -0.17 0.00 -0.33 -0.40
F -0.07 -0.65 -0.57 -0.21 -0.17 0.00 -0.55
G -0.19 -0.56 -0.58 0.04 -0.20 -0.49 0.00
The Quantiles of Coefficients $z$
Quantile Min 0.10 0.25 0.50 0.75 0.90 Max
$z_{c}$2.32 -1.89 -1.70 -1.47 -1.12 -0.67 2.28
Notes: This table provides information about the calibrated parameters of the student utility model. The upper portion of this table presents the horizontal components $\theta$ of students’ utility, with the rows corresponding to the student’s college and the columns corresponding to the student’s year of study. Our normalization ensures that the sum of the parameters in each row weighted with the share of student population across years equals zero. The middle portion presents interaction terms of matrix $\Gamma$. Each non-zero interaction coefficient $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$ corresponds to the increase in student’s utility from taking a course in college $a^{\prime}$ (column) compared to a course in student’s college $a$ (row) with the same vertical component. Note that we normalize $\gamma_{aa}=0,a\in{\cal A}$. The bottom portion presents some quantiles of vertical components $z$. The interpretation of $z_{c}$ is the average popularity of the course among students within the course’s college.
Appendix D Additional Simulations
This appendix provides additional simulation results. First, for the simulations conducted in the main text, we present statistics on the equilibrium prices in the Pseudo-Market with Priorities (PMP) mechanism (Section D.1) and compare performance with some alternative mechanisms (Section D.2). Second, we supplement the main text with simulation results with a department-first priority structure (Section D.3). Last, we report robustness checks by considering several sizes for student choice sets and the standard deviation of the noise utility parameter (Section D.4).
D.1 The Prices in Pseudo-Market with Priority Equilibria
First, we present statistics on the equilibrium prices in the PMP mechanism for the simulations reported in Section 3.3. The top part of Table D.1 provides information about where the cutoff levels of priority are across courses and how high the prices are for students at the cutoff level of priority. The bottom part provides information about the percentage of students allocated different numbers of courses with positive prices within each year of study.
$70.8\%$ of courses have their cutoff level of priority at $1$. The large share of courses at the lowest priority level reflects the non-binding enrollment of most courses in the university data, where just $11.2\%$ of courses are at or above their maximum enrollment capacity (see Table 2 on p. 2). Above level $1$, there is a noticeable difference in the percentage of courses with odd and even cutoff levels of priority. Substantially more courses have odd cutoff levels ($8.0\%$, $8.5\%$, $5.9\%$ for levels $3$, $5$, and $7$, respectively) than even ones ($2.1\%$, $2.6\%$, $1.7\%$, $0.5\%$ for levels $2$, $4$, $6$, and $8$). The hybrid priority structure we utilize places students at odd priority levels in a course if they are ineligible for any of its reserved seats. The students eligible for a course’s reserved seats typically only comprise a small subset of the interested students. Furthermore, some courses do not reserve any seats. As a result, the cutoff level of priority often ends up at an odd level rather than an even one.
Table D.1: Price statistics for the Pseudo-Market with Priorities mechanism.
Cutoff levels of priority and their average prices
Cutoff level of priority $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
% of courses 70.8% 2.1% 8.0% 2.6% 8.5% 1.7% 5.9% 0.5%
Average cutoff price $0.06$ $0.76$ $0.83$ $0.75$ $0.88$ $0.80$ $0.94$ $0.81$
Courses with positive prices taken by students
# of courses paid for Year 1 Year 2 Year 3 Year 4
0 42.6% 36.7% 44.0% 59.01%
1 45.2% 51.2% 45.4% 37.0%
2 10.9% 10.9% 9.7% 3.8%
3 1.2% 1.1% 0.9% 0.2%
4 0.05% 0.06% 0.03% 0.006%
5 0.001% 0.002% 0.001% 0%
Notes: This table presents statistics on the equilibrium prices in the Pseudo-Market with Priorities mechanism for the simulations reported in Section 3.3. Results are averages across $100$ runs with different random component draws.
Among courses with a cutoff level of priority of $1$, the average price at the cutoff is just $0.06$. As many courses are free for all students, the average cutoff price for these courses is considerably lower than for courses with a higher cutoff level. For higher cutoff levels, the average cutoff prices lie between $0.75$ and $0.94$, with higher prices for odd cutoff levels. Student budgets are assigned in $[1,1.25]$, meaning that, at the average cutoff price, each student can afford one but not two courses for which they are at the cutoff level of priority. Still, students manage to take $k=5$ courses by taking courses for which they are above the cutoff level of priority, where they face a priority-specific price of 00.
The bottom part of Table D.1 further explores the number of courses students pay a positive price for. Across all years of study, at least $36\%$ of students obtain all their course seats for free, and fewer than $13\%$ of students pay a positive price for more than one course. Many courses have a zero price for all students, leading some students to obtain their most preferred course schedule for free. In addition, the average price of a course with a cutoff level of priority above $1$ is more than half of the maximum student budget. So, students often cannot afford a second course with a positive price after obtaining one.
These findings differ across years of study. Less than $7\%$ of cutoff levels of priority are $7$ or $8$, implying that Year $4$ students face zero price for the vast majority of courses. Almost $60\%$ of these students pay for none of their assigned courses, with about $4\%$ obtaining multiple course seats with positive prices. These results are in line with the ones presented in Section 3 that show a larger percentage of Year $4$ students being indifferent between their allocation in the PMP mechanism and the RSD benchmark than students of any other year of study (see Table 5). In earlier years of study, a larger percentage of students pay for course seats. Under $45\%$ of first- through third-year students obtain all their course seats for free, with more than $10\%$ needing to pay for multiple courses in each year.
D.2 Alternative Course Allocation Mechanisms
Our main simulation results compare the Pseudo-Market with Priorities mechanism and two variants of the Deferred Acceptance mechanism, each of which respects the course priority structure. That is, the outcomes of these mechanisms contain no student who wants to enroll in a course that assigns a seat to a student of strictly lower priority.
In turn, the Random Serial Dictatorship with course reserves (RSD) benchmark does not respect any priority structure. The cause of such violations is the presence of course reserves that can be interpreted as seat-specific priorities. Each of a course’s reserved seats has a priority order over students that places reserve-eligible students above others in the same year of study. Unreserved seats only prioritize students by seniority. Other mechanisms in the literature have seat-specific priorities. For example, Hafalir, Yenmez and Yildirim (2013) introduces to the school choice literature a version of the Deferred Acceptance mechanism, the Deferred Acceptance mechanism with minority reserves.
The main objective of this section is to compare the performance of the PMP mechanism with six alternative mechanisms. Four of these mechanisms also have seat-specific priorities (DA-STB and DA-MTB with minority reserves, RSD with university course reserves, Probabilistic Serial with Seniority and Reserves), one does not respect any priorities (A-CEEI), and one respects priorities in a kludgy way (A-CEEI with kludgy pricing).
The Deferred Acceptance (DA) with Minority Reserves with single and multiple tie-breakings. The Deferred Acceptance with minority reserves mechanism was proposed by Hafalir, Yenmez and Yildirim (2013) to address affirmative action policies in school choice. This mechanism is similar to the Deferred Acceptance mechanism, but it allows each school to have a number of reserved seats designated for affirmative action students. These seats give higher priority to minority students up to the point that minorities fill the reserves. We extend this mechanism to a many-to-many matching setting with indifferences. We consider two cases: where indifferences are broken with a single tie-breaking rule and where indifferences are broken with multiple tie-breaking rules, one for each course. [^36] We do not claim that the theoretical properties of Hafalir, Yenmez and Yildirim (2013) extend to many-to-many settings with indifferences. We leave the analysis of this important question for further research.
Random Serial Dictatorship (RSD) with Actual Course Reserves. Course reserves are typically set large at the beginning of the course allocation process at major US universities and relaxed at later stages. In the main text, motivated by this observation, we estimate the optimal reserved seats for each course in the RSD mechanism. Here, we run the same mechanism, using the university’s actual course reserves.
Approximate Competitive Equilibrium from Equal Incomes (A-CEEI). This mechanism was originally proposed by Budish (2011) to allocate courses in business schools. It is a special case of the PMP mechanism in which all courses place all students at the same priority level, leading each course to have a single price. It shares many favorable theoretical properties with the PMP mechanism.
Approximate Competitive Equilibrium from Equal Incomes with Kludgy Pricing (A-CEEI with Kludgy Pricing). This variant of the A-CEEI mechanism aims to satisfy course priorities by using “kludgy pricing”. As in the A-CEEI mechanism, this mechanism uses one price $p_{c}$ per course $c$. However, students receive a price discount depending on their priority. A student $s$ at level of priority $r_{s,c}$ in course $c$ faces price $p_{c}\cdot\left(1-(r_{s,c}-1)/R\right)$, where $R$ is the largest possible priority level. This pricing ensures that students always face a lower price than students at a lower level of priority but leaves open the possibility of a lower-priority student taking a seat when a higher-priority student cannot afford one.
Probabilistic Serial (PS) with Seniority and Reserves. The PS mechanism was proposed by Bogomolnaia and Moulin (2001) to randomly assign objects to agents in settings with single-unit demand for objects and single-unit supply of object types. Kojima (2009) extended the mechanism to settings with multi-unit demand for objects, and Budish et al. (2013) extended the mechanism to settings with multi-unit supply of objects. The undergraduate course allocation setting requires both multi-unit demand and multi-unit supply of courses. The only paper that considers both multi-unit demand and multi-unit supply of courses with preferences over individual courses is an unpublished paper by Pycia (2011). [^37] Nguyen, Peivandi and Vohra (2016) also consider multi-unit demand and multi-unit supply of objects but consider agent preferences over bundles of objects. They propose a version of a probabilistic serial mechanism called the Bundled Probabilistic Serial mechanism that has been successfully implemented at the Technical University of Munich by Bichler and Merting (2021). One of the versions of their probabilistic serial mechanism could also be applied to settings with agent preferences over individual objects. The setting of undergraduate course allocation is further complicated by the presence of student seniorities and course reserves. We define an extension of the probabilistic serial mechanism called Probabilistic Serial with Seniority and Reserves.
In this mechanism, each course seat is regarded as a divisible object of probability shares. We start with the most senior students (e.g., Year 4), each of whom “eats” with speed one the most preferred eligible course seat that is available. Each student is eligible for unreserved seats and each reserved seat specifying her department and year of study. We assume that students first consume reserved course seats before consuming unreserved ones of a particular course. If some reserved/unreserved seats are exhausted, or some student has consumed a full probability share of a certain course’s seat, we stop and redetermine the student’s most preferred eligible course to eat next. This proceeds until time $k$, where the most senior students have consumed up to $k$ courses. We then repeat the procedure with the students of the next seniority level (e.g., Year 3, Year 2, and Year 1). The resulting profile of shares of course seats eaten by agents by time $4k$ corresponds to a random assignment.
Table D.2: Each student’s preferred mechanism by year of study for the alternative mechanisms.
Deferred Acceptance with minority reserves (single) Deferred Acceptance with minority reserves (multiple) Random Serial Dictatorship with actual course reserves
[…]
Year 1 47.3 (1.3) 2.2 (0.7) 50.2 (1.2) 8.4 (1.0) 46.4 (1.5) 28.7 (1.3)
Year 2 23.7 (1.3) 33.6 (1.6) 29.9 (1.3) 36.6 (1.5) 19.1 (1.1) 63.6 (1.3)
Year 3 6.5 (0.8) 42.7 (1.5) 14.0 (1.0) 45.4 (1.6) 4.4 (0.5) 75.5 (1.2)
Year 4 2.0 (0.4) 20.3 (1.5) 12.3 (0.8) 28.1 (1.5) 0.8 (0.3) 53.6 (1.3)
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
[…]
Year 1 88.4 (0.9) 3.4 (0.3) 75.8 (1.5) 4.4 (0.4) 54.2 (1.7) 14.7 (1.3)
Year 2 74.6 (1.2) 13.1 (0.8) 63.7 (1.4) 19.0 (1.1) 31.2 (2.0) 49.6 (1.5)
Year 3 34.6 (1.5) 42.9 (1.5) 30.4 (1.4) 36.5 (1.3) 18.2 (1.6) 57.8 (2.1)
Year 4 15.8 (0.9) 47.3 (1.6) 19.0 (1.0) 20.4 (1.0) 17.2 (1.0) 41.6 (1.4)
Notes: This table reports the percentage of students who prefer a certain mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Results are averages across 100 runs with different random component draws.
Simulation results. Our simulation results for the six alternative mechanisms are presented in Tables D.2D.4. Table D.2, parallel to Table 5 in the main text, reports the percentage of students who prefer the allocation of a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Table D.3, parallel to Table 7, presents information about the change in the standard deviation of student utility by year of study for the six alternative mechanisms compared to the Random Serial Dictatorship with optimal course reserves. Table D.4, parallel to Table 8, presents the shares of students who experience schedule envy towards students of the same priority or lower priority. Table D.5 presents the percentage of students who experience a priority violation. A student experiences a priority violation if she wants to enroll in a course being taken by a student of strictly lower priority. The PMP, DA-STB, and DA-MTB mechanisms studied in the main text have no priority violations.
The DA with minority reserves (with single and multiple tie-breakings) result in a larger percentage of junior students (Years 1 and 2) who prefer these mechanisms ($29.9\%-47.3\%$), and a smaller percentage of senior students (Years 3 and 4) who prefer these mechanisms to the RSD benchmark ($15.6\%-25.1\%$) compared to the PMP, DA-STB, and DA-MTB mechanisms (see Table 5). These results are a direct consequence of the presence of course reserves, which ensures better access to course seats for junior students. Course reserves do not provide equal access to seats, leading to an increase in the standard deviation of student utility for junior students ($6.0\%-14.7\%$) compared to the RSD benchmark. We see similar underperformance in other measures of allocation fairness. In both mechanisms, greater than $20\%$ of students experience schedule envy toward a student of weakly lower priority. Furthermore, $40\%$ of students face a priority violation, indicating that many students can benefit from a course seat assigned to a student of lower priority.
Table D.3: The change in the standard deviation of students’ utilities by year of study for the alternative mechanisms.
Deferred Acceptance with minority reserves (single) Deferred Acceptance with minority reserves (multiple) Random Serial Dictatorship with actual course reserves
Year 1 14.7% 13.3% 13.2%
Year 2 7.1% 6.0% 3.7%
Year 3 -11.6% -13.8% -18.2%
Year 4 -3.4% -5.5% -13.0%
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
Year 1 -10.4% -11.4% 9.4%
Year 2 -12.5% -14.5% -0.8%
Year 3 -20.8% -18.4% -22.0%
Year 4 -22.5% -8.5% -14.1%
Notes: The change in the standard deviation of student utility by year of study for the six alternative mechanisms compared to the Random Serial Dictatorship with optimal course reserves. The results are based on $100$ runs with different random component draws.
The RSD with actual course reserves mechanism performs quite poorly compared to the RSD benchmark. For second- through fourth-year students, more than $50\%$ of students prefer the outcome of the RSD mechanism with optimal course reserves over the actual course reserves. The mechanism underperforms in terms of allocation fairness as well, with $27.8\%$ of students preferring the schedule of a student with weakly lower priority in each course and $59\%$ of students with priority violations. Only first-year students benefit from using the university’s actual course reserves. These results justify why we consider the RSD mechanism with optimal course reserves as the benchmark mechanism rather than the RSD mechanism with actual course reserves. Using the actual course reserves presents an unfair comparison, as course reserves are typically set large at the beginning of the process and relaxed later, resulting in the allocation of courses to students that, to a large extent, is determined by an unstructured decentralized process that happens after the main allocation round in practice.
Table D.4: The percentage of students who experience schedule envy for the alternative mechanisms.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
DA with minority reserves (single) 78.9 (0.5) 18.0 (0.5) 2.7 (0.2) 0.4 (0.2) 0.02 (0.03) 0.0007 (0.01)
DA with minority reserves (multiple) 79.5 (0.5) 18.9 (0.6) 1.6 (0.2) 0.1 (0.03) 0.003 (0.01) 0.0 (0.0)
RSD with actual course reserves 72.2 (0.5) 23.9 (0.5) 3.6 (0.2) 0.3 (0.1) 0.01 (0.02) 0.0 (0.0)
A-CEEI 95.1 (0.4) 4.9 (0.4) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
A-CEEI with kludgy pricing 96.3 (0.3) 3.7 (0.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Probabilistic Serial with reserves 89.1 (0.6) 9.9 (0.6) 0.9 (0.2) 0.05 (0.03) 0.0007 (0.003) 0.0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the alternative mechanisms. The first column provides information about students who experience no envy. The other columns show the percentage of students who experience schedule envy bounded by $1,...,5$ courses. Results are averages across $100$ runs with different random component draws.
The A-CEEI and A-CEEI with Kludgy Pricing mechanisms do not strictly enforce course priorities through prices, unlike the PMP mechanism. These relaxed restrictions result in more course seats available to junior students (Year 1 and 2), who overwhelmingly strictly prefer each mechanism to the RSD benchmark ($63.7\%-88.4\%$). The results reverse for senior students (Years 3 and 4), with more students favoring the benchmark. Both mechanisms reduce the standard deviation of student utility across all years of study ($8.5\%-22.5\%$). This decrease is substantially larger than those observed in Section 3 for the PMP mechanism, suggesting that the relaxed priority constraints allow students to receive more equal access to courses. Similarly, over $95\%$ of students experience no envy toward a student of weakly lower priority, with schedule envy bounded by a single course in each mechanism. However, without these constraints, priority violations are common in both mechanisms, with $61.1\%$ of students in the A-CEEI mechanism and $44.6\%$ of students in the A-CEEI with Kludgy Pricing mechanism whose schedule would benefit from a course seat taken by a student of strictly lower priority. Also, the average timing to calculate the allocation of the A-CEEI and A-CEEI with Kludgy Pricing mechanisms (2003 and 1894 seconds, respectively) is almost twice as much as that of the PMP mechanism (1112 seconds).
Table D.5: The percentage of students with a priority violation for the alternative mechanisms.
DA with minority reserves (single) DA with minority reserves (multiple) RSD with actual reserves A-CEEI A-CEEI with kludgy pricing PS w. seniority and reserves
46.1% 47.3% 59.0% 61.1% 44.6% 81.0%
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority for each of the alternative mechanisms. Results are averages across $100$ runs with different random component draws.
The Probabilistic Serial with Seniority and Reserves mechanism is the only one of mechanisms we consider that produces a random assignment of course seats to students. This approach is more flexible, allowing the assignment of probability shares of course seats to students rather than full seats. The mechanism, however, delivers more beneficial outcomes only to students of Year $1$ (see Table D.2). More second- through fourth-year students strictly prefer their schedule in the RSD benchmark. This result is in contrast to the PMP mechanism, where the percentage of students who prefer the PMP mechanism is larger than the percentage of students who prefer the RSD benchmark for second- through fourth-year students. In terms of allocation fairness, the probabilistic serial mechanism reduces the standard deviation of students’ utilities for all years of study and results in just $10.9\%$ of students with envy towards a student of weakly lower priority. [^38] Kojima (2009) and Budish et al. (2013) establish that the probabilistic serial mechanism in multi-unit demand setting and multi-unit supply setting, respectively, results in no envy. This result does not extend to settings where agents demand multiple units of objects and objects are supplied in multiple units simultenously. However, in doing so, the mechanism results in a large percentage of students who experience priority violation ($84.6\%$). This can be explained, in part, by the flexibility of the mechanism. We count a student as experiencing a priority violation if there is a positive probability share of a course that she prefers to some assigned probability share of hers (possibly empty) that is instead assigned to some student at a strictly lower priority level. This matches the definition of ex-ante stability proposed in Kesten and Ünver (2015). We also consider envy in terms of ex-ante assignment, not an ex-post assignment, as for the other alternative mechanisms. Hence, the results are not directly comparable.
D.3 Department-First Priorities
Some U.S. universities consider department-specific priorities as more important than those based on year of study. For example, Dartmouth College explicitly states these priorities on its website. [^39] See https://www.dartmouth.edu/reg/registration/course_priorities.html. Here, we present additional simulation results comparing the performance of all considered mechanisms under a department-first priority structure, where the department-specific priority takes precedence over the year-specific priority. We modify Example 2 from Section 3 to provide an illustration.
Example D1.
The course reservation for Real Analysis in Table 4 results in the following department-first priority structure for the course:
  • $r_{s,c}=8$: 4th-year students in Dept 2
  • $r_{s,c}=7$: 3rd-year students in Dept 2
  • $r_{s,c}=6$: $\varnothing$
  • $r_{s,c}=5$: $\varnothing$
  • $r_{s,c}=4$: 4th-year students in all departments except Dept 2
  • $r_{s,c}=3$: 3rd-year students in all departments except Dept 2
  • $r_{s,c}=2$: 2nd-year students in all departments
  • $r_{s,c}=1$: 1st-year students in all departments
With a department-first priority structure, the Deferred Acceptance mechanism coincides with the Deferred Acceptance mechanism with minority reserves of Hafalir, Yenmez and Yildirim (2013). As in Example D1, this is because students eligible for reserved seats receive the highest priority for all course seats. Hence, we only consider the Deferred Acceptance mechanisms with single and multiple tie-breakings in our simulations. We also drop the analysis of Random Serial Dictatorship with actual course reserves, as its performance is subpar compared to all the other mechanisms. The results of our simulations are presented in Tables D.6D.8.
Table D.6: Each student’s preferred mechanism by year of study under department-first priority structure.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
[…]
Year 1 14.8 (2.1) 5.5 (1.4) 11.4 (1.5) 1.2 (0.6) 19.4 (1.5) 10.5 (1.1)
Year 2 16.3 (1.7) 9.4 (1.1) 9.3 (1.3) 4.1 (1.2) 20.1 (1.2) 16.3 (1.3)
Year 3 19.5 (2.4) 13.3 (1.3) 8.1 (1.5) 5.6 (1.3) 21.9 (1.4) 20.1 (1.4)
Year 4 14.7 (1.4) 13.2 (1.5) 0.5 (1.0) 6.2 (1.3) 18.4 (1.0) 18.6 (1.3)
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with Seniority and Reserves
[…]
Year 1 71.1 (1.1) 12.8 (0.7) 62.9 (1.5) 10.0 (0.6) 24.3 (2.3) 25.0 (2.4)
Year 2 73.8 (0.9) 13.8 (0.7) 66.4 (1.3) 11.5 (1.0) 20.6 (2.9) 44.3 (1.6)
Year 3 56.8 (1.2) 24.0 (1.2) 51.4 (1.4) 19.5 (1.0) 33.9 (1.8) 35.9 (1.7)
Year 4 27.9 (1.0) 35.6 (1.3) 26.2 (1.3) 26.7 (1.3) 30.1 (1.1) 29.4 (1.2)
Notes: This table reports the percentage of students who strictly prefer a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who strictly prefer the benchmark under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
Table D.6 exhibits each students’ preferred mechanism by year of study under the depart- ment-first priority structure. As opposed to the hybrid priority structure, all students eligible for reserved seats are prioritized above more senior but ineligible students. As a result, a larger percentage of junior students (Years 1 and 2) prefer the PMP, DA-STB, DA-MTB, and A-CEEI with kludgy pricing mechanisms over the RSD benchmark. The comparison is the opposite for senior students (Years $3$ and $4$), with the DA-STB and DA-MTB mechanisms underperforming the RSD benchmark for Year $4$ students. Hence, switching from the RSD benchmark to one of the alternative mechanisms would lead to the redistribution of student welfare from senior to junior students.
Table D.7: The change in the standard deviation of students’ utilities by year of study under a department-first priority structure.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breaking
Year 1 -0.4% 1.1% -0.02%
Year 2 -0.6% 0.5% -0.1%
Year 3 -2.9% 0.1% -1.7%
Year 4 -2.9% -0.1% -1.5%
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
Year 1 -24.4% -21.4% -4.5%
Year 2 -18.9% -16.2% -7.3%
Year 3 -8.1% -8.6% -9.3%
Year 4 -17.6% -11.7% -9.2%
Notes: The change in the standard deviation of student utility by year of study for the considered mechanisms compared to the Random Serial Dictatorship with optimal course reserves. The outcomes of the Deferred Acceptance mechanism with minority reserves coincides with Deferred Acceptance mechanism both for single and multiple tie-breakings. Results are averages across $100$ runs with different random component draws.
Table D.7 reports each mechanism’s change in the standard deviation of students’ utilities relative to the RSD benchmark. The largest improvements in standard deviation of students’ utilities happen for the A-CEEI, A-CEEI with kludgy pricing, and Probabilistic Serial with seniority and reserves mechanisms. Table D.9 shows that this happens at the expense of priority violations, which more than $50\%$ of students experience in each of these three mechanisms. As in the hybrid priority results, this is a larger percentage of priority violations than we observe in the benchmark ($32.5\%$). In turn, the PMP, DA-STB, and DA-MTB mechanisms have no priority violations, but perform similarly to the RSD benchmark in terms of the standard deviation of students’ utilities.
Table D.8 shows that the percentage of students experiencing schedule envy in each mechanism is similar in comparison to the hybrid priority structure (see Tables 8 and D.4). The PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms deliver schedule envy bounded by a single course toward students of weakly lower priority, with only $8.2\%$, $4.9\%$, and $4.1\%$ experiencing any envy, respectively. The DA-STB and DA-MTB mechanisms each result in about $13\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses in the DA-STB mechanism and under $1\%$ in DA-MTB. The RSD with optimal course reserves mechanism has $14.5\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses. Just $1\%$ of students experience schedule envy in the random assignment of the Probabilistic Serial with seniority and reserves mechanism.
Table D.8: The percentage of students who experience schedule envy under department-first priority structure.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
Pseudo-Market with Priorities 91.8 (0.8) 8.2 (0.8) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
RSD with optimal course reserves 85.5 (0.6) 12.4 (0.6) 1.8 (0.2) 0.3 (0.1) 0.02 (0.02) 0.0004 (0.002)
DA with single tie-breaking 86.8 (0.6) 11.3 (0.6) 1.6 (0.2) 0.3 (0.2) 0.01 (0.02) 0.0007 (0.007)
DA with multiple tie-breaking 87.7 (0.6) 11.9 (0.6) 0.4 (0.1) 0.002 (0.005) 0.0 (0.0) 0.0 (0.0)
A-CEEI 95.1 (0.4) 4.9 (0.4) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
A-CEEI with kludgy pricing 95.9 (0.3) 4.1 (0.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Probabilistic Serial with seniority and reserves 99.0 (0.2) 1.0 (0.2) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course under a department-first priority structure. The first column provides information about students who experience no envy. The other columns show the percentage of students who experience schedule envy bounded by $1,...,5$ courses. Results are averages across $100$ runs with different random component draws.
Table D.9: The percentage of students with a priority violation under a department-first priority structure.
PMP RSD with opt. reserves DA with single tie-breaking DA with mult. tie-breaking A-CEEI A-CEEI with kludgy pricing PS w. seniority and reserves
0% 32.5% 0% 0% 64.1% 53.2% 84.8%
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority under a department-first priority structure. Results are averages across $100$ runs with different random component draws.
D.4 Robustness: Student Choice Sets and Noise Term
This section performs some additional robustness checks. We take the calibrated student utility model calibrated and consider changes in two parameters: the size of student choice sets ($60$, $70$, $80$, $90$), and the standard deviation of the random component of student utility ($\sigma=0.75,1,1.25,1.5$). Figure D.1 reports the performance of the PMP mechanism compared to the RSD benchmark. The graphs show the percentages of students who are indifferent (solid line), prefer the PMP mechanism (dashed line), and prefer the RSD benchmark (dotted line). In contrast to the main text, the reported simulation results are based on $50$ random component draws.
The top four graphs show comparative statics results for the size of student choice sets. The percentage of students indifferent between the PMP and RSD mechanisms is relatively constant across all sizes ($74.8\%$ to $76.3\%$), slightly lower for student choice sets of size $90$ ($68.5\%$). A relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the choice set.
The bottom four graphs show comparative statics results for the standard deviation of the noisy component $\sigma$. The percentage of students indifferent between the PMP mechanism and RSD mechanism significantly declines from $\sigma=0.75$ to $\sigma=1.5$, with the largest decrease for Year 1 students ($85.7\%$ to $13.8\%$). Again, a relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the standard deviation. The benefits of the PMP mechanism are also larger for more noisy environments, where student preferences for the same schedules could be quite different.
Overall, the PMP mechanism outperforms the RSD benchmark for all considered values of the size of student choice sets and the standard deviation of the random component of student utility. Our results are robust in response to the changes in these parameters.
Figure D.1: Additional Simulations: Student Choice Sets and Noise Term.
Notes: The performance of Pseudo-Market with Priorities (PMP) compared to the Random Serial Dictatorship (RSD) with optimal course reserves. The top four figures present the percentages of students who are Indifferent (solid line), Prefer (dashed line), and Do Not Prefer (dotted line) the allocation of PMP to the RSD benchmark for different sizes of student choice sets: $60$, $70$, $80$, and $90$. The bottom four figures present the same comparison for the standard deviation of student utility $\sigma=0.75,1,1.25,1.5$. Results are averages across $50$ runs with different random component draws.
Appendix E Supplementary Materials
The supplementary materials are organized as follows. Section E.1 describes the Mathematica programs for calibrating the student utility model in Section C and performing the simulations comparing course allocation mechanisms in Section 3.3 and Sections D.1D.4. Section E.2 provides details about the algorithm developed to quickly find the Pseudo-Market Equilibrium with Priorities for large student populations.
E.1 Program
Most simulations have been done using Mathematica 14.0.0.0 on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GB of RAM, 128 GB of SSD hard drive space, and a Windows Server 2019 Standard operating system. There are three simulation programs:
  • File MSM-utility-model.nb implements the simulated method of moments to calibrate the parameters of the student utility model.
  • File simulations.nb determines the outcomes and compares the performance of the four mechanisms described in Section 3.2.
  • File simulations-appendix.nb determines the outcomes and compares the performance of the six alternative mechanisms described in Sections D.2.
The program MSM-utility-model.nb has the following sections:
  • Functions, Data, and Parameters sets the main parameters, loads the data from files, and defines some essential functions.
    • ShortDataCreation creates a version of the data with a small number of students for testing purposes.
    • Utility generates and returns the matrix of student utilities over courses for a given set of parameters $\theta$, $\gamma$, and $z$.
    • RSDallocation finds the outcome of the Random Serial Dictatorship with course reserves for a given matrix of student utilities, tie-breaking, and course reserves.
  • Simulated Method of Moments sets the main parameters of the method of simulation moments and defines all essential functions of the calibration process.
    • EmpiricalMoments produces the empirical moments from the data.
    • W1 and Weight calculate the moment weight matrices for the first and second stages of the calibration process (see Section C.4).
    • RandomParam generates the random student choice sets and random utility components.
    • SimulatedMoments and MomentSelection call RSDallocation to produce simulated allocations and select the relevant simulated moments.
    • MSMObjFn produces the weighted squared distance between the simulated moments and the empirical moments.
    • Optimization runs a process to optimize the student utility parameters with an initial set of parameters and other adjustable attributes.
    • ParameterNormalization implements the normalization of student utility parameters discussed in Section C.1.
    • SimulatedMoments WeightCalibration is an analog of SimulatedMoments function for the optimal weight calculation.
    • CalibrationTest tests whether the utility parameters are well-calibrated.
  • Run calls the Optimization function, tracks the progress, and records the calibrated parameters of the utility model into file “INPUTP.xlsx”. The section also performs a $J$-test on model fit for Section C.5.
The programs simulations.nb and simulations-appendix.nb have the same sections and employ similar functions. We label the functions only used in simulations-appendix.nb with an asterisk.
  • Functions, Data, and Parameters sets the main parameters, loads the data from files, and defines some essential functions.
    • ShortDataCreation, RandomParam, Utility, and RSDallocation are the same as the functions described above.
    • DAmechanism finds the outcome of the Deferred Acceptance mechanism for given student utilities and a given priority order.
    • OptSetAsideFunction generates the optimal set of course reserves using the Deferred Acceptance Mechanism with single tie-breaking.
  • Allocation Generation Functions defines functions that produce the allocations for the mechanisms in Section 3 and D.2.
    • PMPallocation and PMPallocationAPPENDIX* find the outcomes of the Pseudo-Market with Priorities mechanism. They employ some supplementary functions, which are described below.
    • PMPInitialPrice and PMPInitialPriceAPPENDIX* provide an educated guess for the initial price in the search for a PMP equilibrium.
    • PMPPhaseI and PMPPhaseIAPPENDIX* implement Phase I of the algorithm to find a Pseudo-Market Equilibrium with Priorities.
    • PMPPhaseII and PMPPhaseIIAPPENDIX* implement Phase II of the algorithm to find a Pseudo-Market Equilibrium with Priorities.
    • OptAllocation and OptAllocationAPPENDIX* find each student’s demand for courses at a given set of prices.
    • OptAllocationInd and OptAllocationIndAPPENDIX* solve the optimization problem for an individual student.
    • OptAllocationFast and OptAllocationFastAPPENDIX* find each student’s demand for courses at a given set of prices using the information about prices and demand at the previous step.
    • Error calculates the market-clearing error for an allocation at a given set of prices.
    • FinalAllocation, FinalAllocationAPPENDIX*, OptAllocationFinalFeasibilityCheck, and OptAllocationFinalOptimizationCheck provide additional checks that the optimization problem for each student is correctly solved for the final allocation and prices.
    • DASTBallocation and DAMTBallocation find the outcomes of the Deferred Acceptance mechanism with single and multiple tie-breakings, respectively.
    • ACEEIallocation* and Kludgyallocation* find the outcomes of the Approximate Competitive Equilibrium from Equal Incomes (A-CEEI) and the A-CEEI with kludgy pricing mechanisms, respectively.
    • DAreservesmechanism* finds the outcome of the Deferred Acceptance with minority reserves for given student utilities, priority order, and course reserves.
    • DASTBreserveallocation* and DAMTBreserveallocation* find the outcomes of Deferred Acceptance mechanism with minority reserves mechanisms with single and multiple tie-breakings, respectively.
    • PSseniorityreservesallocation* finds the outcome of the Probabilistic Serial Mechanism with Seniority and Reserves mechanism.
  • Simulation Run is the main section of the program. It generates the allocations for each mechanism across multiple simulation rounds. The allocations are stored in the files allocations-data.txt and appendix-data.txt*. Data on course prices and excess demand for the PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms are recorded in files pmp-data.txt, aceei-data.txt*, and kludgy-data.txt*, respectively.
  • Results Setup loads the allocation data from allocation-data.txt for simulations.nb and from appendix-data.txt* for simulations-appendix.nb.
  • Preferred Mechanism Results produces results for Tables 5, D.2, and D.6.
  • Envy Results produces results for Tables 8, D.4, and D.8.
  • Priority Violations produces results for Tables D.5 and D.9.
  • Utility Results loads in pricing data from pmp-data.txt, aceei-data.txt*, and kludgy-data.txt* and produces results for Tables 6, 7, D.1, D.3, and D.7.
E.2 Computational Algorithm
We developed an algorithm to find a Pseudo-Market Equilibrium with Priorities, taking in student utilities $v$, student priorities $priorities$, and a tie-breaking order over students $studentrank$ and outputting an allocation $bestdemand$ and a vector of course prices $bestt$. Its main stages are described below, with the pseudocode shown on page Algorithm.
Initialization. We first allocate budgets of fake money to students. To do so, we reorder the set of student budgets $b$, evenly spread values from $1$ to $1.251+\beta=$ 1.25, according to $studentrank$. The function PMPInitialPrice takes in student utilities $v$ and provides an initial estimate for the price vector based on students’ ideal bundles of courses initdemand. [^40] A student’s ideal bundle is the set of courses they would demand if all course prices were zero. Rather than working in $M\!R$-dimensional space, we leverage the approach in the proof of Theorem 1 and parameterize priority-specific prices as $t\in[0,R\overline{b}]^{M}$ We collect the main parameters of the algorithm in $pmpparam\leftarrow\{v,b,priorities,initdemand\}$ and initialize the variables $boundcheck\leftarrow 0$ and $oversubcheck\leftarrow 0$. These last two parameters serve as exit flags for Phases I and II of the algorithm described below.
[…]
The main innovation of our search process is that course demand is only updated for two subsets of students at each step. The first subset includes students who can no longer afford the allocation determined in the previous step. The second includes students who did not demand their ideal bundle in the previous step and have a strictly positive utility for courses with recently decreased prices, where these prices are within the student’s budget. For any student outside these two subsets, their optimal course demand remains unchanged between consecutive steps of the search process. This innovation has proven particularly effective for large student populations, significantly reducing the time required for the search process compared to having all students update their demand at each step.
Phase II. The second phase starts with the best price of Phase I and works to reduce the excess course demand to levels below those observed in the data. For each course with positive excess demand and each student who demands it, we use a dual program to calculate the minimum price change needed so that the student no longer demands the course. We then sort these price changes and apply the smallest price change that fully eliminates the excess demand for the course. For example, if a course’s excess demand is 2, we apply the second smallest price change among students who demand the course. We repeat this process until two conditions are met: (i) the distribution of excess demand errors stochastically dominates the distribution observed in the data, and (ii) no course is oversubscribed by more than $k-1=4$ seats. If both (i) and (ii) are satisfied we set $oversubcheck=1$. We also update the value of $boundcheck$ as the market-clearing error changes. The output of Phase II includes the price vector $bestt$ and the exit flags $boundcheck$ and $oversubcheck$.
Final Allocation. After Phase II, we provide checks that the optimization problem for each student at the final prices is correctly solved and produce the final course allocation. If errors occur, the FinalAllocation function will correct them. If errors result in a violation of Phase I and II conditions, we return $bestdemand=\emptyset$. If such an event happens, we reset the exit flags and the number of rounds to the maximum value and restart the process with a new initial price vector. The last step is done using the last IF statement.
Last, we describe the two while loops in the algorithm. The outer while loop continues as long as $boundcheck=0$ or $oversubcheck=0$, allowing the search for a price vector to proceed. At each step, we add a small amount of noise to the initial price vector to begin the search from a slightly different starting point, and we reset the number of rounds in the inner loop to zero. The inner loop has the same exit conditions, with an additional condition on the number of rounds. If either $boundcheck=0$ or $oversubcheck=0$ after Phase II, we repeat the process. We allow up to $maxrounds=6$ iterations of Phase I and II within the inner loop making the search process more persistent during additional rounds. If the algorithm fails to find a solution after $maxrounds$ iterations, we restart the process with a new initial price vector. Simulations indicate that typically only 1-4 iterations occur between Phases I and II in the inner loop. We did not encounter cases requiring a restart through the outer loop for the main simulations. The simulations in the appendix Section D.4 checking the robustness of our results do encounter some challenging cases where a restart is necessary.
Algorithm Pseudo-Market with Priorities
1:student utilities $v$, student priorities $priorities$, student ranking $studentrank$
2:course allocation $bestdemand$, prices $bestt$
3:Initialization: $b\leftarrow b(studentrank)$, $\{t_{0},initdemand\}\leftarrow\textbf{PMPInitialPrice}(v)$,
4:$pmpparam\leftarrow\{v,b,priorities,initdemand\}$, $boundcheck\leftarrow 0$, $oversubcheck\leftarrow 0$
5:while $boundcheck=0$ OR $oversubcheck=0$ do
6:   $t\leftarrow t_{0}+noise$, $nrounds\leftarrow 1$
7:   while ($boundcheck=0$ OR $oversubcheck=0$) AND $nrounds\leq maxrounds$ do
8:      $\{bestt,boundcheck\}\leftarrow\textbf{PMPPhaseI}(t,pmpparam,nround)$
9:      $\{bestt,boundcheck,oversubcheck\}\leftarrow\textbf{PMPPhaseII}(bestt,pmpparam)$
10:      $\{bestdemand\}\leftarrow\textbf{FinalAllocation}(bestt,pmpparam)$
11:      if $bestdemand==\emptyset$ then
12:         $boundcheck\leftarrow 0,oversubcheck\leftarrow 0,nround\leftarrow maxrounds$
13:      end if
14:      $t\leftarrow bestt$, $nround\leftarrow nround+1$
15:   end while
16:end while