### v2

# Undergraduate Course Allocation

Abstract
Prevailing methods of course allocation at undergraduate institutions involve reserving seats to give priority to designated groups of students. We introduce a competitive equilibrium-based mechanism that assigns course seats using student preferences and course priorities. This mechanism satisfies approximate notions of stability, efficiency, envy-freeness, and strategy-proofness. We evaluate its performance relative to a mechanism widely used in practice using preferences estimated from university data. Our empirical findings demonstrate an improvement in student satisfaction and allocation fairness. The number of students who envy another student of weakly lower priority declines by 8 percent, or roughly 500 students.
JEL classification: D47, D63, D82, C63, C78, I21
Keywords: market design, many-to-many matching, course allocation, approximate competitive equilibrium, random serial dictatorship, reserves, priorities
Every academic term, over 6,500 undergraduate institutions across the United States assign course schedules to a total of nearly twenty million students. [^1] Based on National Center for Education Statistics, 2017-2018. Based on course timing and prerequisites, the set of schedules a student can be assigned in a given term is limited. Moreover, student preferences over possible schedules are heterogeneous. In the face of room size and teaching constraints, universities use factors such as seniority and student department to complete the challenging task of allocating seats in overdemanded courses.
Course allocation is a well-studied problem in the context of graduate business schools. Many institutions employ course bidding systems that require strategic decision-making from students, often leading to inefficient allocations (Sönmez and Ünver, 2010). Recently, several business schools replaced course bidding systems with the approximate competitive equilibrium from equal incomes (A-CEEI) mechanism (Budish, 2011; Budish et al., 2023), which seeks to alleviate these issues. This mechanism elicits student preferences over course schedules, assigns students almost-equal budgets of fake money, and internally determines an approximate competitive equilibrium, in which each student receives her most-preferred affordable course schedule subject to course prices. The competitive equilibrium is approximate due to a small market-clearing error, measured by the assignment of a course’s seats over and under its capacity. This mechanism treats all business school students equally, as students are randomly assigned budgets and face the same course prices.
At undergraduate institutions, students are not treated equally, making course allocation a two-sided matching problem. Each course divides students into levels of priority based on student characteristics like year of study and department. In commonly-used course allocation mechanisms, students select courses in order of seniority (with ties broken randomly) and priorities based on departments are enforced by reserving seats in each course. Top US universities that closely follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. [^2] In general, there is some heterogeneity in how universities elect to assign seats to students. Some universities use priorities (e.g., Dartmouth) or reserves (e.g., Vanderbilt) to assign seats, whereas others delegate the decision to departments (e.g., Cornell). Some use two phases (e.g., UCLA), only allowing students to enroll in a limited number of courses in the first phase. Seniority can be based on academic year (e.g., Carnegie Mellon), number of earned credits (e.g., University of North Carolina), or time to graduation (e.g., Washington University in St. Louis). At some universities, few classes are over-enrolled (e.g., Caltech), whereas at others, the process is rather congested (e.g., UC Berkeley). In most cases, courses have a small number of priority levels, rather than one (Budish, 2011) or complete strict priorities Roth (1984).
In this paper, we approach the task of allocating course schedules to students in the context of a many-to-many two-sided matching problem with heterogeneous student preferences and weak course priorities. We propose a novel deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that uses fake money and competitive equilibrium to allocate course seats to students without transfers. This mechanism elicits student preferences over course schedules, assigns almost-equal budgets to students, and finds an approximate competitive equilibrium allocation under priority-specific prices. These prices respect course priorities by setting a cutoff priority level for each course, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face some price, and lower priority students cannot afford a seat.
The A-CEEI mechanism is the special case of our mechanism in which no course prioritizes any student over any others. Our first result shows that, with arbitrary course priorities, the PMP mechanism maintains the same, small worst-case bound on the market-clearing error found in Budish (2011) for the A-CEEI mechanism. This bound is increasing in the dimension of the space of prices, which equals the number of courses in settings without priorities. Our result holds despite the expansion from one price per course to priority-specific prices in each course. The key insight of the proof is the introduction of a parameterization within the priority-specific price space. The parameterization simplifies the space of prices to a single value per course, enabling a simultaneous search for both the cutoff priority level and the corresponding price. This strategy simplifies the complexity of the problem and allows for the application of some previously established results.
The PMP mechanism has several advantageous theoretical properties. Adjusting course capacities to accommodate over-enrolled courses, the PMP mechanism’s outcomes ensure that no student can be assigned a preferred schedule without violating course priorities or capacity constraints (approximate stability). It is also not possible to reassign course seats among a group of students, benefitting some members and hurting none, while ensuring that the outcome respects course priorities to the same degree (approximate priority-constrained efficiency). Among similarly prioritized students, its outcome limits envy; if a student prefers the schedule assigned to a student with the same or a lower priority level in each course, by removing one course from the preferred schedule, the student now prefers her own (schedule envy bounded by a single course). Finally, in large populations, the PMP mechanism is resistant to strategic manipulations by students (strategy-proof in the large).
To better guide practitioners in the course allocation process, we compare the performance of the PMP mechanism with several alternative mechanisms. As a benchmark, we consider a mechanism, the Random Serial Dictatorship with course reserves (RSD) mechanism, which closely resembles what is used in practice. Students select courses in order of seniority (with ties broken randomly) and priorities based on student department and year of study are enforced by reserving seats in each course. [^3] In practice, at some universities, courses set a large number of reserved seats at the beginning of the registration process. After students initially register, university departments relax the course reserve constraints and admit students from waiting lists to ensure full enrollment. We do not model this later period of the registration process. Instead, we estimate the optimal course reserves that a university registrar should set at the beginning of the process if the number of reserved seats cannot be changed. We also consider the Deferred Acceptance mechanism with a single and with multiple tie-breakings (denoted DA-STB and DA-MTB, respectively). These mechanisms allocate courses to students by extending the Gale-Shapley algorithm (Gale and Shapley, 1962) to many-to-many matching, where priority ties are resolved by either a single tie-breaking rule or course-specific tie-breakings. The Deferred Acceptance mechanism has been successful in many practical applications, including matching applicants to residency programs, students to schools, and workers to job positions (see Roth, 2018). We analyze six additional mechanisms in Appendix D.
We use university data on the final allocation of 6 0236\,023 students to $756$ courses from the Spring 2018 term at a private institution in the mid-Atlantic region. In total, 23 36923\,369 course seats are occupied, with $11.2\%$ of courses enrolled at or above their reported capacities. We use student and course characteristics to estimate a model of student utilities and evaluate the performance of the PMP, RSD, DA-STB, and DA-MTB mechanisms with a variety of measures of student satisfaction and allocation fairness.
Our results demonstrate that the PMP mechanism delivers a substantially fairer outcome than the benchmark RSD mechanism and the two DA mechanisms. Considerably fewer students - $8.1\%$ of the total population, or almost $500$ students - prefer the course schedule assigned to a student of weakly lower priority in each course. In the PMP mechanism, no student’s envy is greater than a single course, as opposed to $3.3\%$, $2.5\%$, and $0.7\%$ of students in the RSD, DA-STB, and DA-MTB mechanisms, respectively. While the PMP, DA-STB, and DA-MTB mechanisms prevent lower-priority students from taking a seat that a higher-priority student desires, $16.2\%$ of students would benefit from taking a seat that is assigned to a student with strictly lower priority in the RSD mechanism.
These fairness improvements come without sacrificing student satisfaction. Only $31\%$ of students change schedules between the PMP and RSD mechanisms. Among these students, $300$ more students strictly prefer their outcome in the PMP mechanism. In contrast, both DA mechanisms result in approximately the same number of students who strictly prefer the mechanism and who strictly prefer the RSD benchmark. The main difference between the two DA mechanisms is that only $11.3\%$ of students change their schedules for DA-STB, while $40.3\%$ do so for DA-MTB. The PMP mechanism also improves on the RSD benchmark in mean utility for students all years of study, unlike both DA mechanisms.
Literature review. This paper contributes to the literature analyzing the allocation of objects to agents through competitive equilibria. As competitive equilibrium from equal incomes are not guaranteed to exist, many papers focus on random allocation mechanisms. Hylland and Zeckhauser (1979) first proposed the use of fake money and competitive equilibrium to randomly allocate objects to agents through what are referred to as pseudo-market mechanisms. He et al. (2018) incorporate a priority structure into pseudo-markets and analyze random mechanisms with a focus on unit-demand settings such as school choice. [^4] He et al. (2018) explain how their results extend in many-to-many settings with additive utility. Miralles and Pycia (2021), Echenique, Miralles and Zhang (2021) and Nguyen, Nguyen and Teytelboym (2021) consider pseudo-market solutions to random allocation problems with complex constraints. Pycia (2021) provides an excellent survey of the use of pseudo-markets for random allocations in environments without transfers.
In contrast to these papers, we only consider deterministic mechanisms and study a many-to-many matching problem. Budish (2011) introduces the idea of competitive markets that may only approximately clear. He proposes a deterministic mechanism, the approximate competitive equilibrium from equal incomes mechanism, which is approximately Pareto-efficient, strategy-proof in the large, and bounds student envy by a single course. This mechanism was successfully implemented at Wharton Business School and Columbia Business School (see Budish et al., 2017; Budish and Kessler, 2022). [^5] There are few other studies of real-world course allocation mechanisms. Budish and Cantillon (2012) use field data to study a non-strategy proof course allocation mechanism at Harvard Business School. Bichler and Merting (2021) compare the performance of two random mechanisms at the Technical University of Munich. Rusznák, Biró and Fleiner (2021) analyze course allocation at a large university in Hungary. A key contribution of the present study is extending Budish’s approach to two-sided matching settings with weak course priorities, a central feature of the undergraduate course allocation problem.
Two recent papers also analyze deterministic assignments in many-to-many matching settings. Nguyen and Vohra (2022) establish the existence of a competitive equilibrium when all agent preferences satisfy a novel geometric substitutes property, which is a strict generalization of the gross substitutes property (see Kelso and Crawford, 1982). When there is an upper limit on the number of goods $k$ an agent can acquire, they prove that there always exists an approximate competitive equilibrium with no good over-enrolled by more than k1k-1 units. This “good-by-good” bound dominates the aggregate market-clearing bound of Budish (2011) when preferences are close substitutes, whereas the aggregate market-clearing bound dominates when student preferences are close complements. Though our main theoretical results follow the aggregate market-clearing bound approach, we ensure in our simulations that each approximate competitive equilibrium allocation found has no courses that are over-enrolled by more than k1k-1 seats.
Similarly, Lin et al. (2022) analyze deterministic allocation mechanisms through approximate competitive equilibria with a good-by-good bound. They consider complex feasibility constraints and prove the existence of an approximate competitive equilibrium that satisfies stability and fairness properties similar to the ones analyzed in this paper. We establish some additional properties of our mechanism, proving that it is strategy-proof in the large and results in an outcome that is approximately priority-constrained efficient. With simulations on randomly generated data, they compare their approach to draft mechanisms in the context of reassigning sporting event season tickets to families. We address the problem of allocating courses to undergraduate students by estimating preferences from real-world university data and comparing our proposed mechanism to a version of the mechanism used at many top universities in the United States. In Appendix D, we also provide a comparison of our mechanism with the deferred acceptance mechanism with single and multiple tie-breakings Gale and Shapley (1962), the deferred acceptance mechanism with minority reserves Hafalir, Yenmez and Yildirim (2013), some alternative pseudo-market mechanisms Budish (2011), and a variant of the probabilistic serial mechanism Bogomolnaia and Moulin (2001). Many of these mechanisms have been studied in many-to-one matching settings, but their properties in many-to-many matching with indifferences are not well-understood.
The remainder of the paper is organized as follows. We model undergraduate course allocation in Section 1 and investigate the properties of the PMP mechanism in Section 2. In Section 3, we present simulation results using university data. We offer concluding remarks in Section 4. Proofs are in Appendices A and B. Details on the student utility model calibration and additional simulation results are in Appendices C, D, and E.
1 Environment
Course allocation is a many-to-many matching problem described by the tuple $({\cal S},{\cal C},Q,V,\mathcal{R})$.
  • ${\cal S}=\left\{1,...,N\right\}$ is a set of students; in reference to students, we use she/her pronouns.
  • ${\cal C}=\left\{1,...,M\right\}$ is a set of courses.
  • $Q=(q_{1},...,q_{M})$ is a vector of course capacities; each course $c\in{\cal C}$ has $q_{c}$ seats.
  • $V=(\succsim_{1},....,\succsim_{N})$ is a vector of student preferences over course schedules: subsets of $\cal C$}. Students typically only consider a subset of all possible course schedules due to factors such as course meeting times and prerequisites. Additionally, each student can take at most $k$ total course seats. We assume that these restrictions are incorporated into student preferences and that student preferences are strict. We also assume, for simplicity of exposition, that $1\leq k\leq M/2$. These restrictions still permit students to view courses as substitutes or complements.
  • $\mathcal{R}=\{r_{s,c}\}_{s\in{\cal S},c\in{\cal C}}$ is a course priority structure. $r_{s,c}\in\{1,...,R\}$ specifies the level of priority of student $s$ in course $c$, with $R\geq 1$ and a larger number meaning a higher level of priority. Multiple students can share the same level of priority in a course.
Our setting connects two well-studied problems in market design in a natural way. If $r_{s,c}=1$ for all students $s\in{\cal S}$ and courses $c\in{\cal C}$, this is the combinatorial assignment problem of Budish (2011). If $\{r_{s,c}\}_{s\in{\cal S}}$ forms a strict ordering over all students for each course $c$, this is a two-sided matching problem in which students have strict preferences over subsets of courses and courses strictly rank individual students.
Universities often categorize students into four years of study based on semesters or credits completed. In addition, universities place students into departments, each of which offers a set of courses. These practices illustrate two examples of course priorities: year-specific and department-specific priorities. In a year-specific priority structure, $R=4$, with $r_{s,c}$ equal to the year of study of student $s$ in every course $c$. In a department-specific priority structure, $R=2$, with $r_{s,c}=2$ if student $s$ and course $c$ are in the same department and $r_{s,c}=1$ otherwise. Year-specific priorities give an advantage to students closer to graduation in every course, whereas department-specific priorities give an advantage to students that require a given course to graduate over those that do not. We discuss hybrid priority structures that combine year-specific priorities and department-specific priorities in Section 3.
We consider deterministic allocations of courses to students. An allocation $x=(x_{s})_{s\in{\cal S}}$ assigns the course schedule $x_{s}$ to each student $s\in{\cal S}$. For ease of notation, we view $x_{s}$ as both a subset of courses in $\cal C$} and a vector from the set $\{0,1\}^{M}$. Allocation $x$ is feasible if $\sum_{s\in{\cal S}}x_{s,c}\leq q_{c}$ for each $c\in{\cal C}$, assigning no course a greater number of students than its capacity. Allocation $x$ is individually rational if, for every $s\in{\cal S}$, $x_{s}=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\}$, ensuring that $s$ does not want to drop any of her assigned courses. A pair $(s,C)$ of student $s\in{\cal S}$ and course schedule $C\subset{\cal C}$ is a block of $x$ if a student prefers $C$ to her current course schedule $x_{s}$ and each course in $C$ not in $x_{s}$ has an available seat or a lower priority student assigned to a seat. Formally, $(s,C)$ is a block if $C=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\cup C\}$, $C\neq x_{s}$, and for each $c\in C\setminus x_{s}$, $\sum_{s\in{\cal S}}x_{s,c}<q_{c}$ or there is a student $s^{\prime}\in{\cal S}$ with $c\in x_{s^{\prime}}$ and $r_{s,c}>r_{s^{\prime},c}$.
We evaluate allocations based on stability, efficiency, and fairness. An allocation $x$ is stable if it is feasible, individually rational, and admits no blocks (see Roth and Sotomayor, 1990; Echenique and Oviedo, 2006). [^6] A related concept is justified course envy. Allocation $x$ prevents justified course envy if there are no students $s,s^{\prime}\in{\cal S}$ and course $c\in{\cal C}$ such that $r_{s,c}>r_{s^{\prime},c},\,c\notin x_{s},\,c\in x_{s^{\prime}}$ and $c\in\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x_{s}\cup c\}$. The absence of justified course envy prevents envy towards students of a lower priority, but does not account for individual rationality and the possibility of envy towards several students assigned to several courses.
In the next section, we analyze approximate market equilibria, in which some courses might have seats assigned slightly above capacity or empty demanded seats unfilled. We say such allocations have a market-clearing error. To account for allocations with market-clearing errors, we use an approximate version of stability
Definition 1.
An allocation $x$ is approximately stable if it is stable with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$.
The relationship between stability and approximate stability is similar to the one between the notions of Pareto efficiency and approximate Pareto efficiency in Budish (2011). An allocation $y$ Pareto dominates an allocation $x$ if there is at least one student who strictly prefers her course schedule in $y$ and all other students weakly prefer their course schedules in $y$. Allocation $x$ is Pareto efficient if no allocation $y$ Pareto dominates $x$, and approximately Pareto efficient if no $y$ that Pareto dominates $x$ assigns weakly fewer seats in each course.
Approximate Pareto efficiency is a meaningful notion of efficiency in environments where a market-clearing error is present.
In our setting, a meaningful notion of efficiency should also account for course priorities. Following Schlegel and Mamageishvili (2020), we extend course priorities over individual students to priorities over subsets of students using first-order stochastic dominance. We say that allocation $y$ dominates allocation $x$ for course $c$, and write $y_{c}\succeq_{c}x_{c}$, if for all levels of priority $r\in{\{1,...,R\}}$, $\sum_{s\in{\cal S}:r_{s,c}\geq r}y_{s,c}\geq\sum_{s\in{\cal S}:r_{s,c}\geq r}x_{s,c}$. Using this definition, we consider the following notion of priority-constrained efficiency.
Definition 2.
An allocation $x$ is approximately priority-constrained efficient if, in an environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$, for each allocation $y$ that Pareto dominates $x$, there is a course $c^{\prime}\in{\cal C}$ for which $y_{c^{\prime}}\nsucceq_{c^{\prime}}x_{c^{\prime}}$.
As with approximate stability, we account for a possible market-clearing error in the allocation $x$ by adjusting the course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$.
Except for this adjustment, our definition coincides with the priority-constrained efficiency introduced by Schlegel and Mamageishvili (2020).
The seminal measure of fairness is envy-freeness, as introduced by Foley (1967): an allocation $x$ prevents schedule envy if there are no students $s,s^{\prime}\in{\cal S}$ such that $x_{s^{\prime}}\;\succ_{s}\;x_{s}$. However, without using lotteries, we cannot hope to allocate courses in a completely envy-free way among students (e.g., if two students with the same priorities desire a course with a single seat). Extending the notion originally proposed by Budish (2011), we consider the more permissive concept of schedule envy bounded by a single course toward students of weakly lower priority in every course (see also Lin et al., 2022).
Definition 3.
An allocation $x$ has schedule envy bounded by a single course toward students of weakly lower priority if, for any $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$, either $x_{s}\;\succsim_{s}\;x_{s^{\prime}}$ or there exists some course $c^{*}$ such that $x_{s}\;\succsim_{s}\;(x_{s^{\prime}}\setminus\{c^{*}\})$.
This criterion of fairness minimizes each student’s envy towards students with weakly lower priority in every course.
Allocations are found through mechanisms, which systematically elicit student preferences over course schedules. A mechanism is strategy-proof if there is no student, who, by reporting manipulated preferences, receives an allocation she strictly prefers to the course schedule she would get from reporting her true preferences. Evidence from business schools demonstrates that requiring strategic play on behalf of students can complicate efficiency (see Budish and Cantillon, 2012; Budish and Kessler, 2022; Sönmez and Ünver, 2010). We assume that strategic play is only a concern for students, as course priorities are set based on commonly observable factors such as student year of study or department.
Mechanisms that utilize market equilibria fail to be strategy-proof when students can influence course prices. A mechanism is strategy-proof in the large if truthful reporting is approximately optimal in an environment with a large number of students. For market equilibria, this amounts to being strategy-proof in a limit market in which each student regards the market prices as exogenous to her reported preferences (see Azevedo and Budish, 2019; Budish, 2011). [^7] He et al. (2018) and Liu and Pycia (2016) consider two related notions of asymptotic incentive compatibility and asymptotic strategy-proofness for cardinal and ordinal preferences, respectively. Nguyen, Teytelboym and Vardi (2025) recently introduced a stronger notion of uniform strategy-proofness. To avoid unnecessary heavy notation early in the paper, we introduce this concept formally in the proof of Theorem 5.
2 Pseudo-Markets with Priorities
In this section, we present our novel mechanism, which allocates courses to students by extending the concept of approximate competitive equilibrium from equal incomes from Budish (2011) to settings with course priorities. For this purpose, we assign each student $s\in{\cal S}$ a budget of fake money $b^{*}_{s}$, with $1\leq\min_{s}b^{*}_{s}\leq\max_{s}b_{s}^{*}\leq 1+\beta$ for some small $\beta>0$. The parameter $\beta$ is the maximum allowable budget inequality across students. We also allow for slack in the market-clearing condition, which is bounded by $\alpha\geq 0.$
Definition 4.
The allocation $x^{*},$ priority-specific prices $p^{*},$ and budgets $b^{*}$ constitute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities if:
  1. 1.
    Each student $s\in{\cal S}$ is assigned her most-preferred affordable course schedule:
    $x_{s}^{*}=\max_{\left(\succsim_{s}\right)}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}p_{c,r_{s,c}}^{*}x_{s,c}^{\prime}\leq b_{s}^{*}\right\}$
  2. 2.
    Each course $c\in{\cal C}$ has a cutoff priority level $r^{*}_{c}$ such that:
    $p^{*}_{c,r}\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}\\ {}[0,\overline{b})&r=r^{*}_{c}\\ [\overline{b},+\infty)&r<r^{*}_{c}\end{cases},$ ($\ast$)
    where $\overline{b}=1+\beta+\epsilon$ for some small $\epsilon>0$.
  3. 3.
    The market-clearing error is less than $\alpha$: $z^{*}||_{2}\leq\alpha,$ where $z^{*}=(z_{1}^{*},....,z_{M}^{*})$ and:
    1. (a)
      $z_{c}^{*}=\sum_{s\in{\cal S}}x_{s,c}^{*}-q_{c}$ if $p_{c,1}^{*}>0,$
    2. (b)
      $z_{c}^{*}=\max(\sum_{s\in{\cal S}}x_{s,c}^{*}-q_{c},0)$ if $p_{c,1}^{*}=0$.
  4. 4.
    Student budgets are almost equal: $1\leq\min_{s\in{\cal S}}b_{s}^{*}\leq\max_{s\in{\cal S}}b_{s}^{*}\leq 1+\beta$.
Unlike Budish (2011), the above definition of an approximate competitive equilibrium allows for course prices to depend on priority levels, making the vector of prices $p^{*}=\left\{p^{*}_{c,r}\right\}_{c\in{\cal C},r\in\{1,...,R\}}\in\mathbb{R}^{M\!R}$. Condition ($\ast$2) ensures that each course has a cutoff level of priority, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face some price, and lower priority students cannot afford a seat. In doing so, condition ($\ast$2) guarantees no students are allocated a seat in a course when a higher priority student who would like a seat does not receive one. A similar condition appeared in He et al. (2018) in the context of random allocation with priorities.
The market-clearing error for a course depends only on whether its price equals to zero at $r=1$, the lowest priority level. Under-enrollment is only counted as an error if the price at level $r=1$ is positive. This requirement is a non-trivial extension of Budish (2011)’s definition to settings with course priorities. Alternatively, if a course’s market-clearing error depends on whether the price at the cutoff priority level is zero, one can change the cutoff level and artificially lower the error. For example, consider an under-enrolled course with a price of zero at and above the cutoff priority level and a price of $\overline{b}$ below the cutoff. By the alternative definition, the market-clearing error is zero. However, the unfilled course seats should be counted towards the market-clearing error, as they may be eliminated if the price below the cutoff level is decreased.
Budish (2011) establishes the existence of an approximate competitive equilibrium from equal incomes for any budget inequality $\beta>0$ with an upper bound on the market-clearing error of $\alpha=\sqrt{kM/2}$. [^8] Budish (2011) shows the existence of an approximate equilibrium from equal incomes with a market-clearing error at most $\sqrt{\min\{2k,M\}M}/2$. As we assume $k\leq M/2$, the upper bound reduces to $\sqrt{kM/2}$. This bound is proportional to the square root of the dimension of the price space. In the presence of $R$ priority levels for each of the $M$ courses, the dimension of the price space is MRM\!R. This could potentially require a large market-clearing error for a Pseudo-Market Equilibrium with Priorities to exist. The main idea of the proof of Theorem 1 is to reduce the effective dimension of the price space from MRM\!R to $M$, where we can establish the existence of a Pseudo-Market Equilibrium with Priorities. In particular, for any vector $t\in\mathbb{R}^{M}$, we can define priority-specific prices $p\in\mathbb{R}^{M\!R}$ by setting:
$p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$
for each course $c\in{\cal C}$ and priority level $r\in\{1,...,R\}$. This parameterization guarantees that, for any $t\in[0,R\overline{b})^{M}$, the corresponding price vector $p$ satisfies condition ($\ast$2). We then prove the existence of an approximate equilibrium in the $M$-dimensional space.
Theorem 1 (Existence).
For any $\beta>0$, there exists a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities.
Theorem 1 establishes the existence of a Pseudo-Market Equilibrium with Priorities with the same upper bound on the market-clearing error as the approximate competitive equilibrium from equal incomes, extending the main result of Budish (2011) to environments with course priorities. The proof of Theorem 1 is presented in Appendix A.
2.1 The Pseudo-Market with Priorities Mechanism
Next, using the Pseudo-Market Equilibrium with Priorities, we introduce the Pseudo-Market with Priorities (PMP) mechanism and investigate approximate notions of stability, efficiency, fairness, and strategy-proofness it satisfies.
Definition 5.
The Pseudo-Market with Priorities mechanism with market-clearing error $\alpha$ and budget inequality $\beta$ is defined through the following steps:
  1. 1.
    Each student $s$ reports preferences $\succsim^{\prime}_{s}$ over course schedules.
  2. 2.
    Each student $s$ is assigned a random budget $b^{*}_{s}$ drawn uniformly from $1,1+\beta]$.
  3. 3.
    Compute an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities $(x^{*},p^{*},b^{*})$ using an anonymous method. Allocate courses to students according to $x^{*}$.
In general, the PMP mechanism depends on the level of allowable market-clearing error $\alpha$ and the level of budget inequality $\beta$. We avoid this dependence in our exposition for results that hold for all non-negative $\alpha\geq 0$ and $\beta\geq 0$.
We first establish that the PMP mechanism results in an approximately stable allocation. Definitions 4 and 5 ensure that the allocation is individually rational. With courses regarded as being at their full capacity, condition ($\ast$2) on priority-specific prices prevents blocks. [^9] An alternative statement holds. Let $(x^{*},p^{*},b^{*})$ be an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities, and, for each $c\in{\cal C}$, let $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{sc}$ if $p^{*}_{c,1}>0$ and $q^{\prime}_{c}=\max(\sum_{s\in{\cal S}}x^{*}_{sc},q_{c})$ if $p^{*}_{c,1}=0$. Then, allocation $x^{*}$ is stable in the environment with adjusted capacities $q^{\prime}_{c},$ $c\in{\cal C}$.
Theorem 2 (Stability).
The Pseudo-Market with Priorities mechanism results in an allocation that is approximately stable.
The proof is postponed to Appendix A. The main reason that the outcome of the PMP mechanism is not stable is the possibility of over- and under-enrolled courses. Stability requires a feasible allocation and only allows a course to have unfilled seats if no student can benefit from taking one. The PMP mechanism can violate these conditions when the market-clearing error is nonzero. However, Theorem 1 suggests instances will be rare.
The presence of course priorities prevents the PMP mechanism from being Pareto efficient. This is similar to how stable allocations in matching markets can fail to be Pareto efficient (see Roth and Sotomayor, 1990). We illustrate this in the following example.
Example 1.
Consider a setting with two students ${\cal S}=\{1,2\}$ and two courses ${\cal C}=\{A,B\}$, with $q_{A}=q_{B}=k=1$. Student preferences are $\{A\}\succ_{1}\{B\}\succ_{1}\varnothing$ and $\{B\}\succ_{2}\{A\}\succ_{2}\varnothing$. Student budgets are $b^{*}_{1}=1$ and $b^{*}_{2}=1+\beta$ for some $0<\beta<1$. Course priorities are the opposite of student preferences, with $r_{1,B}=r_{2,A}=2$ and $r_{1,A}=r_{2,B}=1$. Budgets $b^{*}$, cutoff priority levels $r^{*}_{A}=r^{*}_{B}=2$, priority-specific prices $p^{*}_{A}=p^{*}_{B}=(2,1)$, and allocation:
Student $x^{*}_{s,A}$ $x^{*}_{s,B}$
1 0 1
2 1 0
constitute a $(0,\beta)$-Pseudo-Market Equilibrium with Priorities. The allocation is not Pareto efficient, as both students would be made better off by exchanging their assigned seats.
The next result shows that the outcome of the PMP mechanism satisfies a notion of constrained efficiency. Adjusting course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$, no allocation that Pareto dominates the outcome of the PMP mechanism respects priorities to the same degree. For each course, the distribution of assigned student priority levels cannot first-order stochastically dominate the corresponding distribution for the PMP outcome (see Definition 2). [^10] Similarly to Theorem 2, an alternative statement holds. Let $(x^{*},p^{*},b^{*})$ be an $(\alpha,\beta)$-Pseudo-Market Equilibrium with Priorities, and, for each $c\in{\cal C}$, let $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{sc}$ if $p^{*}_{c,1}>0$ and $q^{\prime}_{c}=\max(\sum_{s\in{\cal S}}x^{*}_{sc},q_{c})$ if $p^{*}_{c,1}=0$. Then, allocation $x^{*}$ is priority-constrained efficient in the environment with adjusted capacities $q^{\prime}_{c},$ $c\in{\cal C}$. In environments without course priorities, this result reduces to Proposition 2 in Budish (2011).
Theorem 3 (Efficiency).
The Pseudo-Market with Priorities mechanism results in an allocation that is approximately priority-constrained efficient.
The proof of the above result resembles the proof of Schlegel and Mamageishvili (2020) for single-unit settings with random allocations and is postponed to Appendix A.  [^11] We assume each agent has strict preferences over course schedules. So, unlike Schlegel and Mamageishvili (2020) and Miralles and Pycia (2021), we do not require that a student chooses the cheapest course schedule when multiple course schedules are optimal.
Though the concepts of approximate stability and priority-constrained efficiency are related, neither implies the other. If we assume in Example 1 that both students have the same priority in each course, then there are two (approximately) stable allocations: $x_{1}=\{A\}$, $x_{2}=\{B\}$ and $x_{1}=\{B\}$, $x_{2}=\{A\}$. However, only the first allocation is approximately priority-constrained efficient. Here, approximate priority-constrained efficiency selects among approximately stable allocations, favoring more Pareto optimal outcomes. In turn, suppose that student $1$ is the only student in the environment and that $x_{1}=\{A,B\}$. This allocation is approximately priority-constrained efficient, but not approximately stable, as it is not individually rational.
Theorem 2 prevents the possibility that a student can benefit from taking a seat in a course that a student with lower priority obtains. However, priorities in undergraduate course allocation are weak, with hundreds of students often at the same level of priority across many courses. The next result addresses fairness between students whose priorities are weakly ordered in the same way across all courses.
Theorem 4 (Fairness).
If $\beta\leq\frac{1}{k-1}$, the Pseudo-Market with Priorities mechanism results in an allocation that has schedule envy bounded by a single course toward students of weakly lower levels of priority.
Theorem 4 extends the result established in Budish (2011) for settings without priorities. With priorities based on seniority and student department, the above result implies that, with a sufficiently small budget inequality, the PMP mechanism produces an allocation in which any two students from the same department and the same year of study have schedule envy bounded by a single course toward one another. [^12] This result can be strengthened. The PMP mechanism results in an allocation $x^{*}$ such that for any two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for each $c\in x^{*}_{s^{\prime}}$, envy is bounded by a single course. In particular, rather than all courses, $s$ needs only to be at a weakly higher priority level in the courses being taken by $s^{\prime}$.
Our final theorem establishes that students have almost no incentive to manipulate the PMP mechanism in large populations. This is commonly the case at undergraduate institutions, where student bodies often consist of thousands of students.
Theorem 5 (Strategy-Proofness).
The Pseudo-Market with Priorities mechanism is strategy-proof in the large.
We define the concept of strategy-proofness in the large in the proof of Theorem 5 in Appendix A. To prove this theorem, we leverage on a result from Azevedo and Budish (2019), which establishes a connection between strategy-proofness in the large and envy-freeness. Students can be partitioned into groups based on their course priorities, where each student in a group is at the same level of priority in every course. The PMP mechanism is a semi-anonymous mechanism, as, within groups, each student faces the same course prices and the same random lottery over budgets. As a student cannot prefer the allocation of any lower-budget student in the same group, the PMP mechanism is envy-free but for tie-breaking. Our result then follows from Azevedo and Budish (2019), which shows that any semi-anonymous mechanism that is envy-free but for tie-breaking is strategy-proof in the large. [^13] One can also consider a continuum replication of a course allocation problem with course priorities. By leveraging on the price structure as in Theorem 1, the steps of Theorem 4 of Budish (2011) can be adapted to obtain the result. The alternative proof is available upon request.
We conclude this section by considering special cases discussed in previous literature. In the case of no priorities, our mechanism reduces to the approximate competitive equilibrium from equal incomes mechanism of Budish (2011). On the other hand, courses may have strict priorities over students. In this case, we have a two-sided matching problem. As students can view sets of courses as complements, no stable matching is guaranteed to exist. Still, Theorem 1 confirms a Pseudo-Market Equilibrium with Priorities exists with a small error. So, our mechanism does not necessarily coincide with any stable matching mechanism.
If students have strict preferences over individual courses as well, the student-proposing deferred acceptance mechanism results in the student-optimal stable matching (see Roth, 1984), which can be supported with a Pseudo-Market Equilibrium with Priorities where the market clears exactly, the cutoff level of priority is at the lowest-priority student to receive a seat, and the price of each course is zero at the cutoff. The set of equilibria with zero market-clearing error exactly coincides with the set of stable matchings. [^14] See Miralles and Pycia (2021) for a related result showing that every efficient assignment can be decentralized through prices in random allocation settings. In large markets, we expect the set of stable matchings to be small (see Kojima and Pathak, 2009) and, hence, PMP outcomes should not differ much from each other.
When each student can only enroll in one course, our setting reduces to the unit demand variation of the problem, known in the literature as school choice. The standard mechanisms in school choice are ordinal (Abdulkadiroğlu and Sönmez, 2003) or elicit only restricted information about cardinal preferences from agents (Abdulkadiroğlu, Che and Yasuda, 2015). With weak course priorities, the deferred acceptance algorithm may not return a student-optimal stable matching (Erdil and Ergin, 2008). At the same time, any stable matching that is priority-constrained efficient, student-optimal or not, can be supported by a Pseudo-Market Equilibrium with Priorities with zero market-clearing error.
Allowing random allocations, He et al. (2018) show how to extend the pseudo-market approach of Hylland and Zeckhauser (1979) to school choice settings. Their pricing structure ensures that their pseudo-market equilibrium allocation is fair. We adopt a similar pricing structure but analyze deterministic allocations rather than random allocations. While deterministic allocations are preferable from a practical market design point of view, they are associated with additional complications in equilibrium existence.
3 Simulations
Next, we analyze the performance of the Pseudo-Market with Priorities mechanism with student preferences estimated using course allocation data from a private institution. We compare the Pseudo-Market with Priorities mechanism with a version of the Random Serial Dictatorship with course reserves, the mechanism used for course allocation at many U.S. universities, and Deferred Acceptance with single and multiple tie-breakings, mechanisms actively used in the allocation of students to public schools.
3.1 Student and Course Data
Table 1: The number of students and courses in each college.
College A B C D E F G
# of students 853 1642 259 1274 745 741 509
# of courses 180 84 12 269 88 84 39
Notes: We use capital letters to denote college names. There are $1565$, $1611$, $1422$, and $1425$ students in the first, second, third, and fourth year of study, respectively.
We utilize data from the Spring 2018 semester at a private institution in the mid-Atlantic region, covering 6 0236\,023 students across four years of study and $756$ courses across seven colleges (see Table 1). Colleges are the largest constituent units of the university and typically contain several departments. For example, the Department of Economics is housed in the College of Liberal Arts at a typical U.S. university. Colleges differ in the size of their student population and the number and type of courses they offer. Some colleges serve the whole university, whereas others mainly offer courses only for their students. [^15] We exclude graduate and exchange students. Fourth- and fifth-year students are treated as one group.
Table 2: The percentage of courses at or above maximum capacity.
Seats taken above capacity $\geq 0$ $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses 11.2 7.3 4.1 3.3 2.5 1.9
Notes: Maximum capacity is a soft constraint on enrollment for many courses.
Course enrollments. The university allows courses to be over-enrolled. Table 2 shows that $11.2\%$ of courses meet or exceed their listed capacity, with $7.3\%$ of courses exceeding it. The listed maximum capacity in over-enrolled courses is intended to ensure proper class dynamics and balanced course sections. This fits our treatment of capacity constraints in the Pseudo-Market with Priorities mechanism, which allows over-enrollment through the market-clearing error. We set each course’s capacity equal to the maximum of the capacity observed in the data and the true enrollment.
Course seat reservations. In practice, the course registration process follows a Random Serial Dictatorship with course reserves, where students select courses in order of seniority (with ties broken randomly) and courses make seat reservations for students in certain departments and years of study. Seat reservations allow courses to express priorities for students, ensuring that students in specific groups have access to a fixed number of a course’s seats. [^16] See Celebi and Flynn (2024) for a framework analyzing both reservations and priorities. Table 3 presents quantiles of the distributions of course capacities, enrollment, and total number of reserved seats from the university data. 33 45533\,455 course seats were available across courses, 23 36923\,369 seats were occupied, and 13 92213\,922 were reserved at the start of course registration.
Table 3: Course capacities, enrollment, and reservations.
Quantile 0.10 0.25 0.50 0.75 0.90
Maximum Capacity 8 15 25 50 98
Actual Enrollment 3 7 15 35 72
# of Reserved Seats 0 0 3 20 53
Notes: This table presents some quantiles of the distributions of course capacities, enrollment, and number of reserved seats from the university data. Courses with zero enrollment are dropped from the data.
Table 4 provides some examples of course seat reservations. For instance, Microeconomics reserves $25$ seats for Department $1$ and Department $2$ students in Year $1$. In the data, it is possible that these reservations are done independently by the two departments (e.g., among Year $1$ students, $15$ seats are reserved for Department $1$ students and $10$ seats reserved for Department $2$). Whenever this is the case, we combine these reservations into one by uniting the department names and summing the number of reserved seats. After this simplification, each course has at most one reservation for each year of study and one for “all years of study”.
Table 4: Examples of course seat reservations.
Course Departments Year of Study Total Reserved Seats
Microeconomics Dept 1, Dept 2 Year 1 25
Real Analysis Dept 2 Year 3, Year 4 50
Machine Learning All Year 4 10
Communication Dept 3 All 5
Notes: Each row of this table corresponds to a seat reservation for a different course. “All” denotes that any student, regardless of department/year of study, is eligible for a reserved seat.
Using the data on seat reservations, we create a priority structure that reflects the university’s intentions. The registration process favors more senior students by giving them earlier time slots to choose courses, resembling year-specific priorities. In addition, courses reserve seats for students in specific departments and years of study, resembling department-specific priorities. [^17] Universities that follow these priorities include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. However, each university operates slightly differently. Within a year of study, many universities (e.g., Duke) randomly stagger students across time slots rather than having all register at once. In addition, some universities (e.g., Vanderbilt) reserve seats for students in specific majors or degree programs rather than departments or years of study. We combine these notions into a hybrid priority structure, with year-specific priorities taking precedence over department-specific priorities. Example 2 provides an illustration for Real Analysis in Table 4. As no seats are reserved for 1st- and 2nd-year students, there are no students at the fourth and second levels of priority.
Example 2.
The course seat reservations for Real Analysis in Table 4 results in the following hybrid priority structure for the course:
  • $r_{s,c}=8$: 4th-year students in Dept 2
  • $r_{s,c}=7$: 4th-year students in all departments except Dept 2
  • $r_{s,c}=6$: 3rd-year students in Dept 2
  • $r_{s,c}=5$: 3rd-year students in all departments except Dept 2
  • $r_{s,c}=4$: $\varnothing$
  • $r_{s,c}=3$: 2nd-year students in all departments
  • $r_{s,c}=2$: $\varnothing$
  • $r_{s,c}=1$: 1st-year students in all departments
Student utilities. The data does not contain information on student preferences over course schedules. To recover this information, we use a parametric form for student utilities and the method of simulated moments to calibrate its parameters (see, e.g., Agarwal, 2015). We assume that each student can take up to $k=5$ courses and that preferences are additive over courses. [^18] Less than 3% of the student body is enrolled in more than five courses in the data. Student $s$’s utility from taking course $c$, $u_{sc}$, takes the form
$u_{sc}={\color[rgb]{0,0,0}Y^{\prime}_{s}\theta}+Y^{\prime}_{s}\Gamma H_{c}+{\color[rgb]{0,0,0}z_{c}}+\varepsilon_{sc}.$ (1)
The parametric form in equation (1) contains three non-random components. The first, $Y^{\prime}_{s}\theta$, is a horizontal component determined by the student’s characteristics. $Y_{s}$ is a binary vector that indicates student $s$’s college $a\in{\cal A}=\{A,...,G\}$ and year of study $y\in{\cal Y}=\{1,...,4\}$. So, $Y_{s}^{\prime}\theta=\theta_{ay}$, where vector $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $|{\cal A}||{\cal Y}|=28$ elements to calibrate. The larger $\theta_{ay}$ is, the more courses student $s$ wants to enroll in. This accounts for variability in the number of courses taken by students from different colleges and years of study (see Table C.1 in Appendix C).
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. $H_{c}$ is a binary vector of observed course characteristics indicating course $c$’s college. We restrict matrix $\Gamma$ to have only $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. This accounts for patterns of course enrollment across colleges (see Table C.2 in Appendix C).
The third component, $\color[rgb]{0,0,0}z_{c}$}, is a vertical component that measures the popularity or quality of course $c$. The vector $z$ has length $756$, with one entry for each course. The last term $\varepsilon_{sc}$ is a random utility component that is assumed to have a normal distribution $\varepsilon_{sc}\sim N(0,\sigma^{2})$ with unobserved parameter $\sigma$. As normalizations, the value of the outside option from not taking a course is $u_{s0}=0$ and the standard deviation of the noise parameter is $\sigma=1$.
We cannot jointly identify the parameters $\theta$, $\Gamma$, and $z$ without additional normalization as they enter in equation (1) in an additive way. Hence, we normalize the average horizontal component across students in the same college to $\overline{\theta}_{a}=0$ and set the interaction terms $\gamma_{aa}$ to 0 for each $a\in{\cal A}$. These normalizations allow for an interpretation of $z_{c}$ as the average student utility from taking course $c$ among students in the same college as course $c$. The value of the parameter $\gamma_{aa^{\prime}},a\neq a^{\prime}$ is the average increase in student utility from taking course $c^{\prime}$ in college $a^{\prime}$ over taking course $c$ in student’s own college $a$ when both courses have the same values of vertical components (i.e., $z_{c}=z_{c^{\prime}}$). Overall, we have $819$ independent parameters $\theta$, $\gamma$, and $z$ to calibrate.
Last, we assume that each student’s choice set is limited to $80$ courses. This assumption is motivated by the observation that any given student typically only considers taking a subset of all offered courses in a given term and is made for computational purposes. We randomly draw these $80$ courses among those taken by at least one student from the same college and year. The probability that a student in college $a$ and year $y$ draws course $c$ is proportional to the share of seats taken in course $c$ by students in college $a$ and year $y$, among all seats taken by such students. [^19] We expand some students’ choice sets to make sure enough reserve-eligible students are interested in each course. This ensures that our calibration process runs smoothly. We elaborate on this in Appendix C. Not all $80$ utility entries are necessarily positive.
We use the method of simulated moments to identify the model’s parameters. A description of the calibration procedure, interpretations of the calibrated parameter values, and model fit results are left to Appendix C. Here, we discuss two obstacles in the calibration process. First, in practice, course reservations do not remain constant during the registration process. After the first week of registration, demand for courses becomes more or less clear. To ensure complete course enrollment, departments start relaxing course reservations by admitting students from waiting lists. While there are best practices for handling waiting lists, this phase of enrollment is done solely at each department’s discretion. As a result, a course’s final enrollment can violate its seat reservations set at the beginning of the process. We use Hall’s Marriage Theorem to identify course reserve violations and minimally adjust the number of reserved seats. The seat reservations are lowered in 170 of the 756 courses to ensure the number of reserved seats is consistent with the allocation observed in the data.
The second obstacle concerned calibrating the vertical components $z_{c}$ for courses at maximum capacity. For such courses, no value of $z_{c}$ can lead to a higher enrollment than observed in the data. We identified the vertical components for courses at maximum capacity using data on how quickly seats in those courses were occupied during the registration process.
3.2 Course Allocation Mechanisms
We analyze the four mechanisms described below. We assume students report their utilities truthfully, as each mechanism is strategy-proof or strategy-proof in the large.
The Pseudo-Market with Priorities (PMP) mechanism. Students are randomly assigned evenly spaced budgets between $1$ and $1+\beta$ (with $\beta=\frac{1}{k-1}=0.25$) and placed at one of $R=8$ levels of priority in each course. As in the proof of Theorem 1, we parameterize prices with $t\in[0,R\overline{b}]^{M}$, where $\overline{b}=1.251>1+\beta$. For each $t\in[0,R\overline{b}]^{M}$, course $c\in{\cal C}$, and priority level $r\in{\{1,...,R\}}$, we define priority-specific prices as $p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$. Then, we look for an equilibrium in the lower-dimensional $t$-space.
The search process has two phases. Phase I starts with an educated guess for course prices based on students’ utility-maximizing course schedules. The algorithm searches in the lower-dimensional space of $t$ parameters for a market-clearing error smaller than the theoretical bound $\alpha=\sqrt{kM/2}\approx 43.5$, iteratively computing each student’s utility-maximizing schedule and adjusting prices proportional to the number of over and under-enrolled seats. We move to Phase II if the market-clearing error fails to improve within a certain number of iterations or if the improvement is smaller than $1\%$ after the theoretical bound has been reached. Phase II takes the Phase I prices and works to reduce excess demand, ensuring that no course is over-enrolled by more than k1=4k-1=4 seats and the distribution of over-enrolled courses is smaller than the one observed in the data (see Table 2). [^20] Doing so guarantees the allocation has a market-clearing error below the bound in Theorem 1 and meets the good-by-good bound of Lin et al. (2022). A proof on the existence of an equilibirum with the good-by-good bound in our setting is available by request. We return back to Phase I if an allocation with a market-clearing error greater than $\alpha$ is obtained in Phase II.
The Deferred Acceptance mechanism with single and multiple tie-breakings. These mechanisms use the student-proposing deferred acceptance algorithm to allocate course seats and differ in how they break ties between students at the same level of priority in a course. The first variant uses a single tie-breaking order across all courses. To break ties, we use the budgets assigned in the PMP mechanism, placing students with larger budgets earlier. Without any course seat reservations, this mechanism is equivalent to a random serial dictatorship (with students ordered by seniority and ties broken randomly) and can lead to significant envy between students. The second variant uses multiple, course-specific tie-breakings. We refer to these mechanisms as DA-STB and DA-MTB, respectively. Both have received a thorough analysis in the many-to-one matching literature, but their many-to-many counterparts are underexplored (see, e.g., Erdil and Ergin, 2008, 2017).
The Random Serial Dictatorship (RSD) with optimal course reserves. This mechanism orders students by seniority, breaking ties using the same order as in PMP and DA-STB, and assigns each student the (up to) $k$ courses available that offer her the highest utility. Course seat reservations are treated as described in Section 3.1. If a student is eligible for one of a course’s seat reservations, she can enroll in a reserved or regular seat. Otherwise, the student can only be enrolled in a regular seat. Reserved seats must be occupied before a reserve-eligible student occupies a regular seat.
Using the number of reserved seats observed in the data leads to significant under-performance of RSD with course reserves (see Appendix D.2). In practice, the university’s registration system has two phases. The first is a centralized admission process, which uses a version of the random serial dictatorship with course reserves. The second is a decentralized stage, where departments relax reserves to fill unoccupied seats with waitlisted students. This second phase is not well-structured. Importantly, the university administration sets course reserves more generously in the first phase than it would if the second phase did not exist. This guarantees that prioritized students have access to course seats but leaves many unfilled reserved seats to be assigned in the second phase. Hence, comparing the performance of the other three mechanisms to the performance of RSD with these reserved seats would not be a fair exercise. Instead, we estimate the optimal reserved seats by generating $100$ environments with random student choice sets and utilities, setting a random tie-breaking order, and running the DA-STB mechanism. We take the average number of reserve-eligible students assigned to each course across all environments (rounded to the nearest integer) as the optimal number of reserved seats. We call the result of the RSD mechanism operated with these estimated reserved seats the RSD with optimal course reserves. [^21] The term “optimal” is not used here in the sense of maximizing some objective, but rather to denote a reasonable and practical specification of reserve prices within the current setting.
In the next section, we present simulation results comparing the PMP, DA-STB, DA-MTB, and RSD mechanisms across five measures. These include two measures of student satisfaction: (i) each student’s preferred mechanism and (ii) mean student utility; and three measures of allocation fairness: (iii) the standard deviation of students’ utilities, (iv) the amount of envy (toward students with weakly lower priority), and (v) priority violations (students desiring a course being taken by a student of strictly lower priority). From theory alone, we know that no student can experience envy greater than a single course in the PMP mechanism (Theorem 4), and that the PMP, DA-STB, and DA-MTB mechanisms all prevent priority violations. Otherwise, it is not clear how the four mechanisms will perform relative to one another. Our simulations provide comparisons for the remaining measures and allow us to see if the envy and priority violations have important magnitudes.
3.3 Simulation results
We evaluate the performance of the Pseudo-Market with Priorities, Random Serial Dictatorship with optimal course reserves, and Deferred Acceptance with single and multiple tie-breakings mechanisms across 100 simulation runs. [^22] Simulations were run on Mathematica with a server with 12 Cores of CPU (2.6 GHz Intel Xeon Gold 6126 CPU), 128 GBs of RAM, 70 GBs of SSD hard drive, and a Windows Server 2019 Standard operating system. It takes about forty-five minutes to calculate the outcomes of all four mechanisms for each run. Each run corresponds to a random draw of student choice sets, utilities, a single tie-breaking order kept the same across PMP, DA-STB, and RSD, and multiple course-specific tie-breakings for DA-MTB.
Figure 1: Each student’s preferred mechanism by year of study.
Notes: This figure reports the percent of students who strictly prefer each given mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark. Results are averages across $100$ runs with different random component draws.
Student satisfaction. We take the RSD with optimal course reserves as a benchmark. Figure 1 displays results on the percent of students who strictly prefer each given mechanism to RSD and the percent of students who strictly prefer RSD. Bars represent shares of the total student population and are broken down by year of study. The large share of indifferences across mechanisms can be explained by students taking under-enrolled courses, as only $11.2\%$ of courses are at or above capacity in the data. Around $17.9\%$ of students prefer PMP to RSD, as opposed to $12.8\%$ of students preferring RSD to PMP. As a result, over $5\%$ more of the population - or $300$ more students - strictly prefer the PMP mechanism. A greater number of first-year students strictly prefer the benchmark to each alternative mechanism. The weaker result for Year 1 students can be attributed to the gains for more senior students, who, on average, are assigned $135$ more course seats under PMP than RSD.
In contrast, $88.7\%$ of students are indifferent between the DA-STB and RSD mechanisms. The vast majority of students being indifferent between the two mechanisms can be explained by our using DA-STB to select the optimal course reserves for the RSD mechanism. Moreover, the outcomes of RSD and DA-STB coincide in the absence of course reserves, and if both mechanisms employ the same tie-breaking rule. DA-MTB results in only $59.7\%$ of students with indifferences, with $20.4\%$ of students strictly preferring DA-MTB and $19.9\%$ strictly preferring RSD. These results are a direct consequence of DA-MTB breaking ties on a per-course basis, producing both winners and losers among students. Also, direct calculations show that more students strictly prefer the PMP mechanism to each DA mechanism in each year of study. [^23] The results of the direct comparison of DA-STB and DA-MTB with PMP are available upon request.
Figure 2: Improvements in utility among students with changing schedules.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each given mechanism over the Random Serial Dictatorship with optimal course reserves benchmark. On average, the PMP mechanism produces a different schedule for 416, 658, 478, and 298 students in Years 1, 2, 3, and 4, respectively; the DA-STB mechanism does so for 164, 263, 174, and 78 students; and the DA-MTB mechanism does so for 547, 836, 603, and 444 students. Results are averages across $100$ runs with different random component draws.
Figure 2 presents changes in cardinal utility, displaying each mechanism’s percent improvement in mean utility over the RSD mechanism among students with changing schedules. The PMP mechanism improves mean utility for all years of study, with larger gains for students in later years of study. In both DA mechanisms, the percent improvement in mean utility is negative for Year 1 students and close to zero for Year 2 students. Among students with changing schedules, the DA-STB mechanism provides the largest improvement in mean utility for Years 3 and 4. However, the DA-STB mechanism changes the schedules of only 174 and 78 students for Years 3 and 4, respectively. In contrast, the PMP mechanism changes the schedules for 479 and 298 students for Years 3 and 4. DA-MTB delivers almost no improvement in mean utilities for each cohort. For each year of study, across all students, the PMP mechanism has the highest mean utility of the four mechanisms.
Allocation fairness. We compare allocation fairness across mechanisms using three metrics. First, Figure 3 shows that, within each year of study, the PMP and DA-MTB mechanisms decrease the standard deviation of students’ utilities relative to the RSD benchmark. The same result holds for all but fourth-year students in the DA-STB mechanism. For each year of study, this finding is the strongest for the PMP mechanism. For each mechanism, the largest reduction occurs for second-year students.
Figure 3: The change in the standard deviation of students’ utilities.
Notes: This figure reports the percent change in the standard deviation of students’ utilities for each given mechanism compared to the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
Second, Table 5 reports results on schedule envy toward students of weakly lower priority in each course. The first column displays the percent of students without envy. The values in each subsequent column are obtained by removing one course from each envied schedules and checking whether the student still experiences envy. In line with our theoretical results for the PMP mechanism (see Theorem 4), no student has envy greater than a single course toward a student of weakly lower priority. As Table 5 shows, this property is not satisfied by the other three mechanisms. About $9.1\%$ of students experience envy in the PMP mechanism, substantially less than in the RSD, DA-STB, and DA-MTB mechanisms ($17.2\%$, $14.1\%$, and $13.3\%$, respectively). Relative to the RSD benchmark, $8.1\%$ less of the total population (or almost 500 students) experience any envy in the PMP mechanism.
Table 5: The percentage of students who experience schedule envy.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
Pseudo-Market with Priorities 90.9 (0.6) 9.1 (0.6) 0 (0.0) 0 (0.0) 0 (0.0) 0 (0.0)
Random Serial Dictatorship with optimal course reserves 82.8 (1.2) 13.9 (0.9) 2.6 (0.3) 0.6 (0.2) 0.06 (0.05) 0.002 (0.0)
Deferred Acceptance with single tie-breaking 85.9 (0.6) 11.6 (0.5) 2.1 (0.2) 0.4 (0.2) 0.03 (0.03) 0.001 (0.0)
Deferred Acceptance with multiple tie-breakings 86.7 (0.6) 12.6 (0.6) 0.7 (0.2) 0.02 (0.02) 0 (0.0) 0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the four mechanisms. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across 100 runs with different random component draws.
Third, we examine fairness between levels of priority. The PMP mechanism is approximately stable, guaranteeing that no student can improve her schedule by taking a seat assigned to a student with strictly lower priority in the course (see Theorem 2). As students’ preferences are additive over courses, the DA-STB and DA-MTB mechanisms also ensure that no priorities are violated. However, in the RSD mechanism, $16.2\%$ of students would benefit from taking a seat assigned to a student with strictly lower priority. These instances occur when a course reserves too few of its seats, meaning a reserve-eligible student selecting courses late in the RSD mechanism loses out on a seat when a reserve-ineligible student selecting earlier obtains one.
Our simulations present important comparisons that would be difficult to derive from theory. While there is some variation across years of study, the PMP mechanism generally outperforms the other three mechanisms in each measure of student satisfaction and allocation fairness. [^24] In principle, it may be delivering better outcomes due to the market-clearing error, which allows some courses to be over-enrolled. The average market-clearing error across outcomes is $\approx 21.4$, less than half of the worst-case bound of $\approx 43.5$. In line with the bound in Nguyen and Vohra (2022), no course is over-enrolled by more than k1=4k-1=4 seats. About $3\%$ of courses are also over-enrolled, as opposed to $7.3\%$ in the data. The comparison of DA-STB and DA-MTB in many-to-many settings is not readily available in the literature. Hopefully, these results will direct further theoretical investigations of these mechanisms. Appendix D contains additional simulation results, including a description of the equilibrium prices in the PMP mechanism (Section D.1), a comparison with six alternative mechanisms (Section D.2), results under an alternative priority structure (Section D.3), and further robustness checks (Section D.4).
Limitations. Our student utility model has limitations that a more detailed empirical analysis could address. Notably, assuming additive utilities may be too restrictive, as it fails to consider some complementarities and substitutabilities across courses. Courses covering similar content are substitutes, while a course requiring another as a corequisite (i.e., to be taken at the same time) are complements. Additive utilities allow us to compare the performance of the PMP mechanism with the DA-STB and DA-MTB mechanisms. Unlike these mechanisms, the PMP mechanism can be extended to more general preferences.
One step in expanding this analysis beyond additive utilities is to consider preferences over course sections rather than courses. Universities offer the same course at different times or with different instructors, restricting students from enrolling in two sections of the same course. In the data, many courses have several sections, with some much more popular than others. Since we aggregate enrollment and capacities across sections, if one of a course’s sections is not at maximum capacity, the course cannot be either. $30\%$ of sections are at maximum capacity, as opposed to only $11.2\%$ of courses. Consequently, we expect the gains from the PMP mechanism presented earlier in this section to be understated.
Alternatively, we could consider additional factors that influence preferences. Our utility model assumes preferences only depend on a student’s college and year of study. Allowing utilities to depend on a student’s department would capture more heterogeneity in preferences. Information about past enrollment would disentangle preferences over courses from constraints based on the course’s prerequisites, pinning down student choice sets. Past enrollment, along with other characteristics (e.g., demographics), are not available in our data.
4 Discussion
This paper explores undergraduate course allocation as a many-to-many matching problem. We design a deterministic mechanism, the Pseudo-Market with Priorities (PMP) mechanism, which extends the approximate competitive equilibrium from equal incomes mechanism of Budish (2011) to accommodate course priorities and has the same worst-case bound on the market-clearing error. This mechanism satisfies approximate notions of stability, efficiency, fairness, and strategy-proofness. Our empirical findings exhibit that it outperforms an idealized version of the mechanism used in practice in student satisfaction and fairness.
The gains from introducing the PMP mechanism will generally depend on university-specific factors. When a large set of equally-prioritized students compete for a small set of seats across courses, the commonly-used RSD mechanism produces allocations with significant envy. Instead, the PMP mechanism bounds envy among these students. Hence, the fairness improvements from introducing our mechanism will be large in environments with many over-demanded courses. Similarly, if a university uses a coarser priority structure, there will be more equally-prioritized students, increasing the gains from introducing the PMP mechanism. In contrast, for universities that employ finer priority structures (e.g., prioritizing students based on units completed or grade point average), the PMP and RSD mechanisms will result in similar outcomes.
A university’s practical implementation of the PMP mechanism will also play a role in its benefits. Our model assumes students can report preferences over all possible course schedules. In practice, reporting preferences requires a language simple enough for students to use and expressive enough to elicit students’ actual preferences. The implementation of the A-CEEI mechanism in Budish et al. (2017) allows students to report cardinal values for individual courses and adjustments for pairs of courses. Changes to this language are necessary to make it accessible to undergraduate students. In particular, it should accommodate overlapping lecture times, prerequisite constraints, required courses, and credit limits, blocking students from receiving unacceptable schedules. The deployment of machine learning-based techniques may also improve the precision of elicited preferences (Soumalias et al., 2024).
Finding an equilibrium given reported preferences is computationally intense. Budish et al. (2017) estimate that 20 times the time used at Wharton would be necessary to implement the A-CEEI mechanism for the full student body at Ohio State University. However, implementing the PMP mechanism may be less challenging than this estimate suggests. Since a course’s price is zero or unaffordable at all but one of its priority levels, students face few courses with a positive but affordable price. In our simulations, with additive student utilities, the average timing of the PMP mechanism improves on the A-CEEI mechanism by a factor of two (see Appendix D.2). Still, incorporating scheduling constraints and preference complementarities may be challenging and should be carefully addressed in practice.
Finally, we note that our analysis applies to many other matching problems. Two under-explored examples are assigning teaching assistants to courses and referees to soccer matches. Extending our results to these and other applications is left for future research.
Appendix A Appendix: Proofs
Proof of Theorem 1. Consider a course allocation problem $({\cal S},{\cal C},Q,{\cal V},\mathcal{R})$. We show that, for any $\beta>0$, there exists a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities.
Let $b=(b_{1},...,b_{N})$ be a budget vector that satisfies $1\leq\min_{s}(b_{s})\leq\max_{s}(b_{s})\leq 1+\beta$ for some $\beta>0$, and let $\overline{b}=1+\beta+\epsilon$ for a fixed $\epsilon>0$. Consider the $M$-dimensional set ${\cal T}=[0,R\overline{b}]^{M}$, which conveniently parameterizes priority-specific prices and allows us to look for a competitive equilibrium in a lower dimensional space. In particular, for each parameter $t\in{\cal T}$, course $c\in{\cal C}$, and level of priority $r\in{\{1,...,R\}}$, we define priority-specific prices as
$p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0).$ (A1)
With this definition, for each $t\in{\cal T}$ and $c\in{\cal C}$, there is a unique cutoff level of priority $r^{*}_{c}(t)\in{\{1,...,R+1\}}$ [^25] Note that $r^{*}_{c}(t)=R+1$ is only achieved if $t=R\overline{b}$. The cutoff level $r^{*}_{c}$ can always be reduced to $R$. such that for any $r\in\{1,...,R\}$, $p_{c,r}(t)$ satisfies
$p_{c,r}(t)\in\begin{cases}\left\{0\right\}&r>r^{*}_{c}(t)\\ {}[0,\overline{b})&r=r^{*}_{c}(t)\\ [\overline{b},R\overline{b}]&r<r^{*}_{c}(t)\end{cases}.$ (A2)
[…]
$d_{s}(\tilde{t})=\max_{\succsim_{s}}\left\{x^{\prime}_{s}\subseteq{\cal C}:\sum_{c\in{\cal C}}x^{\prime}_{s,c}\max(\tilde{t}_{c}-(r_{s,c}-1)\overline{b},0)\leq b_{s}+\tau_{s,x^{\prime}_{s}}\right\},$
where the $\tau_{s,x_{s}}$ are student- and schedule-specific taxes chosen to ensure that the demand is single-valued. Taxes are chosen to favor more-preferred bundles (i.e., if $x^{\prime}_{s}\succ_{s}x_{s}$, then $\tau_{s,x^{\prime}_{s}}>\tau_{s,x_{s}}$). For each course $c\in{\cal C}$, the excess demand function $z_{c}:\tilde{{\cal T}}\rightarrow\mathbb{Z}$ is defined by
$z_{c}(\tilde{t})=\sum_{s\in{\cal S}}x^{*}_{s,c}-q_{c},$
where $x^{*}_{s}=d_{s}(t^{*})$ for all $s\in{\cal S}$. Excess demand is bounded, as $N\leq z_{c}\leq N-1$ for all $c\in{\cal C}$. We also define a budget surface for each student $s\in{\cal S}$ and schedule $x_{s}\subseteq{\cal C}$ as
$H(s,x_{s})=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s,c}\max(\tilde{t}_{c}-(r_{s,c}-1)\overline{b},0)=b_{s}+\tau_{s,x_{s}}\right\}.$
Unlike the case without priorities, the budget surface $H(s,x_{s})$ may not be a hyperplane (see Budish, 2011). Lemma B1 in Appendix B shows that it is still possible to choose $b_{s}$ and $\tau_{s,x_{s}}$ such that at most $M$ budget constraints intersect for any $\tilde{t}\in\tilde{{\cal T}}$.
[…]
$\text{trunc}(\tilde{t})_{c}=\min\{R\overline{b},\max\{0,\tilde{t}_{c}\}\}.$
Also, for $\gamma\in(0,1/N)$, we define a tâttonnement price adjustment function $f:\tilde{{\cal T}}\rightarrow\tilde{{\cal T}}$ by
$).f(\tilde{t})=\text{trunc}(\tilde{t})+\gamma z(\text{trunc}(\tilde{t}$)).
[…]
  • The definition of the demand function $d_{s}(t^{*})$ implies that any course schedule that student $s$ prefers to $x^{*}_{s}=d_{s}(t^{*})$ must cost strictly more than $b^{*}_{s}=b_{s}+\tau_{s,x^{*}_{s}}$.
  • Prices $\{p_{c,r}(t^{*})\}_{c\in{\cal C},r\in{\{1,...,R\}}}$ and cutoff levels of priority defined by (A1) and (A2) ensure that condition ($\ast$2) of Definition 4 is satisfied.
[…]
  • $p_{c,1}(t^{*})=0$ implies $z_{c}(t^{*})\leq 0$. To see this, consider two cases. If $\tilde{t}_{c}^{*}\in(0,R\overline{b})$, the fixed point equation implies $z_{c}(t^{*})=0$. If $\tilde{t}_{c}^{*}\in[-1,0]$, we have $t_{c}^{*}\equiv\text{trunc}(\tilde{t}^{*}_{c})=0$ and $p_{c,1}(t^{*})=0$, so the fixed point equation ensures that $z_{c}(t^{*})\leq 0$.
[…]
The remainder of the proof uses $F$ to show that there exists budgets for which excess demand at prices $t^{*}$ are at most a distance of $\sqrt{kM/2}$ from a perfect market-clearing excess demand vector. The reduction of the MRM\!R-dimensional space of prices $\{p_{c,r}\}_{c\in{\cal C},{r\in{\{1,...,R\}}}}$ to $M$-dimensional space $\tilde{{\cal T}}=[-1,R\overline{b}+1]^{M}$ allows us to use the steps of Budish (2011) to establish the existence of a $(\sqrt{kM/2},\beta)$-Pseudo-Market Equilibrium with Priorities. For completeness, we provide adapted steps in Appendix B.∎
Proof of Theorem 2. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism and suppose each course $c\in{\cal C}$ instead has capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$. We show that the allocation $x^{*}$ is stable. Since each course is at full capacity, $x^{*}$ must be feasible. Definitions 4 and 5 ensure that $x^{*}_{s}$ is individually rational as well. Last, suppose that $(s,C)$ is a block. Then, $C=\max_{\succeq_{s}}\{x^{\prime}_{s}\,:\,x^{\prime}_{s}\subseteq x^{*}_{s}\cup C\}$, $C\neq x^{*}_{s}$, and for each $c\in{\cal C}$ with $c\notin x^{*}_{s}$, there is a student $s^{\prime}\in{\cal S}$ with $c\in x^{*}_{s^{\prime}}$ and $r_{s^{\prime},c}<r_{s,c}$. There must exist at least one course $c\in{\cal C}$ with $c\notin x^{*}_{s}$ such that $p^{*}_{c,r_{s,c}}>0$; otherwise, student $s$ should have been assigned $C$ instead of $x^{*}_{s}$. The structure of equilibrium prices implies that for any student $s^{\prime}$ with a lower priority than $s$ for course $c$ we must have $p^{*}_{c,r_{s^{\prime},c}}\geq\overline{b}$. Therefore, we obtain a contradiction: $x^{*}$ cannot possibly assign a student $s^{\prime}$ a seat in course $c$. Hence, there are no blocks, and the outcome of every Pseudo-Market with Priorities mechanism is approximately stable. ∎
Proof of Theorem 3. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism. We show that the allocation $x^{*}$ is approximately priority-constrained efficient. Suppose each course $c\in{\cal C}$ has capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x^{*}_{s,c}$ and that $y$ is an allocation that Pareto dominates $x^{*}$. Allocation $y$ cannot assign more students to some course $c$ than $x^{*}$ does as it would imply that $\sum_{s\in{\cal S}}y^{*}_{s,c}>q^{\prime}_{c}$. If $y$ assigns fewer students to some course $c$ than $x^{*}$ does, then $\sum_{s\in{\cal S}:r_{s,c}\geq 1}x^{*}_{s,c}>\sum_{s\in{\cal S}:r_{s,c}\geq 1}y_{s,c}$, confirming that $y_{c}\nsucceq_{c}x_{c}^{*}$.
Now, suppose that allocation $y$ has the same number of seats assigned in each course. Then, there is a student $s^{\prime}$ who strictly prefers their assigned course schedule in $y$ to their schedule in $x^{*}$. For all students $s\in{\cal S}$, we must have that
$\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}y_{s,c}\geq b^{*}_{s}\geq\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}x^{*}_{s,c},$
with the first inequality holding strictly for student $s^{\prime}$. Across all students, we have that
$\sum_{s\in{\cal S}}\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}y_{s,c}>\sum_{s\in{\cal S}}\sum_{c\in{\cal C}}p^{*}_{c,r_{s,c}}x^{*}_{s,c}.$
Rearranging,
$\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}=r}y_{s,c}>\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}=r}x^{*}_{s,c}.$
Hence, we obtain
$0<\sum_{c\in{\cal C}}\sum_{r=1}^{R}p^{*}_{c,r}\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})=\sum_{c\in{\cal C}}\sum_{r=2}^{R}(p^{*}_{c,r}-p^{*}_{c,r-1})\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})+p^{*}_{c,1}\sum_{s:\,r_{s,c}\geq 1}(y_{s,c}-x^{*}_{s,c}).$
The final term must equal zero, as allocations $y$ and $x^{*}$ assign the same number of seats in each course. Also, $p^{*}_{c,r}-p^{*}_{c,r-1}\leq 0$ for all levels $r$, so we must have $\sum_{s:\,r_{s,c}\geq r}(y_{s,c}-x^{*}_{s,c})<0$ for some course $c$ and level $r$. This implies $y_{c}\nsucceq_{c}x^{*}_{c}$ and confirms that allocation $x^{*}$ is approximately priority-constrained efficient.∎
Proof of Theorem 4. Let $(x^{*},p^{*},b^{*})$ be an outcome of the Pseudo-Market with Priorities mechanism and suppose that $\beta\leq\frac{1}{k-1}$. We show that the allocation $x^{*}$ has schedule envy bounded by a single course toward students of weakly lower levels of priority. Consider two students $s,s^{\prime}\in\cal S$ such that $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$. Denote the prices faced by $s$ and $s^{\prime}$ as $p^{*}_{s}=\{p^{*}_{c,r_{s,c}}\}_{c\in{\cal C}}$ and $p^{*}_{s^{\prime}}=\{p^{*}_{c,r_{s^{\prime},c}}\}_{c\in{\cal C}}$ and the course schedule assigned to student $s^{\prime}$ as $x^{*}_{s^{\prime}}=\{c_{j_{1}},...,c_{j_{k^{\prime}}}\}$ for $c_{j_{1}},...,c_{j_{k^{\prime}}}\in\{1,...,M\}$ and $k^{\prime}\leq k$. Suppose that $s$ envies $s^{\prime}$ and $s$ cannot afford the course schedule of $s^{\prime}$ even if one course from $x^{*}_{s^{\prime}}$ is dropped. That is,
$p^{*}_{s}\cdot x^{*}_{s^{\prime}}\backslash\{c_{j_{\ell}}\}>b^{*}_{s}$
for $\ell=1,...,k^{\prime}$. Since $r_{s,c}\geq r_{s^{\prime},c}$ for all $c\in{\cal C}$, $p^{*}_{s^{\prime}}\geq p^{*}_{s}$, guaranteeing that
$p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}\backslash\{c_{j_{\ell}}\}>b^{*}_{s}$
for $\ell=1,...,k^{\prime}$. As $b^{*}_{s^{\prime}}\geq p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}$, summing these inequalities over $\ell$, we obtain
$(k^{\prime}-1)b^{*}_{s^{\prime}}\geq(k^{\prime}-1)p^{*}_{s^{\prime}}\cdot x^{*}_{s^{\prime}}>k^{\prime}b^{*}_{s}.$
This implies $\frac{b^{*}_{s^{\prime}}}{b^{*}_{s}}>\frac{k^{\prime}}{k^{\prime}-1}\geq\frac{k}{k-1}\geq 1+\beta$, which contradicts how budgets are allocated.∎
Proof of Theorem 5. We first formally define a direct mechanism, a semi-anonymous direct mechanism, and the property of being strategy-proof in the large. Next, we show that the PMP mechanism is a semi-anonymous direct mechanism that is envy-free but for tie-breaking. The result then follows from Appendix C in Azevedo and Budish (2019).
Consider a sequence of environments labeled by the number of students $|{\cal S}^{N}|\equiv N$. The set of courses $\cal C$} is fixed, with the number of courses equal to $M$, but the capacity of each course $q^{N}_{c}$ can vary. Each student $s\in{\cal S}^{N}$ has a type $v_{s}=(\succeq_{s},r_{s})$ that describes her preferences $\succeq_{s}$ over course schedules and her priorities $\{r_{s,c}\}_{c\in\cal C}$. The set of all possible types is denoted as $\cal V$}. An allocation $x\in X^{N}_{0}\equiv\{0,1\}^{MN}$ specifies a schedule $x_{s}$ for each student $s$. Letting $X=\Delta X_{0}$ denote the set of random allocations, we assume that each student has a von Neumann-Morgenstern utility function $u_{v}:X\rightarrow[0,1]$ consistent with her type $v\in{\cal V}$.
We consider a sequence of direct mechanisms $\{\Phi^{N}\}_{N\in\mathbb{N}}$ that, for each report of student preferences, assigns a distribution over course allocations; that is, $\Phi^{N}:{\cal V}^{N}\rightarrow\Delta(X_{0}^{N})$. The PMP mechanism is an example of a semi-anonymous direct mechanism, where agents are divided into groups, agents within each group are treated the same way, and agents across groups can be treated differently. Formally, we partition ${\cal S}^{N}$ into groups according to their course priorities $\left\{S^{N}_{g}\right\}_{g\in\{1,...,R\}^{M}}$. This partition places students who have the same level of priority in every course in one group. All students in the same group face the same prices in the PMP mechanism and the same lottery over budgets. However, these students might have different preferences $\succeq_{s}$ and, hence, types $v_{s}$. We assume that each student can misrepresent her preferences, but not her priorities.
To define the property of a direct mechanism being strategy-proof in the large, we consider function $\phi^{N}:{\cal V}\times\Delta{\cal V}\rightarrow X$ for each $N$ according to
$\phi^{N}(v_{s},m)=\sum_{v_{-s}\in{\cal V}^{N-1}}\Phi_{s}^{N}(v_{s},v_{-s})Pr(v_{-s}|v_{-s}\sim iid(m)),$
where $\Phi_{s}^{N}(v_{s},v_{-s})$ denotes the course schedule obtained by student $s$ when she reports type $v_{s}$ and all other students report $v_{-s}$. $Pr(v_{-s}|v_{-s}\sim iid(m))$ denotes the probability that profile $v_{-s}$ is realized when other students’ types $v_{-s}$ are independent and identically distributed according to $m\in\Delta{\cal V}$. In other words, $\phi^{N}(v_{s},m)$ describes the random outcome that student $s$ expects to receive when she reports $v_{s}$ and the other students’ types are distributed independently and identically according to mm—and they report their types truthfully. Using this notation, we have the following definition.
Definition A1.
A semi-anonymous direct mechanism $\{\Phi^{N}\}_{N\in\mathbb{N}}$ is strategy-proof in the large if, for any random distribution of reports by other students $m\in\Delta{\cal V}$ with full support and $\varepsilon>0$, there exists $N_{0}$ such that, for all $N\geq N_{0}$ and all $v^{\prime}_{s},v_{s}\in{\cal V}$ where students can misrepresent only their preferences, we have $u_{v_{s}}[\phi^{N}(v_{s},m)]\geq u_{v_{s}}[\phi^{N}(v^{\prime}_{s},m)]-\varepsilon.$
In other words, in a large enough market, reporting preferences truthfully is approximately optimal, for any independent and identical distribution of the other students’ types that has full support. Students can only misrepresent their preferences, not their priorities.
Appendix C in Azevedo and Budish (2019) provides an argument explaining why any semi-anonymous mechanism that is envy-free but for tie-breaking is strategy-proof in the large. We define envy-freeness but for tie-breaking below.
Definition A2.
A semi-anonymous direct mechanism $\{\Phi^{N}\}_{N\in{\cal N}}$ is envy-free but for tie-breaking if for each $N$ there exists $x^{N}:({\cal V}\times[0,1])^{N}\rightarrow\Delta(X_{0}^{N})$, symmetric over its coordinates, such that $\Phi^{N}(v)=\int_{\ell\in[0,1]^{N}}x^{N}(v,\ell)d\ell,$ and, for all $s,s^{\prime},N,v$, and $\ell$, if $\ell_{s}\geq\ell_{s^{\prime}}$, and $v_{s}$ and $v_{s^{\prime}}$ belong to the same group $S_{g}$ (that is, $r_{s}=r_{s^{\prime}}$), then $u_{v_{s}}(x^{N}_{s}(v,\ell))\geq u_{v_{s}}(x^{N}_{s^{\prime}}(v,\ell))$.
The PMP mechanism assigns each student a budget using a uniformly distributed lottery $\ell_{s}\sim U[0,1]$ as $b_{s}=1+\ell_{s}\beta$. Because budgets are random, the PMP assignment for each $v\in{\cal V}^{N}$ can be represented as $\Phi^{N}(v)=\int_{\ell\in[0,1]^{N}}x^{N}(v,\ell)d\ell,$ where $x^{N}(v,\ell)\in\{0,1\}^{MN}$ assigns a course schedule to each student in ${\cal S}^{N}.$ Definition 4 guarantees that student $s$ with budget $b_{s}$ (lottery number $l_{s}$) prefers her course schedule to the course schedule of student $s^{\prime}$ in the same group with budget $b_{s^{\prime}}<b_{s}$ ($\ell_{s^{\prime}}<\ell_{s}$). This confirms that the PMP mechanism is strategy-proof in the large and completes the proof.∎
Appendix B Proof of Theorem 1 (Omitted Details)
Appendix A shows how Theorem 1 of Budish (2011) can be extended to accommodate priority-specific course prices. Below, we confirm it is possible to choose budgets $b_{s}$ and $\tau_{s,x_{s}}$ such that at most $M$ budget constraints intersect for any $\tilde{t}\in\tilde{{\cal T}}$. This fact is used on p. A.
Lemma B1.
One can choose taxes $\left\{\tau_{s,x_{s}}\right\}_{s\in{\cal S},x_{s}\subseteq{{\cal C}}}$ that satisfy the following conditions:
  • (i)
    Taxes are small: $\varepsilon<\tau_{s,x_{s}}<\varepsilon$.
  • (ii)
    Taxes favor more preferred bundles: $\tau_{s,x_{s}}>\tau_{s,x^{\prime}_{s}}$ for $x^{\prime}_{s}\succ_{s}x_{s}$.
  • (iii)
    The inequality bounds are preserved:
    $1\leq\min_{s,x_{s}}(b_{s}+\tau_{s,x_{s}})\leq\max_{s,x_{s}}(b_{s}+\tau_{s,x_{s}})\leq 1+\beta.$
  • (iv)
    No perturbed budgets are equal: $b_{s}+\tau_{s,x_{s}}\neq b_{s^{\prime}}+\tau_{s^{\prime},x_{s^{\prime}}}$.
  • (v)
    There is no auxiliary price vector $\tilde{t}\in\tilde{\cal T}$ at which more than $M$ budget constraints $H(s,x_{s})$ intersect.
Proof.
[…]
$\forall s\in{\cal I},c\in{\cal C}:r_{s,c}>r^{*}_{c}(\tilde{t}),\hskip 8.5359ptx_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})=0.$
[…]
$\forall s\in{\cal I},c\in{\cal C}:r_{s,c}<r^{*}_{c}(\tilde{t}),\hskip 8.5359ptx_{s,c}\cdot p_{c,r_{s,c}}(\tilde{t})=0.$
[…]
$\left\{\begin{array}[]{cc}x_{s,1}\cdot p_{1,r_{s,1}}(\tilde{t})+x_{s,2}\cdot p_{2,r_{s,2}}(\tilde{t})+...+x_{s,M}\cdot p_{M,r_{s,M}}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ x_{s^{\prime},1}\cdot p_{1,r_{s^{\prime},1}}(\tilde{t})+x_{s^{\prime},2}\cdot p_{2,r_{s^{\prime},2}}(\tilde{t})+...+x_{s^{\prime},M}\cdot p_{M,r_{s^{\prime},M}}(\tilde{t})&=b_{s^{\prime}}+\tau_{s,x^{\prime}_{s}}\end{array}\right$.
Here, $s,s^{\prime}\in{\cal I}$, the coefficients $x_{s,c}\in\{0,1\}$ are in line with the bundles $x_{s}$ on the budget constraints, and each price satisfies $p_{c,r_{s,c}}=p_{c,r^{*}_{c}}$ or is such that $x_{s,c}\cdot p_{c,r_{s,c}}=0$. Equivalently, setting $\tilde{x}_{s,c}=1$ if $x_{s,c}=1$ and $r_{s,c}=r^{*}_{c}$ and $\tilde{x}_{s,c}=0$, we have the following set of equations:
$\left\{\begin{array}[]{cc}\tilde{x}_{s,1}\cdot p_{1,r^{*}_{1}}(\tilde{t})+\tilde{x}_{s,2}\cdot p_{2,r^{*}_{2}}(\tilde{t})+...+\tilde{x}_{s,M}\cdot p_{M,r^{*}_{M}}(\tilde{t})&=b_{s}+\tau_{s,x_{s}}\\ ....&....\\ \tilde{x}_{s^{\prime},1}\cdot p_{1,r^{*}_{1}}(\tilde{t})+\tilde{x}_{s^{\prime},2}\cdot p_{2,r^{*}_{2}}(\tilde{t})+...+\tilde{x}_{s^{\prime},M}\cdot p_{M,r^{*}_{M}}(\tilde{t})&=b_{s^{\prime}}+\tau_{s,x^{\prime}_{s}}\end{array}\right$.
[…]
Our proof of Theorem 1 in Appendix A extends steps 1-3 of the proof of Theorem 1 from Budish (2011) to accommodate priority-specific course prices. The rest of this section adapts the remaining steps from this proof to our setting.
Step 4. Suppose that $t^{*}$ is not on any budget constraint. We show that $t^{*}$ is an exact competitive equilibrium price vector. There is a neighborhood around $t^{*}$ where each agent’s demand is unchanging in price. At price $t^{*}$, $f$ is continuous, and as a result, $F(t^{*})=f(t^{*})$.
[…]
  • If $t^{*}\neq\tilde{t}^{*}$, we establish the following lemma.
    Lemma B2.
    For any $\tilde{t}\in\tilde{\cal T}\setminus{\cal T}$, (i) $f(\tilde{t})=f($trunc(t~))($\tilde${t})) and (ii) $F(\tilde{t})\subseteq F($trunc(t~))($\tilde${t})).
    Proof.
[…]
Step 5. Next, suppose that $t^{*}$ is on $1\leq L\leq M$ budget constraints. We denote $\Phi=\{0,1\}^{L}$ and construct a set of $2^{L}$ price vectors $\{t^{\phi}\}_{\phi\in\Phi}$ that satisfy the following conditions:
  1. 1.
    Each $t^{\phi}$ is close enough to $t^{*}$ such that there is a path from $t^{\phi}$ to $t^{*}$ that does not cross any budget constraint.
  2. 2.
[…]
To construct vectors $\{t^{\phi}\}_{\phi\in\Phi}$, note that each of the $L$ intersecting budget constraints defines two sets:
$H^{0}_{\ell}$ $=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)\leq b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
$H^{1}_{\ell}$ $=\left\{\tilde{t}\in\tilde{{\cal T}}:\sum_{c\in{\cal C}}x_{s_{\ell}c}\max(\tilde{t}_{c}-(r_{s_{\ell},c}-1)\overline{b},0)>b_{s_{\ell}}+\tau_{s_{\ell},x_{s_{\ell}}}\right\}$
The first set delineates the set of prices for which agent $s_{\ell}$ can afford schedule $x_{s_{\ell}}$, whereas the second set delineates the set of prices for which agent $s_{\ell}$ can’t afford $x_{s_{\ell}}$. Let $\phi=(\phi_{1},...,\phi_{L})\in\Phi$ be an $L$-dimensional vector of zeros and ones, and the polytope $\pi(\phi):=\cap_{\ell=1}^{L}H^{\phi_{\ell}}_{\ell}$ be the set of points in $\cal T$} that belongs to the intersection of sets indexed by $\phi$. Let $H=\{H(s,x_{s})\}_{s\in{\cal I},x_{s}\subseteq{\cal C}}$ be the finite set of all budget constraints formed by any student-schedule pair $(s,x_{s})$. We then define the distance
$\delta<\inf_{\tilde{t}^{\prime\prime}\in{\cal T},H\in{\cal H}}\{||(t^{*}-\tilde{t}^{\prime\prime})||_{2}:\tilde{t}^{\prime\prime}\in H,t^{*}\notin H\},$
which is such that any budget constraint that $t^{*}$ does not belong to is further than $\delta$ away from $t^{*}$. Now, for each $\phi\in\Phi$, we define $t^{\phi}$ by an arbitrary element of $\pi(\phi)\cap B_{\delta}(t^{*})$, where $B_{\delta}(t^{*})$ is a $\delta$-ball of $t^{*}$. Then, the set of price vectors $\{t^{\phi}\}_{\phi\in\Phi}$ satisfies the two requirements above.
Step 6. We now show that a perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$. For this purpose, we first show that for any $y\in F(t^{*})$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ Take some $y\in F(t^{*})$. Consider a sequence $t^{w}\rightarrow t^{*},t^{*}\neq t^{w}\in\tilde{\cal T}\text{ such that }f(t^{w})\rightarrow y^{\prime}$. Note that the sequence $t^{w}$ consists of a finite number of subsequences $t^{w,\phi}\in\pi(\phi)\cap B_{\delta}(t^{*})$ for some $\phi\in\Phi$. [^27] Some subsequences can have only a finite number of elements. Since all elements of the subsequence $t^{w,\phi}$ are on the same side of set $H^{\phi}_{\ell}$, every agent has the same choice at every point. Hence, if the subsequence has an infinite number of elements, we have $t^{w,\phi}\rightarrow t^{*}$ and
$f(t^{w,\phi})\rightarrow t^{*}+\gamma z(t^{\phi}).$
Therefore, the limit of $f(t^{w})$ for the original sequence $t^{w}$ must be also $t^{*}+\gamma z(t^{\phi^{\prime}})$ for some $\phi^{\prime}\in\Phi$. So, $y^{\prime}=t^{*}+\gamma z(t^{\phi^{\prime}})$. Since, by definition, $y$ is a convex combination of such $y^{\prime}$, we must have $y=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi})$, where $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$
Lemma B2 implies that $\tilde{t}^{*}\in F(\tilde{t}^{*})\subseteq F(t^{*})$. Hence, we must have
$\tilde{t}^{*}=t^{*}+\sum_{\phi\in\Phi}\lambda^{\phi}\gamma z(t^{\phi}).$
for nonnegative $\{\lambda_{\phi}\}_{\phi\in\Phi}$ with $\sum_{\phi\in\Phi}\lambda^{\phi}=1.$ We also denote
$\zeta=\sum_{\phi\in\Phi}\lambda^{\phi}z(t^{\phi})=\frac{\tilde{t}^{*}-t^{*}}{\gamma}.$
We note that $\zeta\leq 0$ and $\zeta_{c}<0$ imply $t_{c}^{*}=0$. Hence, vector $\zeta$ is a perfect market-clearing excess demand vector, and it lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$.
Steps 7-8. The structure of excess demand has the same geometric structure as in Budish (2011). In particular, let $L^{\prime}$ be the number of agents whose budget constraints intersect at price $t^{*}$, and renumber agents so that $s=1,...,L^{\prime}$. We denote the number of budgets of student $s$ that intersect at $t^{*}$ as $w_{s}$. Since at most $M$ budgets constraints can intersect, we must have $L\equiv\sum_{s=1}^{L^{\prime}}w_{s}\leq M$. We also denote the bundles pertaining to $s$’s budget constraints as $x_{s}^{1}\succ...\succ x_{s}^{w_{s}}$.
Similarly to Budish (2011), we consider bundles that $s$ demands at prices near $t^{*}$. In the set $H^{0}(s,x^{1}_{s})$, agent $s$ can purchase her favorite bundle $x^{1}_{s}$. Hence, one does not need to know whether prices belong to sets $H^{0}(s,x_{s}^{2}),...,H^{0}(s,x_{s}^{w_{s}})$. Let us denote the demand for prices at $H^{0}(s,x^{1}_{s})\cap B_{\delta}(t^{*})\cap\tilde{\cal T}$ as $d_{s}^{0}$. Similarly, we consider prices in $H^{1}(s,x^{m}_{s})\cap H^{0}(s,x^{m+1}_{s})$ and denote the corresponding demands as $d_{s}^{m}$ for $m=1,...,w_{s}$. Overall, agent $s=1,...,L^{\prime}$ purchases $w_{s}+1$ distinct bundles at prices near to $t^{*}$.
Let us denote the excess demand of the remaining agents as
[…]
A perfect market-clearing excess demand vector lies in the convex hull of $\{z(t^{\phi})\}_{\phi\in\Phi}$ with the elements
[…]
where $0\leq a_{s}^{f}\leq 1,s=1,...,L^{\prime},f=1,...,w_{s}$ and $\sum_{f=1}^{w_{s}}a_{s}^{f}=1,$ $s=1,...,L^{\prime}$. Budish (2011) shows in Step 8 of Theorem 1 that there exists an element of the above geometric structure, $z(t^{\phi^{\prime}})$, within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector. This purely mathematical argument remains unchanged.
Step 9. While $z(t^{\phi^{\prime}})$ is within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector, it may be that $t^{\phi^{\prime}}\in\tilde{\mathcal{T}}\setminus\mathcal{T}$ rather than $\mathcal{T}$. Since we know that $t^{*}\in\mathcal{T}$, we can alter student budgets $b$ so that excess demand at $t^{*}$ equals $z(t^{\phi^{\prime}})$.
For any student $s\in\mathcal{S}\setminus\{1,...,L^{\prime}\}$, we know that $s$ selects the same course schedule at $t^{\phi^{\prime}}$ and $t^{*}$. As a result, setting $b^{*}_{s}=b_{s}+\tau^{\prime}_{s,x^{*}_{s}}$, the price of any bundle preferred by $s$ to $x^{*}_{s}$ costs more than $b^{*}_{s}$.
Any student $s\in\{1,...,L^{\prime}\}$ may select a different course schedules at $t^{\phi^{\prime}}$ and $t^{*}$. For each such student, for each intersecting budget constraint $m=1,...,w_{s}$, if $t^{\phi^{\prime}}\in H^{1}(s,x^{m}_{s})$, then the course schedule $x^{m}_{s}$ is unaffordable for student $s$ at $t^{\phi^{\prime}}$. To guarantee that $x^{m}_{s}$ is unaffordable at $t^{*}$ as well, as in Budish (2011), we change the reverse tax $\tau^{\prime}_{s,x^{m}_{s}}$ to be $\tau^{\prime\prime}_{s,x^{m}_{s}}=\tau^{\prime}_{s,x^{m}_{s}}-\delta_{2}$, where $\delta_{2}>0$ is small enough to ensure that conditions (i), (ii), and (iii) from Lemma B1 still hold. If $t^{\phi^{\prime}}\in H^{0}(s,x^{m}_{s})$ (or, for any bundle $x\neq x^{m}_{s}$ for $m=1,...,w_{s}$), set $\tau^{\prime\prime}_{s,x^{m}_{s}}=\tau^{\prime}_{s,x^{m}_{s}}$. Taxes still favor more preferred bundles, i.e., $\tau^{\prime\prime}_{s,x_{s}}>\tau^{\prime\prime}_{sx^{\prime}_{s}}$ if $x^{\prime}_{s}\succ_{s}x_{s}$. Then, letting $x^{*}_{s}=d_{s}(t^{*},b_{s},\tau^{\prime\prime}_{s})$ and setting $b^{*}_{s}=b_{s}+\tau^{\prime\prime}_{s,x^{*}_{s}}$, the price of any bundle preferred by $s$ to $x^{*}_{s}$ costs more than $b^{*}_{s}$.
As a result, with all taxes equal to 0, at budgets $b^{*}$ and priority-specific course prices $p^{*}$ derived from $t^{*}$ using equation (A1), the resulting allocation $x^{*}$ generates excess demand equal to $z(t^{\phi^{\prime}})$, which has a market-clearing error no greater than $\sqrt{kM/2}$. ∎
Appendix C The Student Utility Model Calibration
This appendix describes the method of the simulated moments approach used to calibrate the student utility model. First, Section C.1 outlines the student utility model. Then, Section C.2 discusses adjustments to the university’s course reserves data. The empirical moments used to calibrate the model are described in Section C.3. Sections C.4 and C.5 provide the calibration process, calibration results, and details on model fit.
C.1 Model
We consider a model with student horizontal preferences, course vertical components, and an interaction term. In particular, $u_{sc}$, student $s$’s utility from taking course $c$, takes the following parametric form:
$u_{sc}=Y^{\prime}_{s}\theta+Y^{\prime}_{s}\Gamma H_{c}+z_{c}+\varepsilon_{sc},$ (C1)
where:
  • $Y_{s}$ and $H_{c}$ are binary vectors of student $s$’s observed characteristics and course $c$’s observed characteristics, each with one nonzero element. Denote the set of colleges by ${\cal A}=\{A,...,G\}$ and years of study by ${\cal Y}=\{1,...,4\}$. Then, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ sets the college and year of study of student $s$ and $H_{c}\in\{0,1\}^{|{\cal A}|}$ sets the college of course $c$.
  • $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ and $z=\{z_{c}\}_{c\in{\cal C}}$ are two vectors of unobserved coefficients; $\Gamma$ is a matrix of unobserved coefficients conformable with vectors $Y_{s}$ and $H_{c}$.
  • $\{\varepsilon_{sc}\}_{s\in{\cal S},c\in{\cal C}}$ are random utility components that are assumed to be identically and independently distributed according to the standard normal distribution $\varepsilon_{sc}\sim N(0,1)$.
Each student can be enrolled in up to a maximum of five courses. We assume that the utility from taking multiple courses equals the sum of utilities of taking individual courses. The utility of taking no course is normalized to $u_{s0}=0$.
The parametric form for student utilities in equation (C1) has three non-random components. The first, $Y^{\prime}_{s}\theta$ is a horizontal component that is determined solely by the student’s characteristics. The vector of coefficients $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $28$ elements in total to calibrate. The horizontal component determines the importance of taking courses for student $s$, with a larger value meaning student $s$ wants to enroll in more courses. This accounts for variability in the data in the number of courses taken by students from different colleges and years of study, as shown in Table C.1.
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. $\Gamma$ accounts for the heterogeneity in patterns of course enrollment across colleges. For example, Table C.2 shows students from college $A$ take $71\%$ of courses from their home college, while more than $95\%$ of courses taken by students from college $C$ are courses outside their home college.
The third component $z_{c}$ is a vertical component that determines the attractiveness of course $c$. This component is common among all students.
Table C.1: The average number of courses taken by students.
Year 1 Year 2 Year 3 Year 4
A 3.6 3.4 3.1 2.9
B 4.0 4.3 3.9 3.0
C 4.7 4.3 4.2 3.5
D 4.4 4.3 4.1 3.4
E 4.3 4.3 3.8 3.3
F 4.2 4.0 3.8 2.9
G 4.6 4.7 4.7 4.1
Notes: Each row presents the average number of courses taken by students from the same college and different years of study.
Table C.2: Percentages of courses taken by students across colleges
% A B C D E F G
A 71.4 0.5 1.9 16.9 1.7 5.0 2.7
B 1.9 43.8 0.6 18.4 19.0 12.2 4.1
C 10.2 1.3 4.5 37.2 29.2 15.2 2.4
D 2.5 1.0 0.8 63.9 12.0 13.0 6.8
E 1.2 2.6 0.1 24.1 53.7 16.1 2.1
F 1.5 1.8 0.8 23.3 17.9 52.4 2.3
G 1.3 0.4 0.3 31.3 7.7 6.1 53.0
Notes: This table reports the percentage of courses taken by students across colleges. The rows correspond to students’ colleges, and the columns correspond to courses’ colleges.
Parameter normalization. To correctly interpret the interaction term $\Gamma$ and the vertical component $z$, we need to further normalize the model’s parameters. As described above, for student $s$ from college $a$ and course $c$ from college $a^{\prime}$, we have $Y^{\prime}_{s}\Gamma H_{c}=\gamma_{aa^{\prime}}$. Let $S_{a}$ and $S_{ay}$ be the set of students in college $a$ and the set of students in college $a$ and year $y$, respectively. The average utility of students in college $a$ from taking course $c$ in college $a^{\prime}$ equals
$\overline{u}_{ac}=\frac{1}{|S_{a}|}\sum_{s\in{\cal S}_{a}}u_{sc}=\overline{\theta}_{a}+\gamma_{aa^{\prime}}+z_{c},$
where $\overline{\theta}_{a}=\frac{1}{|S_{a}|}\sum_{y}|S_{ay}|\theta_{ay}$ is the average student-specific component among students in college $a$. We have three additive parameters that lead to the same value of the utility components $\overline{u}_{ac}$ and need a normalization.
One possible normalization is $\overline{\theta}_{a}=0$ and $\gamma_{aa}=0$ for a=1,,7a=1,...,7. This normalization leads to the interpretation of $z_{c}$ as the average popularity of the course among students from the same college. Then, $\gamma_{aa^{\prime}}$ is an increase in student’s utility from taking a course in college $a^{\prime}$ compared to a course in student’s college $a$ with the same vertical component.
Students’ choice sets. The student choice sets are unobserved in the data. We do not track students’ past enrollments, and as a result, we do not know each student’s set of courses for which they satisfy course prerequisites. This complication requires us to take a stance on how students’ choice sets are formed. We assume that each student’s choice set is limited to $80$ courses. The courses in the student’s choice set are drawn randomly such that the probability that a course from college $a^{\prime}$ is drawn equals the share of students from the same college-year who are enrolled in courses in college $a^{\prime}$ in the actual data. Additionally, we make adjustments for courses with course reservations. For each course reservation with $r$ reserved seats, we ensure that the course is included in the choice sets of at least $2r$ students who qualify for the reservation. This adjustment prevents course reservations from artificially limiting a course’s capacity due to an insufficient number of qualified students. [^28] If there are fewer than $2r$ qualified students, the course is included in the choice set of all such students.
Random choice sets are a limitation of our model. Any changes in the way we draw the course lists could potentially change the calibrated parameter values. Crawford, Griffith and Iaria (2021) provide an insightful discussion of preference estimation techniques with unobserved choice sets. Still, the size of the student choice sets we consider is reasonable, as previous studies report that students typically consider only a limited subset of courses in a given semester (see Budish and Cantillon, 2012; Diebold and Bichler, 2017). We explore how the performance of mechanisms change for various sizes of choice sets in Appendix D.4.
C.2 Course Reserve Adjustment
We use the method of simulated moments to calibrate the student utility model parameters. Some adjustments to the course reserves need to be performed first. Course reserves are typically set large at the beginning of the actual course allocation process and relaxed at later stages. Hence, the final course allocation can violate the initial course reserves. In fact, $170$ out of $756$ courses violate initial course reservations in the data.
We next describe the necessary and sufficient conditions for the existence of a matching that assigns a given population of students to a course and satisfies several reservation criteria. We begin by outlining the conditions for two reservation criteria and then extend these conditions to the general case. Afterward, we describe the process of adjusting course reserves to ensure that the number of reserved seats is consistent with the course seat assignment observed in the data.
Two course reserve criteria. Consider a course with a capacity of $q$ seats. Denote the number of course reserves for students who satisfy criterion $A_{i}$ (student’s department and year of study) as $\mathfrak{r}_{A_{i}}$, $i=1,2$. We assume that $\mathfrak{r}_{A_{1}}\geq 0$ or $\mathfrak{r}_{A_{2}}\geq 0$ and $\mathfrak{r}_{A_{1}}+\mathfrak{r}_{A_{2}}\leq q$. One should interpret the reservation such that if a student satisfies both $A_{1}$ and $A_{2}$, she is eligible to take a seat satisfying either criterion, but if she takes $A_{1}$-type seats it does not decrease the number of $A_{2}$-type seats. This interpretation is consistent with the actual course enrollment process in the data. The number of students who satisfy criterion $A_{i}$ and who are matched to the course in the final allocation is denoted by $x_{A_{i}}$, $i=1,2$. Students who do not satisfy any course reserve criteria are called 0-students, and we denote the number of such students assigned to the course as $x_{0}$. Then, the assignment satisfies course reserves if and only if
$\left\{\begin{array}[]{l}x_{0}+x_{A_{1}\cup A_{2}}\leq q\\[-2.84526pt] x_{0}\leq q-\mathfrak{r}_{A_{1}}-\mathfrak{r}_{A_{2}}\\[-2.84526pt] x_{0}+x_{A_{1}\backslash A_{2}}\leq q-\mathfrak{r}_{A_{2}}\\[-2.84526pt] x_{0}+x_{A_{2}\backslash A_{1}}\leq q-\mathfrak{r}_{A_{1}}\end{array}\right.$. (C2)
To see that the above conditions are necessary, we consider the criteria one by one. The first inequality corresponds to the criterion that puts no restrictions on students, the second inequality corresponds to students who do not satisfy $A_{1}$ and $A_{2}$, the third corresponds to students who do not satisfy $A_{2}$, and the fourth corresponds to students who do not satisfy $A_{1}$. To show sufficiency, we first enroll students $x_{0}$, $x_{A_{1}\backslash A_{2}}$, and $x_{A_{2}\backslash A_{1}}$. It is clearly possible to do so satisfying all criteria. The remaining students $x_{A_{1}\cap A_{2}}$ are constrained by the total course capacity. The first inequality guarantees that they could be enrolled as well.
The general case. The number of constraints increases exponentially with the number of separate seat reservations a course has. However, there are 4 or fewer seat reservations per course in the university data, which makes the analysis computationally manageable. Below, we present the necessary and sufficient conditions that guarantee the existence of a matching that satisfies multiple course reserve criteria that are based on the seminal Hall’s Marriage Theorem (see Hall, 1935).
To state the theorem, let us consider a set of students $S$ and course reserves criteria $A_{1},...,A_{I}$ for some $I\geq 1$. The number of reserved seats corresponding to criterion $A_{i}$ equals $\mathfrak{r}_{A_{i}}\geq 0$. We denote the set of available course seats as $Y$ with $|Y|=q$ and assume $\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}\leq q$. Each student could potentially satisfy multiple criteria. We consider a bipartite graph $(S,Y,E)$ where $E$ is the set of edges connecting student and seats. A student $s\in{\cal S}$ is connected to seat $y\in Y$ with an edge if student $s$ is eligible to take seat $y$. For example, if there are no course reserves, each student is connected with an edge to each seat ($q$ edges). If there is one course reservation criterion $A$, each 0-student has $q-\mathfrak{r}_{A}$ edges, and each student who satisfies criterion $A$ has $q$ edges.
For a subset of students $S^{\prime}\subseteq S$, let $N(S^{\prime})$ denote the set of seats in $Y$ that are connected to at least one student in $S$. We call these seats the neighbors of $S^{\prime}$. A perfect matching is a matching between students and seats in which each student is assigned exactly one seat.
Theorem (Hall’s Marriage Theorem).
The necessary and sufficient condition for the existence of a perfect matching is that every subset $S^{\prime}\subseteq S$ is connected to as many neighbors as the number of its elements
$|S^{\prime}|\leq|N(S^{\prime})|.$ (C3)
It is not hard to see that the conditions of Hall’s Marriage Theorem transform to the system of inequalities (C2) when there are only two criteria. To obtain the necessary and sufficient constraints in the general case, we need to consider only those inequalities where for a fixed set of seats $N(S^{\prime})$ there is the maximum possible $|S^{\prime}|$.
  • The number of students who satisfy no criteria, that is, 0-students
    $x_{0}\leq q-\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}.$
  • The number of students who satisfy only criterion $A_{i}$ or no criteria (adding 0-students only increases the number on the left-hand side of inequality (C3) and does not change the number on the right-hand side of the inequality)
    $x_{A_{i}\backslash\cup_{j\neq i}A_{j}}+x_{0}\leq q-\sum_{j\neq i}\mathfrak{r}_{A_{j}}$
[…]
    $x_{\cup_{i\in{\cal I^{\prime}}}A_{i}\backslash\cup_{j\notin{\cal I^{\prime}}}A_{j}}+x_{0}\leq q-\sum_{j\notin{\cal I^{\prime}}}\mathfrak{r}_{A_{j}}.$ (C4)
[…]
Course reserve adjustment. Using the above conditions, we have identified 170 out of 756 courses that violate the initial course reserve criteria in the university data. For these courses, we order course seat reservations by year of study in decreasing order. We then take the first course reserve and decrease the number of its reserves by one. If the Hall’s inequalities are still violated, we move to the next one and also decrease the number of its reserves by one. If decreasing all course reserves by one does not make matching possible, we repeat this procedure. Clearly, the procedure terminates in a finite number of steps, as a matching with zero course reserves always possible. [^29] For over-enrolled courses, we first set $q$ equal the maximum between course capacity and the number of students assigned in the university data.
C.3 Moment Conditions
For each set of parameter values and realization of random components of the student utility model described in Section C.1, we use the Random Serial Dictatorship with adjusted course reserves to simulate the allocations of courses to students. Then, we use these allocations to calculate the simulated moments specified below.
For each horizontal component of students’ preferences $\theta_{ay},a\in{\cal A},y\in{\cal Y}$, we consider the number of courses taken by students with the same characteristics as a moment, i.e.,
$\breve{m}_{ay}=\sum_{s\in{\cal S}_{ay}}\sum_{c\in{\cal C}}x_{s,c},$
where $x_{s,c}\in\{0,1\}$ indicates student $s$’s enrollment into course $c$ and $S_{ay}$ is the set of students from college $a$ and year $y$.
Similarly, the number of students from college $a$ who enrolled in courses in college $a^{\prime}$ is a natural moment for the interaction term $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$; that is,
$\hat{m}_{aa^{\prime}}=\sum_{s\in{\cal S}_{a}}\sum_{c\in{\cal C}_{a^{\prime}}}x_{s,c},$
where $C_{a^{\prime}}$ is the set of courses in college $a^{\prime}$ and $S_{a}$ is the set of students from college $a$.
For the vertical component of course-specific qualities $z_{c},c\in{\cal C}$, it is natural to consider the course enrollment as a moment, i.e.,
$\tilde{m}_{c}=\sum_{s\in{\cal S}}x_{s,c}.$
Courses at maximum capacity. A critical aspect of the data is that $85$ out of $756$ courses are at maximum capacity in the data, which creates a problem with identifying course-specific parameters $z_{c}$ for such courses. To address this issue, we replace moments $\tilde{m}_{c}$ for such courses with moments based on the timing of course enrollment. We first define a ranking of students.
Definition C1.
The rank of student $s$, $\text{rank}_{s}$, is her position in the order of students to enroll in the Random Serial Dictatorship with course reserves after breaking ties.
Hence, more popular courses are filled up earlier with students of a smaller rank. Less popular courses are filled up later with students of a larger rank. [^30] More senior students have earlier time slots in the registration process. Hence, measuring popularity by how fast a course fills up is biased towards the popularity of courses among more senior students.
To define student ranking in the university data, we considered the earliest timestamp of each student in our dataset; that is, the time students registered for their first course. Using these timestamps, we rank students in the university data. [^31] An alternative measure of how fast courses fill up is course registration times. Both student ranking and course registration time are closely related in our simulations, as each student registers for all desired courses simultaneously. This connection is less evident in the data. Many students do not register for courses at the same time. This choice pattern is probably connected with students changing their minds or new possibilities arising during the actual course registration process. We found that student ranking better connects our simulations to the actual university data than course registration time. To measure course popularity, we use the mean student rank enrolled in the course:
$Mean(\{\text{rank}_{s}:x_{s,c}=1,\,s\in{\cal S}\}).$
For courses at maximum capacity, we replace moments $\tilde{m}_{c}$ with the mean rank across the registered students. [^32] We also tried to use the largest student rank and some quantiles of students registered for the course to measure course popularity. These measures did not work very well. Last-minute course registrations are a noisy measure that do not reflect course popularity well. Measures based on quantiles proved to be discontinuous, with a small change in utility parameters often resulting in jumps in the student ranks at a specific quantile and a bad model fit. We keep using the symbol $\tilde{m}_{c}$ for all moments corresponding to vertical components $z$ to avoid clutter in notation. Overall, there are $756+28+49=833$ moments m=(m~,m˘,m^)m=($\tilde$$\breve{m}$$\hat${m}) and $819$ independent parameters $\pi=(\theta,\Gamma,z)$ to calibrate. Hence, the model is over-identified, allowing us to test for model fit in Section C.5.
C.4 Calibration
We use the Random Serial Dictatorship with adjusted course reserves to calculate simulated moments and find $\pi^{*}=(\theta^{*},\Gamma^{*},z^{*})$ that minimize a criterion function:
$m-m^{s}(\pi)||_{W}^{2}=(m-m^{s}(\pi))^{\prime}W(m-m^{s}(\pi)),$ (C5)
where $m$ is a set of moments constructed from the empirical data, $m^{s}(\pi)$ is the average of moments constructed from $100$ simulation rounds for a given vector $\pi=(\theta,\Gamma,z)$ (drawing random student utility components and choice sets for each round), and $W$ is a weight matrix. We use the following procedure (see Jalali, Rahmandad and Ghoddusi, 2015):
  • First Stage. We calibrate the optimal parameters $\pi$ using an initial weight matrix $W_{0}$. $W_{0}$ is a diagonal matrix with diagonal element $d$ equal to $1/m^{2}_{d}$, where $m_{d}$ is the value of the moment constructed from the empirical data. Hence, the first stage minimizes the sum of percentage errors for each moment.
    Using estimated $\pi$, we calculate the variance-covariance matrix of simulated moments
    $\hat{S}=\frac{1}{L_{1}}\sum^{L_{1}}_{\ell_{1}=1}\left[m^{\ell_{1}}(\pi,\varepsilon^{\ell_{1}})-\frac{1}{L_{2}}\sum^{L_{2}}_{\ell_{2}=1}m^{\ell_{2}}(\pi,\varepsilon^{\ell_{2}})\right]\left[m^{\ell_{1}}(\pi,\varepsilon^{\ell_{1}})-\frac{1}{L_{2}}\sum^{L_{2}}_{\ell_{2}=1}m^{\ell_{2}}(\pi,\varepsilon^{\ell_{2}})\right],$ (C6)
    where $\varepsilon^{\ell_{1}}$ and $\varepsilon^{\ell_{1}}$ are draws of random component for $\ell_{1}=1,...,L_{1}$ and $\ell_{2}=1,...,L_{2}$. Following best practices, we take $L_{1}=L_{2}=1000$ (see Adda and Cooper, 2003). The optimal weight matrix is the inverse of the variance-covariance matrix $W^{*}=\hat{S}^{-1}.$
  • Second Stage. We start with the optimal weight matrix $W^{*}_{\ell}=W^{*}$ for $\ell=1$. We then use the optimal weight matrix to obtain more precise estimates for parameters $\pi^{*}_{\ell}$. Using the more precise estimate $\pi^{*}_{\ell}$, we compute the optimal weight matrix $W^{*}_{\ell+1}$ and repeat the process $\ell=2,3,$... until the maximum change of calibrated parameters $|\pi^{*}_{l+1}-\pi^{*}_{l}|$ becomes smaller than $0.01$ to ensure convergence.
Calibration results. Table C.3 provides the calibrated parameters of the student utility model. The top portion of the table displays the horizontal components $\theta$, with rows corresponding to colleges and columns corresponding to year of study. The coefficient $\theta_{ay}$ represents the overall importance of taking courses for students in college $a$ and year $y$. Our normalization ensures that the sum of the coefficients in each row, weighted by the share of the student population across years, equals zero.
The horizontal components are positive for students in their first and second years and mostly negative in their third and fourth years. This observation suggests that students prefer enrolling in more courses during their initial years of study. The tendency to enroll in fewer than five courses in later years may be due to several factors. Such students may choose part-term courses that are not in our data or have already accumulated enough credits for their degree, allowing them to ease their course load in the final years.
The middle portion of the table presents interaction terms of matrix $\Gamma$. The general trend of negative utility changes emphasizes a preference for internal college courses, but specific positive changes indicate potential areas for beneficial inter-college course offerings. For example, students from colleges $C$ and $D$ value highly courses in college $A$.
The bottom portion of the table presents quantiles of $756$ coefficients of vertical component $z$, with each coefficient $z_{c}$ being responsible for the average popularity of the course among students from the same college. The values range from a minimum of 2.32-2.32 to a maximum of $2.28$. This wide range indicates significant variation in the popularity of courses within colleges. Most of the quantile values are negative, up to the $90th$ percentile, indicating that a majority of the courses have negative values. This result is most likely manifested by the student utility model that relies on college-year parameters to identify student preferences. Students randomly draw $80$ courses in their choice set but take no more than $5$ courses. Our model, which relies on college-level and year-level parameters, matches this empirical pattern by putting the values of the vertical components of most courses in the negative region.
C.5 Model Fit
We test for model fit using the calibrated parameters. Our student utility model is over-identified, as we have $833$ moments and $819$ independent parameters. The $J$-test statistic is given by
$J=\frac{K}{K+1}(m-m^{s}(\pi^{*}))^{\prime}W^{*}(m-m^{s}(\pi^{*}))\sim\chi^{2}_{14}.$
where $W^{*}$ is the variance-covariance matrix of moments estimated at the terminal step of stage $2$ in the calibration process. This statistic has chi-square distribution with $14$ degrees of freedom under the null hypothesis that the model fits the data. The $J$-test statistic is $J=2.42$, which is less than the critical value of $29.14$ at $99\%$ cut-off level. Hence, the $J$-test does not reject the null hypothesis that the model fits the data at $99\%$ confidence level.
Table C.3: The calibrated parameters of the student utility model.
Coefficients $\theta$
Year 1 Year 2 Year 3 Year 4
A 0.12 0.20 -0.04 -0.27
B 0.13 0.19 0.01 -0.31
C 0.28 0.21 0.01 -0.44
D 0.09 0.19 -0.03 -0.32
E 0.20 0.15 -0.13 -0.29
F 0.17 0.08 -0.09 -0.28
G 0.19 0.11 0.01 -0.37
Coefficients $\gamma$
A B C D E G F
A 0.00 -0.65 -0.58 -0.28 -0.55 -0.70 -0.52
B -0.11 0.00 -0.54 -0.24 -0.09 -0.46 -0.48
C 0.37 -0.22 0.00 -0.01 0.01 -0.28 -0.26
D 0.14 -0.12 -0.39 0.00 -0.16 -0.32 -0.27
E 0.02 -0.55 -0.34 -0.17 0.00 -0.33 -0.40
F -0.07 -0.65 -0.57 -0.21 -0.17 0.00 -0.55
G -0.19 -0.56 -0.58 0.04 -0.20 -0.49 0.00
The Quantiles of Coefficients $z$
Quantile Min 0.10 0.25 0.50 0.75 0.90 Max
$z_{c}$ -2.32 -1.89 -1.70 -1.47 -1.12 -0.67 2.28
Notes: This table provides information about the calibrated parameters of the student utility model. The upper portion of this table presents the horizontal components $\theta$ of students’ utility, with the rows corresponding to the student’s college and the columns corresponding to the student’s year of study. Our normalization ensures that the sum of the parameters in each row weighted with the share of student population across years equals zero. The middle portion presents interaction terms of matrix $\Gamma$. Each non-zero interaction coefficient $\gamma_{aa^{\prime}},a,a^{\prime}\in{\cal A}$ corresponds to the increase in student’s utility from taking a course in college $a^{\prime}$ (column) compared to a course in student’s college $a$ (row) with the same vertical component. Note that we normalize $\gamma_{aa}=0,a\in{\cal A}$. The bottom portion presents some quantiles of vertical components $z$. The interpretation of $z_{c}$ is the average popularity of the course among students within the course’s college.
Appendix D Additional Simulations
This appendix provides additional simulation results. First, for the simulations conducted in the main text, we present additional results on the equilibrium outcomes of the Pseudo-Market with Priorities (PMP) mechanism (Section D.1) and compare performance with some alternative mechanisms (Section D.2). Second, we supplement the main text with simulation results with a department-first priority structure (Section D.3). Last, we report robustness checks by considering several sizes for student choice sets and the standard deviation of the noise utility parameter (Section D.4).
D.1 Outcomes of the Pseudo-Market with Priorities Mechanism
First, we present statistics on the equilibrium prices in the PMP mechanism for the simulations reported in Section 3.3. The top part of Table D.1 provides information about where the cutoff levels of priority are across courses and how high the prices are for students at the cutoff level of priority. The bottom part provides information about the percentage of students within each year of study allocated different numbers of courses with positive prices.
Around $70.8\%$ of courses have their cutoff priority at level $1$. The large share of courses at the lowest priority level reflects the non-binding enrollment of most courses in the university data, where just $11.2\%$ of courses are at or above their maximum enrollment capacity (see Table 2 on p. 2). Above level $1$, there is a noticeable difference in the percentage of courses with odd and even cutoff levels of priority. Substantially more courses have odd cutoff levels ($8.0\%$, $8.5\%$, $5.9\%$ for levels $3$, 55, and $7$, respectively) than even ones ($2.1\%$, $2.6\%$, $1.7\%$, $0.5\%$ for levels $2$, 44, 66, and $8$). The hybrid priority structure we utilize places students at odd priority levels in a course if they are ineligible for any of its reserved seats. The students eligible for a course’s reserved seats typically only comprise a small subset of the interested students. Furthermore, some courses do not reserve any seats. As a result, the cutoff level of priority often ends up at an odd level rather than an even one.
Table D.1: Price statistics for the Pseudo-Market with Priorities mechanism.
Cutoff levels of priority and their average prices
Cutoff level of priority 11 22 33 44 55 66 77 88
% of courses 70.8% 2.1% 8.0% 2.6% 8.5% 1.7% 5.9% 0.5%
Average cutoff price $0.06$ $0.76$ $0.83$ $0.75$ $0.88$ $0.80$ $0.94$ $0.81$
Courses with positive prices taken by students
# of courses paid for Year 1 Year 2 Year 3 Year 4
0 42.6% 36.7% 44.0% 59.01%
1 45.2% 51.2% 45.4% 37.0%
2 10.9% 10.9% 9.7% 3.8%
3 1.2% 1.1% 0.9% 0.2%
4 0.05% 0.06% 0.03% 0.006%
5 0.001% 0.002% 0.001% 0%
Notes: This table presents statistics on the equilibrium prices in the Pseudo-Market with Priorities mechanism. Results are averages across $100$ runs with different random component draws.
Among courses with a cutoff priority level of $1$, the average price at the cutoff is just $0.06$. As many courses are free for all students, the average cutoff price for these courses is considerably lower than for courses with a higher cutoff level. For higher cutoff levels, the average cutoff prices lie between $0.75$ and $0.94$, with higher prices for odd cutoff levels. Student budgets are assigned in $[1,1.25]$, meaning that, at the average cutoff price, each student can afford one but not two courses for which they are at the cutoff level of priority. Still, students manage to take $k=5$ courses by taking courses for which they are above the cutoff level of priority, where they face a priority-specific price of 0.
The bottom part of Table D.1 further explores the number of courses students pay a positive price for. Across all years of study, at least $36\%$ of students obtain all their course seats for free, and fewer than $13\%$ of students pay a positive price for more than one course. Many courses have a zero price for all students, leading some students to obtain their most preferred course schedule for free. In addition, students often cannot afford a second course with a positive price after obtaining one.
These findings differ across years of study. Less than $7\%$ of cutoff levels of priority are $7$ or $8$, implying that Year $4$ students face zero price for the vast majority of courses. Almost $60\%$ of these students pay for none of their assigned courses, with about $4\%$ obtaining multiple course seats with positive prices.
In earlier years of study, a larger percentage of students pay for course seats. Under $45\%$ of first- through third-year students obtain all their course seats for free, with more than $10\%$ needing to pay for multiple courses in each year.
Table D.2: Over-enrolled courses in the Pseudo-Market with Priorities mechanism.
# seats above max capacity $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses $2.5$ $0.4$ $0.1$ $0.02$ $0.0$
Notes: This table provides information about the distribution of over-enrolled courses in the Pseudo-Market with Priorities mechanism. Results are averages across 100 runs with different random component draws.
We also present results on course over-enrollment in the PMP mechanism’s outcomes. The average market-clearing error in our simulations is $21.4$, less than half of the theoretical worst-case bound of $\alpha\approx 43.5$. Table D.2 reports that $2.5\%$, or roughly 19 courses, are over-enrolled by one seat. In addition, $0.4\%$ (3 courses) are over-enrolled by two seats, and $0.12\%$ (<<1 course) are over-enrolled by at least three seats. No courses are over-enrolled by more than four seats. In contrast, the university’s original allocation has $7.3\%$ of courses over-enrolled, with $1.9\%$ over-enrolled by at least five seats (see Table 2). A lower market-clearing error and over-enrollment are possible by allowing a longer runtime for our algorithm.
D.2 Alternative Course Allocation Mechanisms
Our main simulation results compare the Pseudo-Market with Priorities mechanism and two variants of the Deferred Acceptance mechanism, each of which respects the course priority structure. That is, the outcomes of these mechanisms contain no student who wants to enroll in a course that assigns a seat to a student of strictly lower priority.
In turn, the Random Serial Dictatorship with course reserves (RSD) benchmark does not respect any priority structure. The cause of such violations is the presence of course reserves that can be interpreted as seat-specific priorities. Each of a course’s reserved seats has a priority order over students that places reserve-eligible students above others in the same year of study. Unreserved seats only prioritize students by year of study. Other mechanisms in the literature have seat-specific priorities. For example, the Deferred Acceptance with minority reserves mechanism introduced in Hafalir, Yenmez and Yildirim (2013) is a modification of the Deferred Acceptance mechanism with seat-specific priorities.
The main objective of this section is to compare the performance of the PMP mechanism with six alternative mechanisms. Four of these mechanisms also have seat-specific priorities (DA-STB and DA-MTB with minority reserves, RSD with university course reserves, Probabilistic Serial with Seniority and Reserves), one does not respect any priorities (A-CEEI), and one respects priorities in a kludgy way (A-CEEI with kludgy pricing).
Deferred Acceptance (DA) with Minority Reserves with single and multiple tie-breakings. The Deferred Acceptance with minority reserves mechanism was proposed by Hafalir, Yenmez and Yildirim (2013) to address affirmative action policies in school choice. This mechanism is similar to the Deferred Acceptance mechanism, but it allows each school to have a number of reserved seats designated for affirmative action students. These seats give higher priority to minority students up to the point that minorities fill the reserves. We extend this mechanism to a many-to-many matching setting with indifferences, using the optimal course reserves discussed in Section 3.2. We consider two cases: where indifferences are broken with a single tie-breaking rule and where indifferences are broken with multiple tie-breaking rules, one for each course. [^33] We do not claim that the theoretical properties of Hafalir, Yenmez and Yildirim (2013) extend to many-to-many settings with indifferences. We leave the analysis of this important question for further research.
Random Serial Dictatorship (RSD) with Actual Course Reserves. Course reserves are typically set large at the beginning of the course allocation process at major US universities and relaxed at later stages. In the main text, motivated by this observation, we estimate the optimal reserved seats for each course in the RSD mechanism. Here, we run the same mechanism, using the university’s actual course reserves.
Approximate Competitive Equilibrium from Equal Incomes (A-CEEI). This mechanism was originally proposed by Budish (2011) to allocate courses in business schools. It is a special case of the PMP mechanism in which all courses place all students at the same priority level, leading each course to have a single price. It shares many favorable theoretical properties with the PMP mechanism.
Approximate Competitive Equilibrium from Equal Incomes with Kludgy Pricing (A-CEEI with Kludgy Pricing). This variant of the A-CEEI mechanism aims to satisfy course priorities by using “kludgy pricing”. As in the A-CEEI mechanism, this mechanism uses one price $p_{c}$ per course $c$. However, students receive a price discount depending on their priority. A student $s$ at level of priority $r_{s,c}$ in course $c$ faces price $p_{c}\cdot\left(1-(r_{s,c}-1)/R\right)$, where $R$ is the largest possible priority level. This pricing ensures that students always face a lower price than students at a lower level of priority but leaves open the possibility of a lower-priority student taking a seat when a higher-priority student cannot afford one.
Probabilistic Serial (PS) with Seniority and Reserves. The PS mechanism was proposed by Bogomolnaia and Moulin (2001) to randomly assign objects to agents in settings with single-unit demand for objects and single-unit supply of object types. Kojima (2009) extended the mechanism to settings with multi-unit demand for objects, and Budish et al. (2013) extended the mechanism to settings with multi-unit supply of objects. The undergraduate course allocation setting requires both multi-unit demand and multi-unit supply of courses. The only paper that considers both multi-unit demand and multi-unit supply of courses with preferences over individual courses is an unpublished paper by Pycia (2011). [^34] Nguyen, Peivandi and Vohra (2016) also consider multi-unit demand and multi-unit supply of objects but consider agent preferences over bundles of objects. They propose a version of a probabilistic serial mechanism called the Bundled Probabilistic Serial mechanism that has been successfully implemented at the Technical University of Munich by Bichler and Merting (2021). One of the versions of their probabilistic serial mechanism could also be applied to settings with agent preferences over individual objects. The setting of undergraduate course allocation is further complicated by the presence of student seniorities and course reserves. We define an extension of the probabilistic serial mechanism called Probabilistic Serial with Seniority and Reserves, using the optimal course reserves discussed in Section 3.2.
In this mechanism, each course seat is regarded as a divisible object of probability shares. We start with the most senior students (e.g., Year 4), each of whom “eats” with speed one the most preferred eligible course seat that is available. Each student is eligible for unreserved seats and each reserved seat specifying her department and year of study. We assume that students first consume reserved course seats before consuming unreserved ones of a particular course. If some reserved/unreserved seats are exhausted, or some student has consumed a full probability share of a certain course’s seat, we stop and redetermine the student’s most preferred eligible course to eat next. This proceeds until time $k$, where the most senior students have consumed up to $k$ courses. We then repeat the procedure with the students of the next seniority level (e.g., Year 3, Year 2, and Year 1). The resulting profile of shares of course seats eaten by agents by time $4k$ corresponds to a random assignment.
Figure D.1: Each student’s preferred mechanism for the alternative mechanisms.
Notes: This figure reports the percent of students who strictly prefer each alternative mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark. Results are averages across 100 runs with different random component draws.
Simulation results. Our simulation results for the six alternative mechanisms are presented below. Figure D.1, as in Figure 1 in the main text, reports the percentage of students who prefer the allocation of a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who prefer the benchmark. Figure D.2, as in Figure 2, reports the percent improvement in mean utility among students with changing schedules for the six mechanisms. [^35] On average, the DA-STB with minority reserves mechanism changes schedules for 775, 923, 698, and 319 students in Years 1, 2, 3, and 4, respectively; the DA-MTB with minority reserves mechanism does so for 918, 1071, 845, 576 students; the RSD mechanism with actual course reserves does so for 1176, 1131, 1135, and 774 students; the A-CEEI mechanism does so for 1437, 1413, 1102, and 898 students; the A-CEEI with kludgy pricing mechanism does so for 1256, 1131, 951, and 562 students; and the PS mechanism does so for 1078, 1302, 1080, and 838 students. Figure D.3, as in Figure 3, presents the change in the standard deviation of student utility for each mechanism by year of study. Table D.3, parallel to Table 5, reports students who experience schedule envy towards students of the same priority or lower priority. Table D.4 presents the percentage of students who experience a priority violation: who want to enroll in a course being taken by a student of strictly lower priority. The PMP, DA-STB, and DA-MTB mechanisms studied in the main text each have no priority violations.
Figure D.2: Improvements in utility among students with changing schedules for the alternative mechanisms.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each alternative mechanism over the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
The DA with minority reserves (with single and multiple tie-breakings) mechanisms each have a smaller percentage of students who prefer the mechanism than students who prefer the benchmark. This negative result contrasts with the result for their deferred acceptance counterparts, as studied in Section 3. These mechanisms only benefit Year 1 students, who significantly improve in mean utility ($39.1\%-46.9\%$ among students with changing schedules). These results are a direct consequence of the presence of course reserves, which ensure better access to course seats for junior students. Course reserves do not provide equal access to seats, leading to an increase in the standard deviation of student utility for junior students ($6.0\%-14.7\%$) compared to the RSD benchmark. We see similar underperformance in other measures of allocation fairness. In both mechanisms, more than $20\%$ of students experience schedule envy toward a student of weakly lower priority. Furthermore, more than $40\%$ of students face a priority violation, indicating that many students can benefit from a course seat assigned to a student of lower priority.
Figure D.3: The change in the standard deviation of students’ utilities for the alternative mechanisms.
Notes: This figure reports the percent change in the standard deviation of students’ utilities for each alternative mechanism compared to the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
The RSD with actual course reserves mechanism performs quite poorly compared to the benchmark. More than $50\%$ of students strictly prefer the RSD mechanism with optimal course reserves over the RSD mechanism with actual course reserves. Using the actual course reserves only improves outcomes for first-year students, who have greater access to reserved seats. The mechanism underperforms in terms of allocation fairness as well, with $27.8\%$ of students preferring the schedule of a student with a weakly lower priority in each course and with $59\%$ of students experiencing priority violations. These results justify why we consider the RSD mechanism with optimal course reserves as the benchmark mechanism rather than the RSD mechanism with actual course reserves. Using the actual course reserves presents an unfair comparison, as the university sets reserves large at the beginning of the process and relaxes them later. This results in an allocation of courses to students, which, to a large extent, happens after the main registration process in practice.
Table D.3: The percentage of students who experience schedule envy for the alternative mechanisms.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
DA with minority reserves (single) 78.9 (0.5) 18.0 (0.5) 2.7 (0.2) 0.4 (0.2) 0.02 (0.03) 0.0007 (0.01)
DA with minority reserves (multiple) 79.5 (0.5) 18.9 (0.6) 1.6 (0.2) 0.1 (0.03) 0.003 (0.01) 0.0 (0.0)
RSD with actual course reserves 72.2 (0.5) 23.9 (0.5) 3.6 (0.2) 0.3 (0.1) 0.01 (0.02) 0.0 (0.0)
A-CEEI 95.1 (0.4) 4.9 (0.4) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
A-CEEI with kludgy pricing 96.3 (0.3) 3.7 (0.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Probabilistic Serial with reserves 89.1 (0.6) 9.9 (0.6) 0.9 (0.2) 0.05 (0.03) 0.0007 (0.003) 0.0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the alternative mechanisms. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across 100 runs with different random component draws.
The A-CEEI and A-CEEI with Kludgy Pricing mechanisms do not strictly enforce course priorities through prices, unlike the PMP mechanism. These relaxed restrictions result in more course seats available to junior students (Years 1 and 2), who overwhelmingly prefer each mechanism to the RSD benchmark (see the bottom section of each bar in Figure D.1). While the results are reversed for senior students (Years 3 and 4), many more students overall prefer each of these mechanisms ($48.4\%-54.8\%$) than the number of students who prefer RSD ($19.7\%-25.7\%$). Both mechanisms reduce the standard deviation of student utility across all years of study ($8.5\%-22.5\%$). This decrease is substantially larger than that observed in Section 3 for the PMP mechanism, suggesting that the relaxed priority constraints allow students to receive more equal access to courses. Similarly, over $95\%$ of students experience no envy toward a student of weakly lower priority, with schedule envy bounded by a single course in each mechanism. However, without these constraints, priority violations are common in both mechanisms, with $61.1\%$ of students in the A-CEEI mechanism and $44.6\%$ of students in the A-CEEI with Kludgy Pricing mechanism whose schedule would benefit from a course seat taken by a student of strictly lower priority. Additionally, the average time to calculate the allocation for the A-CEEI and A-CEEI with Kludgy Pricing mechanisms (2003 and 1894 seconds, respectively) is almost twice that of the PMP mechanism (1112 seconds).
Table D.4: The percentage of students with a priority violation for the alternative mechanisms.
DA with minority reserves (single) DA with minority reserves (multiple) RSD with actual reserves A-CEEI A-CEEI with kludgy pricing PS w. seniority and reserves
46.1% 47.3% 59.0% 61.1% 44.6% 81.0%
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority for each of the alternative mechanisms. Results are averages across $100$ runs with different random component draws.
The PS mechanism is the only mechanism considered here that produces random assignments. Random mechanisms are more flexible, allowing the assignment of probability shares of course seats to students rather than full seats. The PS mechanism, however, delivers more beneficial outcomes only to Year $1$ students. More second- through fourth-year students prefer their schedule in the RSD benchmark than the number of students who prefer their PS outcome. These outcomes are also reflected in utility improvements among students with changing schedules. There is a $31.8\%$ improvement in average utility for Year 1 students. In contrast, all the other student cohorts experience a decrease in average utility (see Figure D.2). In terms of allocation fairness, the probabilistic serial mechanism reduces the standard deviation of students’ utilities for all years of study and results in just $10.9\%$ of students with envy towards a student of weakly lower priority. [^36] Kojima (2009) and Budish et al. (2013) prove that the probabilistic serial mechanism results in no envy in multi-unit demand and multi-unit supply settings, respectively. This result does not extend to settings where agents demand multiple units of objects and objects are supplied in multiple units simultaneously. In doing so, the mechanism results in a large percentage of students who experience priority violation ($84.6\%$). This can be explained, in part, by the flexibility of the mechanism. We count a student as experiencing a priority violation if there is a positive probability share of a course that she prefers to some assigned (possibly empty) probability share that is instead assigned to some student at a strictly lower priority level. This matches the definition of ex-ante stability proposed in Kesten and Ünver (2015). We also consider envy in terms of the ex-ante assignment, not an ex-post assignment. Hence, the results are not directly comparable.
D.3 Department-First Priorities
Some U.S. universities consider department-specific priorities as more important than those based on year of study. For example, Dartmouth College explicitly states these priorities on its website. [^37] See https://www.dartmouth.edu/reg/registration/course_priorities.html. Here, we present additional simulation results comparing the performance of all considered mechanisms under a department-first priority structure, where the department-specific priority takes precedence over the year-specific priority. We modify Example 2 from Section 3 to provide an illustration.
Example D1.
The course reservation for Real Analysis in Table 4 results in the following department-first priority structure for the course:
  • $r_{s,c}=8$: 4th-year students in Dept 2
  • $r_{s,c}=7$: 3rd-year students in Dept 2
  • $r_{s,c}=6$: $\varnothing$
  • $r_{s,c}=5$: $\varnothing$
  • $r_{s,c}=4$: 4th-year students in all departments except Dept 2
  • $r_{s,c}=3$: 3rd-year students in all departments except Dept 2
  • $r_{s,c}=2$: 2nd-year students in all departments
  • $r_{s,c}=1$: 1st-year students in all departments
Figure D.4: Each student’s preferred mechanism under the department-first priority structure.
Notes: This figure reports the percent of students who strictly prefer each given mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark under the department-first priority structure. Results are averages across 100 runs with different random component draws.
With a department-first priority structure, the Deferred Acceptance mechanism coincides with the Deferred Acceptance mechanism with minority reserves of Hafalir, Yenmez and Yildirim (2013). As in Example D1, this is because students eligible for reserved seats receive the highest priority for all course seats. Hence, we only consider the Deferred Acceptance mechanisms with single and multiple tie-breakings in our simulations. We also drop the analysis of Random Serial Dictatorship with actual course reserves, as its performance is subpar compared to all the other mechanisms. The results of our simulations are presented in Figures D.4D.6 and Tables D.5D.6.
Figure D.5: Improvements in utility among students with changing schedules under the department-first priority structure.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each given mechanism over the Random Serial Dictatorship with optimal course reserves benchmark under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
Figure D.4 displays the results on each student’s preferred mechanism under the department-first priority structure. We can observe that an aggregate relative comparison is similar to the results for the hybrid priority structure (Figures 1 and D.1). In particular, there is a larger percentage of students that strictly prefer each of the given mechanisms than students who prefer the RSD benchmark. The only exception is the PS mechanism, which is consistent across both priority structures.
As opposed to the hybrid priority structure, the department-first priority structure has all students eligible for reserved seats prioritized above more senior but ineligible students. For PMP, DA-STB, and DA-MTB, this preferential treatment benefits more junior students (Years $1$ and $2$) and harms senior students (Years $3$ and $4$). This is illustrated in Figure D.5 by the magnitudes in improvements in utility among students with changing schedules. [^38] On average, the PMP mechanism changes schedules for 317, 414, 466, and 397 students in Years 1, 2, 3, and 4, respectively; the DA-STB mechanism does so for 198, 216, 196, and 169 students; the DA-MTB mechanism does so for 468, 586, 599, and 527 students; the A-CEEI mechanism does so for 1313, 1412, 1149, and 904 students; the A-CEEI with kludgy pricing mechanism does so for 1140, 1254, 1008, and 753 students; and the PS mechanism does so for 772, 1044, 992, and 848 students. The results are less sharp for the A-CEEI and A-CEEI with kludgy pricing mechanisms. These do not respect course priorities, leading to the benefit being spread across years of study. We also observe the opposite effect for the PS mechanism: an improvement in utility is larger for more senior students compared to those who are more junior.
Figure D.6 reports each mechanism’s change in the standard deviation of students’ utilities relative to the RSD benchmark. The largest improvements happen for the A-CEEI, A-CEEI with kludgy pricing, and Probabilistic Serial with seniority and reserves mechanisms. Table D.6 shows that this happens at the expense of priority violations, which more than $50\%$ of students experience in each of these three mechanisms. As in the hybrid priority structure results, this is a larger percentage of priority violations than we observe in the benchmark ($32.5\%$). In turn, the PMP, DA-STB, and DA-MTB mechanisms have no priority violations, but perform closer to the RSD benchmark in terms of the standard deviation of students’ utilities.
Figure D.6: The change in the standard deviation of students’ utilities under the department-first priority structure.
Notes: This figure reports the percent change in the standard deviation of students’ utilities compared to the Random Serial Dictatorship with optimal course reserves benchmark under the department-first priority structure. Results are averages across 100 runs with different random component draws.
Table D.5 shows that the percentage of students experiencing schedule envy in each mechanism is similar in comparison to the hybrid priority structure (see Tables 5 and D.3). The PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms deliver schedule envy bounded by a single course toward students of weakly lower priority, with only $8.2\%$, $4.9\%$, and $4.1\%$ experiencing any envy, respectively. The DA-STB and DA-MTB mechanisms each result in about $13\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses in the DA-STB mechanism and under $1\%$ in DA-MTB. The RSD with optimal course reserves mechanism has $14.5\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses. Just $1\%$ of students experience schedule envy in the random assignment of the Probabilistic Serial with seniority and reserves mechanism.
Table D.5: The percentage of students who experience schedule envy under the department-first priority structure.
0 courses 1 course 2 courses 3 courses 4 courses 5 courses
Pseudo-Market with Priorities 91.8 (0.8) 8.2 (0.8) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
RSD with optimal course reserves 85.5 (0.6) 12.4 (0.6) 1.8 (0.2) 0.3 (0.1) 0.02 (0.02) 0.0004 (0.002)
DA with single tie-breaking 86.8 (0.6) 11.3 (0.6) 1.6 (0.2) 0.3 (0.2) 0.01 (0.02) 0.0007 (0.007)
DA with multiple tie-breaking 87.7 (0.6) 11.9 (0.6) 0.4 (0.1) 0.002 (0.005) 0.0 (0.0) 0.0 (0.0)
A-CEEI 95.1 (0.4) 4.9 (0.4) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
A-CEEI with kludgy pricing 95.9 (0.3) 4.1 (0.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Probabilistic Serial with seniority and reserves 99.0 (0.2) 1.0 (0.2) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each mechanism under the department-first priority structure. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across $100$ runs with different random component draws.
Table D.6: The percentage of students with a priority violation under the department-first priority structure.
PMP RSD with opt. reserves DA with single tie-breaking DA with mult. tie-breaking A-CEEI A-CEEI with kludgy pricing PS w. seniority and reserves
0% 32.5% 0% 0% 64.1% 53.2% 84.8%
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority for each mechanism under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
D.4 Robustness: Student Choice Sets and Noise Term
This section performs some additional robustness checks. We take the calibrated student utility model and consider changes in two parameters: the size of student choice sets ($60$, $70$, $80$, $90$) and the standard deviation of the random component of student utility ($\sigma=0.75,1,1.25,1.5$). Figure D.7 reports the performance of the PMP mechanism compared to the RSD benchmark. The graphs show the percentages of students who are indifferent (solid line), prefer the PMP mechanism (dashed line), and prefer the RSD benchmark (dotted line). Reported simulation results are based on $50$ random component draws.
The top four graphs show comparative statics results for the size of student choice sets. The percentage of students indifferent between the PMP and RSD mechanisms is relatively constant across all sizes ($68.5\%$ to $76.3\%$), slightly lower for student choice sets of size $90$. A relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the choice set.
The bottom four graphs show comparative statics results for $\sigma$. The percentage of students indifferent between the PMP mechanism and RSD mechanism significantly declines from $\sigma=0.75$ to $\sigma=1.5$, with the largest decrease for Year 1 students ($85.7\%$ to $13.8\%$). Again, a relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the standard deviation. The benefits of the PMP mechanism are larger for noisier environments.
Overall, the PMP mechanism outperforms the RSD benchmark for all considered values of the size of student choice sets and the standard deviation of the random component of student utility. Our results are robust in response to the changes in these parameters.
Figure D.7: Additional Simulations: Student Choice Sets and Noise Term.
Notes: The performance of Pseudo-Market with Priorities (PMP) compared to the Random Serial Dictatorship (RSD) with optimal course reserves. The top four figures present the percentages of students who are Indifferent (solid line), Prefer (dashed line), and Do Not Prefer (dotted line) the allocation of PMP to the RSD benchmark for different sizes of student choice sets: $60$, $70$, $80$, and $90$. The bottom four figures present the same comparison for the standard deviation of student utility $\sigma=0.75,1,1.25,1.5$. Results are averages across $50$ runs with different random component draws.
Appendix E Supplementary Materials
The supplementary materials are organized as follows. Section E.1 describes the Mathematica programs for calibrating the student utility model in Section C and performing the simulations comparing course allocation mechanisms in Section 3.3 and Sections D.1D.4. Section E.2 provides details about the algorithm developed to quickly find the Pseudo-Market Equilibrium with Priorities for large student populations.
E.1 Program
Most simulations have been done using Mathematica 14.0.0.0 on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GB of RAM, 128 GB of SSD hard drive space, and a Windows Server 2019 Standard operating system. There are three programs:
  • File MSM-utility-model.nb implements the simulated method of moments to calibrate the parameters of the student utility model.
  • File simulations.nb determines the outcomes and compares the performance of the four mechanisms described in Section 3.2.
  • File simulations-appendix.nb determines the outcomes and compares the performance of the six alternative mechanisms described in Sections D.2.
The program MSM-utility-model.nb has the following sections:
  • Functions, Data, and Parameters sets the main parameters, loads the data from files, and defines some essential functions.
    • ShortDataCreation creates a version of the data with a small number of students for testing purposes.
    • Utility generates and returns the matrix of student utilities over courses for a given set of parameters $\theta$, $\gamma$, and $z$.
    • RSDallocation finds the outcome of the Random Serial Dictatorship with course reserves for a given matrix of student utilities, tie-breaking, and course reserves.
  • Simulated Method of Moments sets the main parameters of the method of simulation moments and defines all essential functions of the calibration process.
    • EmpiricalMoments produces the empirical moments from the data.
    • W1 and Weight calculate the moment weight matrices for the first and second stages of the calibration process (see Section C.4).
    • RandomParam generates random student choice sets and utility components.
    • SimulatedMoments and MomentSelection call RSDallocation to produce simulated allocations and select the relevant simulated moments.
    • MSMObjFn produces the weighted squared distance between the simulated moments and the empirical moments.
    • Optimization runs a process to optimize the student utility parameters with an initial set of parameters and other adjustable attributes.
    • ParameterNormalization implements the normalization of student utility parameters discussed in Section C.1.
    • SimulatedMoments WeightCalibration is an analog of SimulatedMoments function for the optimal weight calculation.
    • CalibrationTest tests whether the utility parameters are well-calibrated.
  • Run calls the Optimization function, tracks the progress, and records the calibrated parameters of the utility model into file “INPUTP.xlsx”. The section also performs a $J$-test on model fit for Section C.5.
Programs simulations.nb and simulations-appendix.nb have the same sections and similar functions. We label functions only used in simulations-appendix.nb with an asterisk.
  • Functions, Data, and Parameters sets the main parameters, loads the data from files, and defines some essential functions.
    • ShortDataCreation, RandomParam, Utility, and RSDallocation are the same as the functions described above.
    • DAmechanism finds the outcome of the Deferred Acceptance mechanism for given student utilities and a given priority order.
    • OptSetAsideFunction generates the optimal set of course reserves using the Deferred Acceptance Mechanism with single tie-breaking.
  • Allocation Generation Functions defines functions that produce the allocations for the mechanisms in Section 3 and D.2.
    • PMPallocation and PMPallocationAPPENDIX* find the outcomes of the Pseudo-Market with Priorities mechanism. They employ some supplementary functions, which are described below.
    • PMPInitialPrice and PMPInitialPriceAPPENDIX* provide an educated guess for the initial price in the search for a PMP equilibrium.
    • PMPPhaseI and PMPPhaseIAPPENDIX* implement Phase I of the algorithm to find a Pseudo-Market Equilibrium with Priorities.
    • PMPPhaseII and PMPPhaseIIAPPENDIX* implement Phase II of the algorithm to find a Pseudo-Market Equilibrium with Priorities.
    • OptAllocation and OptAllocationAPPENDIX* find each student’s demand for courses at a given set of prices.
    • OptAllocationInd and OptAllocationIndAPPENDIX* solve the optimization problem for an individual student.
    • OptAllocationFast and OptAllocationFastAPPENDIX* find each student’s demand for courses at a given set of prices using the information about prices and demand at the previous step.
    • Error calculates the market-clearing error for an allocation at a given set of prices.
    • FinalAllocation, FinalAllocationAPPENDIX*, OptAllocationFinalFeasibilityCheck, and OptAllocationFinalOptimizationCheck provide additional checks that the optimization problem for each student is correctly solved for the final allocation and prices.
    • DASTBallocation and DAMTBallocation find the outcomes of the Deferred Acceptance mechanism with single and multiple tie-breakings, respectively.
    • ACEEIallocation* and Kludgyallocation* find the outcomes of the Approximate Competitive Equilibrium from Equal Incomes (A-CEEI) and the A-CEEI with kludgy pricing mechanisms, respectively.
    • DAreservesmechanism* finds the outcome of the Deferred Acceptance with minority reserves for given student utilities, priority order, and course reserves.
    • DASTBreserveallocation* and DAMTBreserveallocation* find the outcomes of Deferred Acceptance mechanism with minority reserves mechanisms with single and multiple tie-breakings, respectively.
    • PSseniorityreservesallocation* finds the outcome of the Probabilistic Serial Mechanism with Seniority and Reserves mechanism.
  • Simulation Run is the main section of the program. It generates the allocations for each mechanism across multiple simulation rounds. The allocations are stored in the files allocations-data.txt and appendix-data.txt*. Data on course prices and excess demand for the PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms are recorded in files pmp-data.txt, aceei-data.txt*, and kludgy-data.txt*, respectively.
  • Results Setup loads the allocation data from allocation-data.txt for simulations.nb and from appendix-data.txt* for simulations-appendix.nb.
  • Preferred Mechanism Results produces results for Figures 1, D.1, and D.4.
  • Envy Results produces results for Tables 5, D.3, and D.5.
  • Priority Violations produces results for Tables D.4 and D.6.
  • Utility Results loads in data from pmp-data.txt, aceei-data.txt*, and kludgy-data.txt* and produces results for Tables D.1 and D.2 and Figures 2, 3, D.2, D.3, D.5, and D.6.
E.2 Computational Algorithm
We developed an algorithm to find a Pseudo-Market Equilibrium with Priorities, taking in student utilities $v$, student priorities prioritiespriorities, and a tie-breaking order over students studentrankstudentrank, and outputting an allocation bestdemandbestdemand and a vector of course prices besttbestt. Its main stages are described below, with the pseudocode shown on page Algorithm.
Initialization. We first reorder the set of student budgets $b$, evenly spread values from $1$ to $1+\beta=1.25$, and assign them to students in order of studentrankstudentrank. The function PMPInitialPrice takes in student utilities $v$ and provides an initial estimate for the price vector based on students’ most-preferred course schedules initdemand. Rather than working in MRM\!R-dimensional space, we leverage the approach in the proof of Theorem 1 and parameterize priority-specific prices as $t\in[0,R\overline{b}]^{M}$ . We collect these parameters in $pmpparam\leftarrow\{v,b,priorities,initdemand\}$ and initialize $boundcheck\leftarrow 0$ and $oversubcheck\leftarrow 0$, which serve as exit flags for Phases I and II of the algorithm.
Phase I. The first phase of the algorithm is a $\hat{\text{a}}t$onnement process, searching for a price vector by repeatedly adjusting prices proportionally to excess demand. If the market-clearing error remains above the theoretical bound THEORYBOUND $\approx 43.5$, this phase ends when there is no improvement in the error over a fixed number of consecutive steps. If the market-clearing error is below the theoretical bound, we stop once there is no substantial improvement (more than $1\%$) in error over the same number of consecutive steps. The algorithm sets the flag $boundcheck=1$ if the error falls below THEORYBOUND and $boundcheck=0$ otherwise. As output, Phase I returns the price vector besttbestt and the exit flag boundcheckboundcheck. The upper limit on the consecutive steps with no improvement and some other parameters of Phase I depend on nroundnround making the search process more persistent for later rounds.
The main innovation in Phase I is that demands are only updated for two subsets of students in each step. The first includes students who can no longer afford the allocation determined in the previous step. The second includes students who did not demand their most-preferred course schedule in the previous step and have a positive utility for courses with recently decreased prices (within the student’s budget). Any student outside these two subsets has her demand unchanged between steps. This has proven particularly effective, significantly shortening the search process compared to updating every student’s demand.
Phase II. The second phase starts with the best price of Phase I and reduces excess course demand to levels below those observed in the data. For each positive excess demand course and each demanding student, we use a dual program to calculate the minimum price change needed so the student no longer demands the course and apply the smallest change that fully eliminates the course’s excess demand. We repeat this process until two conditions are met: (i) the distribution of excess demand errors stochastically dominates the distribution observed in the data, and (ii) no course is oversubscribed by more than k1=4k-1=4 seats. If both (i) and (ii) are satisfied we set $oversubcheck=1$. We also update the value of boundcheckboundcheck as the market-clearing error changes. The output of Phase II includes price vector besttbestt and the exit flags boundcheckboundcheck and oversubcheckoversubcheck.
Final Allocation. After Phase II, we check each student’s demand is correctly computed at the final prices and produce the final course allocation. If errors occur, the FinalAllocation function will correct them. If errors violate the Phase I and II conditions, we set $bestdemand=\emptyset$ and restart the process with a new initial price vector.
The outer while loop of the algorithm continues as long as $boundcheck=0$ or $oversubcheck=0$. At each step, we add a small amount of noise to the initial price vector to begin the search from a slightly different starting point. The inner loop has the same exit conditions, with an additional condition on the number of rounds. If either boundcheckboundcheck or $oversubcheck=0$ after Phase II, we repeat the process. We allow up to $maxrounds=6$ iterations of Phase I and II within the inner loop making the search process more persistent during additional rounds. If the algorithm fails to find a solution after maxroundsmaxrounds iterations, we restart with a new initial price vector. Typically only 1-4 iterations occur between Phases I and II in the inner loop. We did not encounter cases requiring a restart through the outer loop for the main simulations. The simulations in the appendix Section D.4 checking the robustness of our results do encounter some challenging cases where a restart is necessary.
Algorithm Pseudo-Market with Priorities
1:student utilities $v$, student priorities prioritiespriorities, student ranking studentrankstudentrank
2:course allocation bestdemandbestdemand, prices besttbestt
3:Initialization: $b\leftarrow b(studentrank)$, $\{t_{0},initdemand\}\leftarrow\textbf{PMPInitialPrice}(v)$,
[…]
5:while $boundcheck=0$ OR $oversubcheck=0$ do
[…]
7:while ($boundcheck=0$ OR $oversubcheck=0$) AND $nrounds\leq maxrounds$ do
8:  $\{bestt,boundcheck\}\leftarrow\textbf{PMPPhaseI}(t,pmpparam,nround)$
9:  $\{bestt,boundcheck,oversubcheck\}\leftarrow\textbf{PMPPhaseII}(bestt,pmpparam)$
10:  $\{bestdemand\}\leftarrow\textbf{FinalAllocation}(bestt,pmpparam)$
11:  if $bestdemand==\emptyset$ then
12:   $boundcheck\leftarrow 0,oversubcheck\leftarrow 0,nround\leftarrow maxrounds$
13:  end if
14:  $t\leftarrow bestt$, $nround\leftarrow nround+1$
15:end while
16:end while